### v1

# Double-coset zeta functions for groups acting on trees

Abstract.
We study the double-coset zeta functions for groups acting on trees, focusing mainly on weakly locally $\infty$-transitive or (P)-closed actions. After giving a geometric characterisation of convergence for the defining series, we provide explicit determinant formulae for the relevant zeta functions in terms of local data of the action. Moreover, we prove that evaluation at $-1$ satisfies the expected identity with the Euler–Poincaré characteristic of the group. The behaviour at $-1$ also sheds light on a connection with the Ihara zeta function of a weighted graph introduced by A. Deitmar.
Key words and phrases:
trees, graphs, group actions, zeta functions, t.d.l.c. groups, Euler–Poincaré characteristic, local–to–global approach
2010 Mathematics Subject Classification:
05C05, 20J05, 20J06, 22D05, 30B50
1. Introduction
Background and motivation
Double cosets play a prominent role in multiple aspects of group theory and beyond. For instance, they are the building blocks of Hecke algebras. Regarded as collections of cosets, they describe spheres with respect to the Weyl distance in the building associated with a Bruhat decomposition of a group [1]. Strictly related objects, namely the suborbits, are also widely studied in permutation group theory. It is common to arrange the suborbit sizes – provided they are all finite – in a non-decreasing sequence $(a_{n})_{n\geq 1}$ and estimate, for example, the growth of $a_{n}$ as a function of $n$ [4] [21, §5].
The present paper focuses on an alternative approach to the suborbit (or double-coset) growth, recently introduced by I. Castellano, G. Chinello and T. Weigel [8]. We briefly outline it here in the slightly more general framework we employ. Let $G$ be a group with subgroups $H,K\leq G$ satisfying $|HgK/K|<\infty$ for every $g\in G$. Note that $|HgK/K|$ is the size of the $H$-orbit of $gK$ in the coset space $G/K$. Such a triple $(G,H,K)$ has the double-coset property if, for every $n\geq 1$,
(1.1) $a_{n}(G,H,K):=|\{HgK\in H\backslash G/K\,:\,|HgK/K|=n\}|<\infty.$
If for every $n\geq 1$, $a_{n}(G,H,K)$ is finite and polynomially bounded as a function in $n$, we say that $(G,H,K)$ has polynomial double-coset growth.
If $(G,H,K)$ has polynomial double-coset growth, we can consider the Dirichlet series generated by $(a_{n}(G,H,K))_{n\geq 1}$, that is,
(1.2) $\zeta_{G,H,K}(s):=\sum_{n=1}^{\infty}a_{n}(G,H,K)\cdot n^{-s}=\sum_{HgK\in H\backslash G/K}|HgK/K|^{-s},$
where $s$ is a complex variable. By [14, §I.1, Theorem 3], this series converges in some half-plane $\{s\in\mathbb{C}\mid\mathrm{Re}(s)>\alpha\}$ and the function it determines is called the double-coset zeta function of $(G,H,K)$.
Zeta functions initially arose in number theory, although nowadays they are also an established tool in studying groups, rings and algebras (see M. du Sautoy’s survey [19] for motivation). In group theory, the seminal work of F. Grunewald, D. Segal, G. Smith [13] initiated the study of numerous zeta functions associated to finitely generated nilpotent groups or profinite groups. Double-coset zeta functions have been one of the first instances of zeta functions in the class of totally disconnected locally compact (= t.d.l.c.) groups which are possibly neither discrete nor profinite. Their introduction was motivated by an interesting behaviour at $s=-1$. I. Castellano, G. Chinello and T. Weigel [8] have provided examples of unimodular t.d.l.c. groups $G$ with a compact open subgroup $K\leq G$ for which the meromorphic continuation of $\zeta_{G,K,K}(s)$, evaluated at $s=-1$, recovers the Euler–Poincaré characteristic $\widetilde{\chi}_{G}$ of $G$ (in the sense of [8, §5]). Namely, one has
(1.3) $\widetilde{\chi}_{G}=\zeta_{G,K,K}(-1)^{-1}\mu_{K},$
where $\mu_{K}$ denotes the left Haar measure on $G$ normalised with respect to $K$. A pair $(G,K)$ for which (1.3) holds is said to satisfy the Euler–Poincaré identity.
The connection between growth series and Euler–Poincaré characteristics is not an isolated phenomenon, see for instance [11, 12, 23] or the introduction of [8]. One of the main goals of the present paper is to prove that the Euler–Poincaré identity holds in two relevant classes of unimodular t.d.l.c. groups acting on trees. More generally, we present a systematic study of the double-coset zeta functions for groups (not necessarily t.d.l.c.) acting on trees, including convergence criteria and explicit formulae.
In what follows, every graph $\Gamma=V\Gamma\sqcup E\Gamma$ is meant in the sense of J-P. Serre [20], and $V\Gamma$ and $E\Gamma$ denote the set of vertices and the set of edges of $\Gamma$, respectively (cf. Section 2.2). Every group action on a tree is without inversion of edges and without global fixed points. The tree of the action is always leafless and with at least one edge (cf. Section 3.1). Moreover, at least for the theorems of the introduction, the stabilisers of adjacent vertices are assumed to be incomparable with respect to the inclusion.
Two relevant properties for group actions on trees
In the present paper, we mainly consider group actions on trees with one of the following two properties.
The first property is weak local $\infty$-transitivity. A group action on a tree $(G,T)$ is weakly locally $\infty$-transitive if, for every $v\in VT$ and every path $\mathfrak{p}$ in the quotient graph $G\backslash T$, the stabiliser of $v$ acts transitively on the set of all geodesics in $T$ starting at $v$ and lifting $\mathfrak{p}$. This condition, which we introduce in Section 3.5, generalises the well-known notion of local $\infty$-transitivity to non-edge transitive group actions.
The second property we consider is (P)-closedness. The concept of (P)-closed actions on trees has been introduced by C. Banks, M.  Elder and G. Willis [2, §3]. It stems from the slightly more general concept of actions on trees with Tits’ independence property, introduced by J. Tits [25, §4.2]. The latter properties play a central role in the theory of groups acting on trees and even beyond. For instance, one may use a (P)-closed action to produce simple groups acting on trees (cf. [25, Théorème 4.5], [2, Theorem 7.3] or [17, Theorem 1.8]). Remarkably, C. Reid and S. Smith [17] provide a complete classification of (P)-closed actions on trees by their local action diagram, which is a local datum attached to each group action on a tree. In Section 3.4 we briefly recall it. With this local description, one may study several global properties of the group action through more accessible features of the associated local action diagram. In the present paper, we exploit this approach more than once.
The reader may find it convenient to keep the following example in mind.
Example 1.1.
Let $\Gamma$ be a connected graph with a function $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$ (called edge weight). Following [6, §3.5], the pair $(\Gamma,\omega)$ admits an essentially unique universal cover $(T,\pi)$, which consists of a tree $T$ and a graph epimorphism $\pi\colon T\to\Gamma$ with the following property: for all $a\in E\Gamma$ and $v\in VT$ with $\pi(v)=o(a)$, the number of edges $e\in ET$ with origin $v$ and $\pi(e)=a$ is exactly $\omega(a)$. Note that $\pi$ is not a covering map in the usual sense. Indeed, the degree $\mathrm{deg}_{T}(v)=|o^{-1}(v)|$ of every $v\in VT$ in $T$ might be greater than the degree $\mathrm{deg}_{\Gamma}(\pi(v))=|o^{-1}(\pi(v))|$ of $\pi(v)$ in $\Gamma$, namely
$\mathrm{deg}(v)=\sum_{a\in E\Gamma\,:\,o(a)=\pi(v)}\omega(a)\geq\mathrm{deg}_{\Gamma}(\pi(v)).$
The group of deck transformations of $(\Gamma,\omega)$ is
$\mathrm{Aut}_{\pi}(T):=\{\varphi\in\mathrm{Aut}(T)\mid\pi\circ\varphi=\pi\},$
where $\mathrm{Aut}(T)$ denotes the group of automorphisms of $T$. The group $\mathrm{Aut}_{\pi}(T)$ is t.d.l.c. with respect to the subspace topology induced by $\mathrm{Aut}(T)$. Moreover, the $\mathrm{Aut}_{\pi}(T)$-action on $T$ is both weakly locally $\infty$-transitive and (P)-closed (cf. Example 3.12(i) and [25, §4.2]).
Polynomial double-coset growth
The first step towards the study of double-coset zeta functions is to determine under which conditions a triple $(G,H,K)$ has the double-coset property or polynomial double-coset growth. Here we only consider the case of $G$ acting on a locally finite tree and we choose $H,K$ among the stabilisers of vertices or edges of $T$ (written $H=G_{t_{1}}$ and $K=G_{t_{2}}$, for $t_{1},t_{2}\in T$). In this setting, we provide the following characterisation.
Theorem A (cf. Theorems 5.6 and 5.7).
Let $G$ be a group that acts on a locally finite tree $T$ with a finite quotient graph. Assume that the action $(G,T)$ is weakly locally $\infty$-transitive or (P)-closed. Then the following are equivalent, for all $t_{1},t_{2}\in T$:
  • (i)
    $(G,G_{t_{1}},G_{t_{2}})$ has the double-coset property;
  • (ii)
    $(G,G_{t_{1}},G_{t_{2}})$ has polynomial double-coset growth;
  • (iii)
    there is $k\geq 1$ such that, for every geodesic $\mathfrak{p}$ in $T$ of length $l\geq 1$, the pointwise stabiliser of $\mathfrak{p}$ does not fix any geodesic in $T$ of length $l+k$ extending $\mathfrak{p}$.
Theorem A implies that both the double-coset property and the polynomial double-coset growth are independent from the choice of $t_{1}$ and $t_{2}$. A similar independence has been shown in [8, Proposition 6.2].
It is worth mentioning that the chain of implications (iii)$\Rightarrow$(ii)$\Rightarrow$(i) is true even when dropping the hypothesis of weak local $\infty$-transitivity or (P)-closedness (cf. Proposition 5.5).
Explicit determinant formulae
The main motivation for considering weakly locally $\infty$-transitive or (P)-closed actions on locally finite trees is the following: provided $H,K$ are either vertex or edge stabilisers, we can count the $(H,K)$-double-cosets (or the $K$-cosets) and compute each size $|HgK/K|$ in terms of convenient local data of the action (cf. Sections 4).
In the weakly locally $\infty$-transitive case, counting the $(H,K)$-double cosets is rephrased in a more accessible counting of certain paths in the quotient graph $\Gamma$ with a suitably defined weight (cf. Section 3.2). This also suggests the definition of a more general Dirichlet series $\mathcal{Z}_{\Gamma,u_{1}\to u_{2}}(s)$ associated to an arbitrary graph $\Gamma$ with edge weight $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$ and to $u_{1},u_{2}\in\Gamma$ (cf. Definition 5.9). The series $\mathcal{Z}_{\Gamma,u_{1}\to u_{2}}(s)$ recovers $\zeta_{G,G_{t_{1}},G_{t_{2}}}(s)$ whenever $G$ is a group acting weakly locally $\infty$-transitively on a locally finite tree $T$ with quotient graph $\Gamma$ and standard edge weight $\omega$ (cf. Section 3.2), and whenever $t_{1},t_{2}\in T$ satisfy $G\cdot t_{1}=u_{1}$ and $G\cdot t_{2}=u_{2}$.
In the (P)-closed case, we proceed in a similar manner, except that we count $K$-cosets instead of $(H,K)$-double cosets. This has only a minor impact of $\zeta_{G,H,K}(s)$, as discussed in Section 5.1. In contrast to $(H,K)$-double-cosets, the $K$-cosets can be enumerated using suitable weighted paths in the local action diagram associated to the action (cf. Definitions 3.4 and 4.8).
In both cases, we can borrow ideas and techniques from graph theory (e.g., counting paths in graphs by their weight, see for instance [9] or [10]) and provide explicit formulae for $\zeta_{G,G_{t_{1}},G_{t_{2}}}(s)$, for all $t_{1},t_{2}\in T$.
In the following, we label $\zeta_{G,G_{t_{1}},G_{t_{2}}}(s)$ with a superscript $\bullet\in\{(w),(p)\}$ to distinguish whether $(G,T)$ is weakly locally $\infty$-transitive ($\bullet=(w)$) or (P)-closed ($\bullet=(p)$).
Theorem B (cf. Theorems 5.12 and 5.19).
Let $(G,T)$ be a group action on a tree that is weakly locally $\infty$-transitive or (P)-closed. Let $t_{1},t_{2}\in T$ be such that $(G,G_{t_{1}},G_{t_{2}})$ has polynomial double-coset growth. Then,
$\zeta_{G,G_{t_{1}},G_{t_{2}}}^{\bullet}(s)=\frac{\det(I^{\bullet}-\mathcal{E}^{\bullet}(s)+\mathcal{U}^{\bullet}_{t_{1},t_{2}}(s))}{\det(I-\mathcal{E}^{\bullet}(s))}+\epsilon^{\bullet}_{t_{1}}(t_{2}),$
for explicitly defined square matrices $\mathcal{E}^{\bullet}(s)$ and $\mathcal{U}^{\bullet}_{t_{1},t_{2}}(s)$ whose entries are entire functions in $s\in\mathbb{C}$, and for a determined integer $\epsilon^{\bullet}_{t_{1}}(t_{2})$. Here $I^{\bullet}$ denotes the identity matrix of the same dimension of $\mathcal{E}^{\bullet}(s)$. In particular, $\zeta^{\bullet}_{G,G_{t_{1}},G_{t_{2}}}(s)$ extends to a meromorphic function over $\mathbb{C}$.
In Theorem B, the matrix $\mathcal{E}^{\bullet}(s)$ can be interpreted as a weighted adjacency matrix of the local structure in which we count the paths (cf. Definitions 5.10 and 5.17). The matrix $\mathcal{U}_{t_{1},t_{2}}^{\bullet}(s)$ and the integer $\epsilon^{\bullet}_{t_{1}}(t_{2})$ can be regarded as “perturbation data” given by the choice of $t_{1}$ and $t_{2}$.
The explicit formulae in Theorem B and the fact that $\epsilon_{t_{1}}(t_{1})=0$, for every $t_{1}\in T$, yield the following:
Corollary C.
Under the hypotheses of Theorem B, let $t\in T$. Then the poles (resp. zeros) of $\zeta_{G,G_{t},G_{t}}^{\bullet}(s)$ are all those $s\in\mathbb{C}$ such that $1$ is an eigenvalue of $\mathcal{E}^{\bullet}(s)$ (resp. $\mathcal{E}^{\bullet}(s)-\mathcal{U}_{t,t}^{\bullet}(s)$) but not an eigenvalue of $\mathcal{E}^{\bullet}(s)-\mathcal{U}_{t,t}^{\bullet}(s)$ (resp. $\mathcal{E}^{\bullet}(s)$).
After Corollary C, the following question arises.
Question D.
Under the hypotheses of Theorem B, let $t\in T$. For which $s\in\mathbb{C}$ do the matrices $\mathcal{E}^{\bullet}(s)$ and $\mathcal{E}^{\bullet}(s)-\mathcal{U}_{t,t}^{\bullet}(s)$ have $1$ as an eigenvalue? Provided $\zeta_{G,G_{t},G_{t}}^{\bullet}(s)$ is an infinite series, what is its abscissa of convergence (that is, the maximal $r\in\mathbb{R}$ such that $\zeta_{G,G_{t},G_{t}}^{\bullet}(s)$ has a pole at $s=r$)?
The behaviour at $s=-1$ and the Euler–Poincaré characteristic
One of the main goals of the paper is the study of the local behaviour at $s=-1$ of the relevant double-coset zeta functions. Unless it is a finite sum, the Dirichlet series in (1.2) does not converge at $s=-1$. This underlines the importance of having a continuation of $\zeta_{G,H,K}(s)$ at least to $s=-1$ – which in our context is provided by Theorem B – to carry out such an evaluation.
Addressing [8, Question G(b)], we prove that the Euler–Poincaré identity also holds in our framework. A crucial step towards this goal is to reduce the evaluation at $s=-1$ to the more accessible weakly locally $\infty$-transitive case (cf. Lemma 7.10). More specifically, we can only focus on $\mathcal{Z}_{\Gamma,u\to u}(-1)^{-1}$ for a finite connected edge-weighted graph $(\Gamma,\omega)$. After introducing suitable notions of unimodularity and Euler–Poincaré characteristic $\chi(\Gamma,u)$ at $u\in\Gamma$ on $(\Gamma,\omega)$ (cf. Definition 7.4), we deduce the following.
Theorem E.
Let $\Gamma$ be a finite connected non-empty graph with no cycles of length $\geq 2$, and let $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 2}$ be an edge weight satisfying $\omega(a)\geq 3$ or $\omega(\bar{a})\geq 3$, for every $a\in E\Gamma$. If $(\Gamma,\omega)$ is unimodular, then
$\chi(\Gamma,u)=\mathcal{Z}_{\Gamma,u\to u}(-1)^{-1},\quad\forall\,u\in\Gamma.$
In Theorem E, the hypothesis on $\omega$ guarantees that a formula analogous to the one in Theorem B is applicable to $\mathcal{Z}_{\Gamma,u\to u}(s)$, for all $u\in\Gamma$. The proof of Theorem E is based on some splitting formulae satisfied by $\mathcal{Z}_{\Gamma,u\to u}(s)^{-1}$, $u\in\Gamma$, that we discuss in Section 6. The assumption that $\Gamma$ has no cycles of length $\geq 2$ guarantees their general applicability. By Theorem E and Lemma 7.10, we deduce the following.
Corollary F.
Let $G$ be a unimodular t.d.l.c. group acting on a locally finite tree $T$ with compact open vertex stabilisers. Assume that the quotient graph is finite and does not have cycles of length $\geq 2$. Suppose also that $(G,T)$ is weakly locally $\infty$-transitive or (P)-closed. Then, for every $t\in T$ such that $(G,G_{t},G_{t})$ has polynomial double-coset growth, we have
$\widetilde{\chi}_{G}=\zeta_{G,G_{t},G_{t}}(-1)^{-1}\mu_{G_{t}}.$
The behaviour at $s=-1$ and the weighted Ihara zeta function
The zeta function $\mathcal{Z}_{\Gamma,u\to u}(s)$ is not the only growth series that has been considered for a weighted graph $(\Gamma,\omega)$. Another relevant example is the weighted Ihara zeta function $Z_{(\Gamma,W)}(x)$ introduced by A. Deitmar [9] for every graph $\Gamma$ with transition weight $W$ (cf. Section 7.3 for a brief recap). In Section 7.3, we provide a canonical way to construct a transition weight $W_{(\Gamma,\omega)}$ on a finite graph $\Gamma$ starting from an edge weight $\omega$ (cf. Example 7.11). The main result of the section focuses on a finite connected non-empty graph $\Gamma$ and subgraphs $\Gamma_{1},\Gamma_{2}\subseteq\Gamma$ satisfying the following: $\Gamma=\Gamma_{1}\cup\Gamma_{2}$, $\Gamma_{1}\cap\Gamma_{2}$ is a $1$-segment graph with edge set $\{a,\bar{a}\}$, and $t(a)$ and $o(a)$ are terminal vertices of $\Gamma_{1}$ and $\Gamma_{2}$, respectively. The graphs $\Gamma_{1}$ and $\Gamma_{2}$ carry restricted edge weights $\omega_{1}$ and $\omega_{2}$ from $\omega$, respectively. Hence we prove the following.
Theorem G (cf. Theorem 7.12).
In the setting before, assume that $\omega(E\Gamma)\subseteq\mathbb{Z}_{\geq 2}$ and that $\omega(a)\geq 3$ or $\omega(\bar{a})\geq 3$, for every $a\in E\Gamma$. Then,
(1.4) $\frac{\mathcal{Z}_{\Gamma,a\to a}(-1)}{\mathcal{Z}_{\Gamma_{1},a\to a}(-1)\cdot\mathcal{Z}_{\Gamma_{2},a\to a}(-1)}=\frac{1}{\omega(a)\omega(\bar{a})}\cdot\frac{Z_{(\Gamma,W)}(1)}{Z_{(\Gamma_{1},W_{1})}(1)\cdot Z_{(\Gamma_{2},W_{2})}(1)}.$
In particular, if $(\Gamma,\omega)$ is unimodular then
(1.5) $\frac{\chi(\Gamma_{1},a)\cdot\chi(\Gamma_{2},a)}{\chi(\Gamma,a)}=\frac{1}{\omega(a)\omega(\bar{a})}\cdot\frac{Z_{(\Gamma,W)}(1)}{Z_{(\Gamma_{1},W_{1})}(1)\cdot Z_{(\Gamma_{2},W_{2})}(1)}.$
Structure of the paper
Sections 2 and 3 collect background knowledge for the paper. In particular, in Section 3.5 we introduce the new concept of weakly locally $\infty$-transitive actions on trees.
In Section 4, we give geometric descriptions of the coset spaces $G/G_{t}$ and the size $|G_{r}gG_{t}/G_{t}|$ of a group $G$ acting on a tree $T$ with respect to vertex or edge stabilisers $G_{r},G_{t}$. This description is furthermore refined if the action is weakly locally $\infty$-transitive (cf. Section 4.2) or (P)-closed (cf. Section 4.3).
Section 5 follows a similar pattern. It begins with general results on the double-coset property and polynomial double-coset growth for groups acting on trees (cf. Proposition 5.5). Afterwards, we specialise the discussion to the cases of weakly locally $\infty$-transitive or (P)-closed actions (cf. Sections 5.4 and 5.5, respectively). In these two cases, we characterise the polynomial double-coset growth and give explicit determinant formulae of the relevant double-coset zeta functions.
Section 6 collects some splitting formulae for $\mathcal{Z}_{\Gamma,u\to u}(s)^{-1}$ which are key for the proofs in the next Section 7.
Finally, in Section 7 we discuss the behaviour at $s=-1$ for the double-coset zeta functions studied in Sections 5.4 and 5.5. In particular, we provide connections with the Euler–Poincaré characteristic of the group (cf. Section 7.2) and the Ihara zeta function of a weighted graph (cf. Section 7.3).
Acknowledgements
The author would like to thank her PhD advisors Christopher Voll and Thomas Weigel for the continuous support and instructive inputs throughout the preparation of this paper. The author also would like to thank Ilaria Castellano, Jan Moritz Petschick, José Pedro Quintanilha and Karthika Rajeev for interesting conversations and several useful comments on an early draft of the paper, and Evgeny Khukhro and Roman Sauer for interesting questions that inspired the results in Section 6.
The author was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 491392403 – TRR 358. The author is a member of Gruppo Nazionale per le Strutture Algebriche, Geometriche e le loro Applicazioni (GNSAGA), which is part of the Istituto Nazionale di Alta Matematica (INdAM).
2. Notation and preliminaries
2.1. Generalities
Whenever there is no ambiguity, we denote a $1$-point set $\{x\}$ by the element $x$ itself. Given a set $X$, for every subset $A\subseteq X$ denote by $\mathbbm{1}_{A}$ the indicator function of $A$. For every set $X$, the symmetric group $\mathrm{Sym}(X)$ is always regarded as a topological group with the permutation topology, i.e., the topology generated by the local basis at $1$ given by all possible pointwise stabilisers of finite subsets of $X$.
Moreover, let $G$ and $H$ be groups acting on the sets $X$ and $Y$, respectively. Denote by $\sigma_{G}\colon G\to\mathrm{Sym}(X)$ and $\sigma_{H}\colon H\to\mathrm{Sym}(Y)$ the homomorphisms induced by the two actions. The $G$-action on $X$ is said to be permutational isomorphic to the $H$-action on $Y$ if there are a group isomorphism $\varphi\colon\sigma_{G}(G)\to\sigma_{H}(H)$ and a bijection $f\colon X\to Y$ satisfying $f(g\cdot x)=\varphi(g)\cdot f(x)$, for all $x\in X$ and $g\in\sigma_{G}(G)$.
2.2. Graphs
A graph (in the sense of J-P. Serre [20]) consists of a set $\Gamma=V\Gamma\sqcup E\Gamma$ partitioned into two subsets $V\Gamma$ and $E\Gamma$ (called the set of vertices and the set of edges of $\Gamma$, respectively), together with two maps $o,t\colon E\Gamma\to V\Gamma$ (called origin and terminus maps, respectively) and an involution ¯$\overline{\cdot}\colon E\Gamma\to E\Gamma$ (called edge inversion) satisfying $\bar{e}\neq e$ and $o(\bar{e})=t(e)$, for every $e\in E\Gamma$. We introduce the following notation for a graph $\Gamma$:
Notation 2.1.
Given $u\in\Gamma$, we use the associated capital letter $U$ to denote the set $\{u\}$ if $u\in V\Gamma$, and the set $\{u,\bar{u}\}$ if $u\in E\Gamma$.
An orientation in a graph $\Gamma$ is a set $E\Gamma^{+}\subseteq E\Gamma$ satisfying |$|\{e,\bar{e}\}\cap E\Gamma^{+}|=1$ for every $e\in E\Gamma$, and $E\Gamma=E\Gamma^{+}\cup\{\bar{e}\mid e\in E\Gamma^{+}\}$. A graph $\Gamma$ is non-empty if it has at least one vertex, and is locally finite if $|o^{-1}(c)|<\infty$ for all $c\in V\Gamma$. A subgraph of a graph $\Gamma$ is a subset $\Lambda$ such that $o(E\Gamma\cap\Lambda),t(E\Gamma\cap\Lambda)\subseteq V\Gamma\cap\Lambda$ and $\overline{e}\in E\Gamma\cap\Lambda$ for every $e\in E\Gamma\cap\Lambda$. The subset $\Lambda$ inherits a graph structure from $\Gamma$. A subgraph $\Lambda$ of $\Gamma$ is proper if $\Lambda\neq\Gamma$. A vertex $v$ in $\Gamma$ is said to be terminal if $|o^{-1}(v)|=1$. An edge $e$ in $\Gamma$ with $o(e)=t(e)$ is called $1$-loop. A $n$-bouquet of loops (based at $c$) is a graph with one vertex $c$ and edge-set $\{a_{i},\bar{a}_{i}\mid 1\leq i\leq n\}$, where each $a_{i}$ is a $1$-loop starting at $c$. A $1$-segment is a graph $\Gamma$ with two distinct vertices and an edge-couple $\{e,\bar{e}\}$ connecting them.
Given two graphs $\Gamma$ and $\Lambda$, a graph morphism is a map $\varphi\colon\Gamma\to\Lambda$ satisfying $\varphi(V\Gamma)\subseteq V\Lambda$, $\varphi(E\Gamma)\subseteq E\Lambda$, $\varphi(o(e))=o(\varphi(e))$ and $\varphi(\bar{e})=\overline{\varphi(e)}$ for every $e\in E\Gamma$. A graph monomorphism (resp. epimorphism, isomorphism) is a graph morphism which is injective (resp. surjective, bijective). Given a graph $\Gamma$, let $\mathrm{Aut}(\Gamma)$ be the group of all automorphisms (= self-isomorphisms) of $\Gamma$. We always regard $\mathrm{Aut}(\Gamma)$ as a topological group with the subspace topology induced by $\mathrm{Sym}(\Gamma)$.
Let $\Gamma$ be a graph. A path in $\Gamma$ is a sequence of vertices and edges $\mathfrak{p}=(v_{0},e_{1},v_{1},\ldots,e_{n},v_{n})$, $n\geq 0$, with $o(e_{i})=v_{i-1}$ and $t(e_{i})=v_{i}$ for every $1\leq i\leq n$. We say that $\mathfrak{p}$ starts at $v_{0}$ (or at $e_{1}$) and ends at $v_{n}$ (or at $e_{n}$), has reverse path is $\overline{\mathfrak{p}}=(v_{n},\bar{e}_{n},v_{n-1},\ldots,\bar{e}_{1},v_{0})$, and length $n$ (written $\ell(\mathfrak{p})=n$). If $n\geq 1$, we may without ambiguity specify only the sequence of edges. If $n=0$, the $1$-term sequence $\mathcal{O}_{v_{0}}=(v_{0})$ is called the trivial path at $v_{0}$. Denote by $\mathcal{P}_{\Gamma}$ the set of all paths in $\Gamma$. Given non-empty subsets $X,Y\subseteq\Gamma$, let $\mathcal{P}_{\Gamma}(X\to Y)$ be the set of all paths in $\Gamma$ starting at some $x\in X$ and ending at some $y\in Y$. The product of two paths $\mathfrak{p}=(e_{1},\ldots,e_{m})$ and $\mathfrak{q}=(f_{1},\ldots,f_{n})$ is defined only if $t(e_{m})=o(f_{1})$ and it is the path $\mathfrak{p}\cdot\mathfrak{q}=(e_{1},\ldots,e_{m},f_{1},\ldots,f_{n})$. If $\mathfrak{p}$ is a path starting and ending at the same vertex, denote by $\mathfrak{p}^{d}$ the $d$-th power of $\mathfrak{p}$ with respect to the product defined before. A path $\mathfrak{p}$ is reduced if either $\ell(\mathfrak{p})=0$ or $\mathfrak{p}=(e_{1},\ldots,e_{n})$ and $e_{i+1}\neq\bar{e}_{i}$ for every $11\leq i\leq n-$ 1. For $n\geq 1$, an $n$-cycle is a reduced path $\mathfrak{p}=(e_{1},\ldots,e_{n})$ with $o(e_{1})=t(e_{n})$ and $t(e_{i})\neq t(e_{j})$ for all $11\leq i,j\leq n-$ 1 with $i\neq j$.
A graph $\Gamma$ is connected if for all $v,w\in V\Gamma$ there is a path from $v$ to $w$. A subgraph of $\Gamma$ is a connected component if it is a maximal connected subgraph of $\Gamma$. A graph is the disjoint union of all its connected components. A tree $T$ is a connected graph with no $n$-cycles, for every $n\geq 1$. If $T$ is a tree and $e\in ET$, then the graph $T\setminus\{e,\bar{e}\}$ has two connected components, $T_{e}^{+}\ni t(e)$ and $T_{e}^{-}\ni o(e)$. Set $T_{\geq e}:=T_{e}^{+}\sqcup\{e\}$ and $T_{\geq\bar{e}}:=T_{e}^{-}\sqcup\{\bar{e}\}$. A tree $T$ is uniquely geodesic, i.e., for all $v,w\in VT$ there is a unique reduced path $[v,w]$ from $v$ to $w$, which we call geodesic from $v$ to $w$. Recall that $[v,w]$ is the path of minimal length in $T$ from $v$ to $w$. Moreover, given $e,f\in ET$, there is a geodesic $(e_{1},\ldots,e_{n})$ in $T$ with $e_{1}=e$ and $e_{n}=f$ if, and only if, $f\in T_{\geq e}$. In general, for $t_{1},t_{2}\in T$ we denote by $[t_{1},t_{2}]$ the geodesic from $t_{1}$ to $t_{2}$ in $T$ (whenever it exists). Moreover, for non-empty subsets $X,Y\subseteq T$, denote by $\mathrm{Geod}_{T}(X\to Y)$ the set of all geodesics in $T$ from some $x\in X$ to some $y\in Y$. Finally, a ray in a tree is a sequence of edges $(e_{i})_{i\in\mathbb{Z}_{\geq 1}}$ such that $o(e_{i})\neq o(e_{j})$ and $t(e_{i})=o(e_{i+1})$, for all $i,j\in\mathbb{Z}_{\geq 1}$ with $i\neq j$.
Remark 2.1.1.
Let $\Gamma$ be a connected graph without $n$-cycles for every $n\geq 2$. Then $\Gamma$ has a unique maximal subtree: the subgraph $\Lambda$ obtained from $\Gamma$ by removing all its $1$-loops. In particular, for all $v,w\in V\Gamma$ the geodesic $[v,w]=(v=v_{0},e_{1},v_{1},\ldots,e_{n},v_{n}=w)$ in $\Lambda$ is the path of minimal length in $\Gamma$ from $v$ to $w$. Thus, if $v\neq w$ then $v_{i}\neq v_{j}$ for all $0\leq i,j\leq n$ with $i\neq j$.
3. Group actions on trees
In this section, we introduce the two main classes of group actions on trees considered in this paper. The first is the class of (P)-closed actions on trees (cf. Section 3.3). In Section 3.4, we briefly recall a local–to–global approach due C. Reid and S. Smith [17] to study these kinds of group actions, which will be largely used in the paper. We also add some new vocabulary (cf. Definition 3.4) that will be exploited in Section 4.3 to count geodesics on the tree of those actions. Section 3.5 introduces the second class we focus on, the one of weakly locally $\infty$-transitive actions on trees. Therein we discuss the connection with locally $\infty$-transitive actions (cf. Lemma 3.10), and provide a local characterisation and some explicit examples (cf. Proposition 3.11 and Example 3.12).
3.1. Group actions on trees
Let $T$ be a tree with $ET\neq\emptyset$ and without leaves (i.e., no vertices $v\in VT$ have $|o^{-1}(v)|=1$), and let $G$ be a topological group. A $G$-action $(G,T)$ on $T$ is a continuous group homomorphism $G\to\mathrm{Aut}(T)$ satisfying, for all $g\in G$ and $e\in ET$, that $g\cdot e\neq\bar{e}$ and, for all $v\in VT$, that $g\cdot v\neq v$ for some $g\in G$. In the literature, the latter two requirements are often added as separate conditions called “acting without edge inversions” and “acting without global fixed points”, respectively.
Given $t\in T$, denote by $G_{t}$ the stabiliser of $t$ in $G$. More generally, for every subset $X\subseteq T$, let $G_{X}$ denote the pointwise stabiliser of $X$. If $\mathfrak{p}=(e_{i})_{1\leq i\leq n}$ is a path, $G_{\mathfrak{p}}$ denotes the pointwise stabiliser of the set $\{e_{1},\ldots,e_{n}\}$.
An action $(G,T)$ is edge-transitive if $ET=G\cdot e\sqcup G\cdot\bar{e}$ for some (and hence every) $e\in ET$. Moreover, $(G,T)$ is locally $\infty$-transitive if, for every $v\in VT$ and $d\geq 0$, the stabiliser $G_{v}$ acts transitively on $\{\mathfrak{p}\in\mathrm{Geod}_{T}(v\to T)\mid\ell(\mathfrak{p})=d\}$ (cf. [5, §0.2]). One checks that locally $\infty$-transitive actions are edge-transitive. Examples of groups admitting a locally $\infty$-transitive action on a tree are the $k$-points of simple simply connected algebraic $k$-group of relative rank $1$, where $k$ is a non-Archimedean local field (cf. [20, pp. 91 and 95]), and the Burger–Mozes universal groups $U(F)$ associated to $2$-transitive groups $F\leq\mathrm{Sym}(\{1,\ldots,d\})$ acting on the barycentric subdivision of a $d$-regular tree (cf. the lines before [5, §3.1]).
3.2. The quotient graph and its standard edge weight
Let $(G,T)$ be a group action on a tree. The quotient graph $\Gamma=G\backslash T$ of the action is the graph with $V\Gamma:=G\backslash VT$, $E\Gamma:=G\backslash ET$ and, given $G\cdot e\in E\Gamma$, its origin is $G\cdot o(e)$, its terminus is $G\cdot t(e)$ and its inverse edge is $G\cdot\bar{e}$. One checks that these definitions are independent from the choice of $e$ in $G\cdot e$.
The assignment $\pi\colon t\in T\longmapsto G\cdot t\in\Gamma$ yields a graph epimorphism which is called the quotient map of $(G,T)$. The map $\pi$ entrywise extends to a map (denoted with the same symbol) from the set of all paths in $T$ to the set of all paths in $\Gamma$.
The standard edge weight on $\Gamma$ is the map $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}\cup\{\infty\}$ defined on every $a\in E\Gamma$ by choosing $v\in VT$ with $\pi(v)=o(a)$ and setting
(3.1) $\omega(a):=|\{e\in ET\,:\,o(e)=v\text{ and }\pi(e)=a\}|.$
In other words, $\omega(a)$ counts how many edges in $T$ starting at $v$ lift $a$ via $\pi$. It is straightforward to check that the assignment in (3.1) does not depend on the choice of the vertex $v$.
Starting from a connected graph $\Gamma$ and a function $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}\cup\{\infty\}$, one always has a group action on a tree with quotient graph $\Gamma$ and standard edge weight $\omega$ (cf. Example 1.1).
3.3. (P)-closed actions on trees
The study of the group $\mathrm{Aut}_{\pi}(T)$ as in Example 1.1 initiated the study of (P)-closed group actions on trees (cf. [2]), a class which stems from the more general class of group actions on trees with the Tits’ independence property (cf. [25, §4.2]). A group action on a tree $(G,T)$ is (P)-closed if $G\leq\mathrm{Aut}(T)$ is closed and, for every $e\in ET$,
(3.2) $G_{e}=G_{T_{\geq\bar{e}}}\cdot G_{T_{\geq e}}.$
Note that the inclusion $\supseteq$ is automatic.
Proposition 3.1.
Let $(G,T)$ be a (P)-closed action on a tree and $(e_{1},\ldots,e_{n})$ be a geodesic in $T$ of length $n\geq 2$. Then, for every $k<n$ we have
(3.3) $G_{(e_{1},\ldots,e_{k})}\cdot(e_{k+1},\ldots,e_{n})=G_{e_{k}}\cdot(e_{k+1},\ldots,e_{n}).$
Proof.
The inclusion $\subseteq$ is clear. Moreover, we note that $e_{1},\ldots,e_{k-1}\in ET_{\geq\bar{e}_{k}}$ and $e_{k+1},\ldots,e_{n}\in ET_{\geq e_{k}}$. Hence,
$\begin{split}G_{(e_{1},\ldots,e_{k})}\cdot(e_{k+1},\ldots,e_{n})&\supseteq G_{T_{\geq\bar{e}_{k}}}\cdot(e_{k+1},\ldots,e_{n})\\ =G_{T_{\geq\bar{e}_{k}}}\cdot G_{T_{\geq e_{k}}}\cdot(e_{k+1},\ldots,e_{n})&=G_{e_{k}}\cdot(e_{k+1},\ldots,e_{n}).\qed\end{split}$
C. Reid and S. Smith [17] provide a parametrisation of (P)-closed group actions on trees in terms of local action diagrams, that we now recall.
3.4. Local action diagrams and their associated universal groups
Following [17, Definition 3.1], a local action diagram is a triple
$\Delta=(\Gamma,(X_{a})_{a\in E\Gamma},(G(c))_{c\in V\Gamma})$
consisting of the following data:
  • (i)
    a connected graph $\Gamma$;
  • (ii)
    a family of non-empty pairwise disjoint sets $(X_{a})_{a\in E\Gamma}$. For every $c\in V\Gamma$, set $X_{c}=\bigsqcup_{a\in o^{-1}(c)}X_{a}$ and $X=\bigsqcup_{a\in E\Gamma}X_{a}$;
  • (iii)
    for every $c\in V\Gamma$, a closed subgroup $G(c)$ of $\mathrm{Sym}(X_{c})$ whose orbits are given by $G(c)\backslash X_{c}=\{X_{a}\}_{a\in o^{-1}(c)}$.
Notation 3.2.
If there is no ambiguity, we write $\Delta=(\Gamma,(X_{a}),(G(c)))$ in place of $\Delta=(\Gamma,(X_{a})_{a\in E\Gamma},(G(c))_{c\in V\Gamma})$. Moreover, given $u\in\Gamma$, set $X_{U}:=X_{u}$ if $u\in V\Gamma$ and $X_{U}:=X_{u}\sqcup X_{\bar{u}}$ if $u\in E\Gamma$.
Local action diagrams can be constructed from a group action on a tree $(G,T)$ as follows: Let $\pi$ be the quotient map on $(G,T)$ and choose a set of representatives $V^{*}$ of the $G$-orbits on $VT$. Following [17, Definition 3.6], the local action diagram associated to $(G,T)$ and $V^{*}$ is defined as follows:
  • (i)
    $\Gamma=G\backslash T$ is the quotient graph of $(G,T)$;
  • (ii)
    for every $a\in E\Gamma$, let $v^{*}\in V^{*}$ be such that $\pi(v^{*})=o(a)$ and define
    $X_{a}:=\{e\in ET\mid o(e)=v^{*}\text{ and }\pi(e)=a\}.$
  • (iii)
    for every $c\in V\Gamma$ with representative $c^{*}\in V^{*}$, let $G(c)$ be the closure in $\mathrm{Sym}(X_{c})$ of the permutation group induced by $G_{c^{*}}$ acting on $X_{c}$.
Note that the standard edge weight on the quotient graph $\Gamma$ is given by $\omega(a)=|X_{a}|$, for every $a\in E\Gamma$. Up to isomorphism of local action diagrams (cf. [17, Definition 3.2]), every group action on a tree $(G,T)$ has a unique associated local action diagram (cf. [17, Lemma 3.7]). Thus we refer to the local action diagram associated to $(G,T)$. Moreover, one of the key results in [17] (recalled in Theorem 3.8) shows that every local action diagram arises as the local action diagram associated to a group action on a tree.
Example 3.3.
  • (i)
    Let $\Gamma$ be a connected graph with an edge weight $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$. Let $T$ and $G=\mathrm{Aut}_{\pi}(T)$ as in Example 1.1. Then the local action diagram $\Delta=(\Gamma,(X_{a}),(G(c)))$ associated to $(G,T)$ is given by taking, for every $a\in E\Gamma$, a set $X_{a}$ of cardinality $\omega(a)$, and by setting
    $G(c):=\{\sigma\in\mathrm{Sym}(X_{c})\mid\forall\,a\in o^{-1}(c),\,\sigma(X_{a})=X_{a}\},$
    for every $c\in V\Gamma$. By design, $G(c)\backslash X_{c}=\{X_{a}\}_{a\in o^{-1}(c)}$. More precisely, for every $a\in o^{-1}(c)$, the $G(c)$-action on $X_{a}$ is permutational isomorphic to the action of $\mathrm{Sym}(X_{a})$ on the same set.
    Note that $G(c)$ is closed in $\mathrm{Sym}(X_{c})$. Indeed, if $\sigma\in\mathrm{Sym}(X_{c})$ satisfies $\sigma(x)\not\in X_{a}$ for some $x\in X_{a}$, then $\sigma\cdot\mathrm{Sym}(X_{c})_{x}$ is an open neighbourhood of $\sigma$ in $\mathrm{Sym}(X_{c})$ which is contained in $\mathrm{Sym}(X_{c})\setminus G(c)$.
  • (ii)
    Given a prime $p$, denote by $\mathbb{Q}_{p}$ the field of $p$-adic numbers, by $\mathbb{Z}_{p}$ the ring of $p$-adic integers, and by $\mathbb{F}_{p}\simeq\mathbb{Z}_{p}/{p\mathbb{Z}_{p}}$ the field of size $p$. Here below, we report and rephrase some results from [20, §II.1.4]. The action of $G=\mathrm{SL}_{2}(\mathbb{Q}_{p})$ on its Bruhat–Tits tree $T$ – which is a $(p+1)$-regular tree – has a $1$-segment as quotient graph $\Gamma$. Set $E\Gamma=\{a,\bar{a}\}$. Moreover, there is $e\in ET$ with $G\cdot e=a$ satisfying $G_{o(e)}=\mathrm{SL}_{2}(\mathbb{Z}_{p})\simeq G_{t(e)}$. Both the $\mathrm{SL}_{2}(\mathbb{Z}_{p})$-action on $\{f\in ET\mid o(f)=o(e)\}$ and the $G_{t(e)}$-action on $\{f\in ET\mid o(f)=o(\bar{e})\}$ are permutational isomorphic to the faithful action of $\mathrm{PSL}_{2}(\mathbb{F}_{p})$ on the projective line $\mathbb{P}^{1}(\mathbb{F}_{p})$. Hence, the local action diagram $\Delta=(\Gamma,(X_{b})_{b\in E\Gamma},(G(c))_{c\in V\Gamma})$ associated to $(G,T)$ is given by setting $X_{a}=X_{\bar{a}}=\mathbb{P}^{1}(\mathbb{F}_{p})$ and $G(o(a))=G(t(a))=\mathrm{PSL}_{2}(\mathbb{F}_{p})$.
We now expand the vocabulary of local action diagrams, introducing tools of key importance for the discussion.
Definition 3.4.
Let $\Delta=(\Gamma,(X_{a}),(G(c)))$ be a local action diagram. For $n\geq 1$, an $n$-path in $\Delta$ is a sequence $\xi=(x_{1},\ldots,x_{n})$ obtained by starting with a path $\mathfrak{p}_{\xi}=(a_{1},\ldots,a_{n})$ in $\Gamma$ and selecting, for each $1\leq i\leq n$, an element $x_{i}\in X_{a_{i}}$. One says that $\xi$ starts at $x_{1}$, ends with $x_{n}$, has length $n$ (written $\ell(\xi)=n$), and $\mathfrak{p}_{\xi}$ is called the underlying path of $\xi$ in $\Gamma$. The 00-path $O_{c}$ in $\Delta$ at $c\in V\Gamma$ is the empty sequence of elements in $X$ with the trivial path at $c$ as underlying path in $\Gamma$. The path $O_{c}$ has length zero. A path in $\Delta$ is an $n$-path for some $n\geq 0$. Given paths $\xi=(x_{1},\ldots,x_{m})$ and $\eta=(y_{1},\ldots,y_{n})$ in $\Delta$ with $\mathfrak{p}_{\xi}=(a_{1},\ldots,a_{m})$ and $\mathfrak{p}_{\eta}=(b_{1},\ldots,b_{n})$, the product $\xi\cdot\eta$ is defined only if $t(a_{m})=o(b_{1})$ and it is the path $(x_{1},\ldots,x_{m},y_{1},\ldots,y_{n})$ in $\Delta$ with underlying path $\mathfrak{p}_{\xi}\cdot\mathfrak{p}_{\eta}$. Put also $\xi\cdot O_{t(a_{m})}=\xi$ and $O_{o(b_{1})}\cdot\eta=\eta$. Given a path $\xi$ and a non-empty set of paths $\mathscr{E}$ in $\Delta$ such that the product $\xi\cdot\eta$ is defined for every $\eta\in\mathscr{E}$, set $\xi\cdot\mathscr{E}=\{\xi\cdot\eta\mid\eta\in\mathscr{E}\}$. If $\xi=(x)$ has length $1$, we write $x\cdot\mathscr{E}$ in place of $(x)\cdot\mathscr{E}$.
A map $\iota\colon X\to X$ is said to be an inversion in $\Delta$ if $\iota(X_{a})\subseteq X_{\bar{a}}$ for every $a\in E\Gamma$. A path $\xi$ in $\Delta$ is reduced in $(\Delta,\iota)$ if it either $\ell(\xi)=0$ or $\xi=(x_{1},\dots,x_{n})$, for some $n\geq 1$, and $x_{i+1}\neq\iota(x_{i})$ for every $11\leq i\leq n-$ 1. Note that, even if $\xi$ is reduced, the underlying path $\mathfrak{p}_{\xi}$ needs not to be reduced. Denote by $\mathcal{P}_{(\Delta,\iota)}$ the set of all reduced paths in $(\Delta,\iota)$. For non-empty subsets $X_{1},X_{2}\subseteq X$, let also $\mathcal{P}_{(\Delta,\iota)}(X_{1}\to X_{2})$ be the collection of all reduced paths in $(\Delta,\iota)$ starting at some $x_{1}\in X_{1}$ and ending at some $x_{2}\in X_{2}$.
An inversion in a local action diagram is not required to be an involution. In fact, the sizes of $X_{a}$ and $X_{\bar{a}}$ might differ. The term “inversion” here refers to the edge inversion on the labels in the partition $X=\bigsqcup_{a\in E\Gamma}X_{a}$.
Let $\Delta=(\Gamma,(X_{a}),(G(c)))$ be a local action diagram. Following [17, Definition 3.4], a $\Delta$-tree $(T,\pi,\mathcal{L})$ consists of a tree $T$, a graph epimorphism $\pi\colon T\to\Gamma$ and a map $\mathcal{L}\colon ET\to X$ which restricts, for all $v\in VT$ and $a\in E\Gamma$ with $o(a)=\pi(v)$, to a bijection
$\mathcal{L}_{v,a}\colon\{e\in ET\mid o(e)=v\text{ and }\pi(e)=a\}\longrightarrow X_{a}.$
In particular, for every $v\in VT$, the map $\mathcal{L}$ restricts to a bijection
$\mathcal{L}_{v}\colon o^{-1}(v)\longrightarrow X_{\pi(v)}.$
Note that the definition of a $\Delta$-tree is independent from $(G(c))_{c\in V\Gamma}$.
According to [6, §3.5], the pair $(T,\pi)$ is the universal cover of the edge-weighted graph $(\Gamma,\omega)$, where $\omega(a)=|X_{a}|$ for every $a\in E\Gamma$. Therefore, for every two $\Delta$-trees $(T,\pi,\mathcal{L})$ and $(T^{\prime},\pi^{\prime},\mathcal{L}^{\prime})$ there is a graph isomorphism $\varphi\colon T\to T^{\prime}$ such that $\pi=\pi^{\prime}\circ\varphi$ (cf. [17, Lemma 3.5]). Moreover, for every $\Delta$-tree we may define $\mathrm{Aut}_{\pi}(T)$ as in Example 1.1.
3.4.1. The standard $\Delta$-tree associated to $\iota$ and $c_{0}$
Let $\Delta=(\Gamma,(X_{a}),(G(c)))$ be a local action diagram. Following the proof of [17, Lemma 3.5], we recall the construction of an explicit family of $\Delta$-trees which plays an important role in the next discussion.
Set an inversion $\iota\colon X\to X$ in $\Delta$ and $c_{0}\in V\Gamma$, and define a graph $T=T(\Delta,\iota,c_{0})$ as follows: The set of vertices of $T$ is
(3.4) $VT:=\mathcal{P}_{(\Delta,\iota)}(X_{c_{0}}\to X).$
The vertex $v_{0}=O_{c_{0}}$ is called the root of $T$. The edges of $T$ are the pairs $(v,w)$ and $(w,v)$ of reduced paths in $\Delta$ of the form $v=(x_{1},\dots,x_{n})$ and $w=(x_{1},\dots,x_{n},x_{n+1})$, for some $n\geq 0$. Every edge $(v,w)$ of $T$ as before is said to be a positive edge. Denote by $ET^{+}$ the set of all positive edges of $T$. The origin, the terminus, and the inversion maps of $T$ are given by $o(v,w)=v$, $t(v,w)=w$ and $\overline{(v,w)}=(w,v)$, for every $(v,w)\in ET$. For every $(v,w)\in ET^{+}$ with $w=(x_{1},\dots,x_{n+1})$, set $\mathcal{L}(v,w)=x_{n+1}$ and $\mathcal{L}(w,v)=\iota(x_{n+1})$.
Remark 3.4.1.
Every $e\in ET(\Delta,\iota,c_{0})^{+}$ satisfies $\mathcal{L}(\bar{e})=\iota(\mathcal{L}(e))$. This is generally not true if $e\in ET(\Delta,\iota,c_{0})\setminus ET(\Delta,\iota,c_{0})^{+}$, as a given $x\in X$ might differ from $\iota(\iota(x))$.
More generally, for every path $\mathfrak{p}=(e_{1},\dots,e_{n})$ in $T=T(\Delta,\iota,c_{0})$ we define
(3.5) $\mathcal{L}(\mathfrak{p}):=(\mathcal{L}(e_{1}),\ldots,\mathcal{L}(e_{n})).$
Remark 3.4.2.
By Remark 3.4.1, a path $\mathfrak{p}=(e_{1},\ldots,e_{n})$ in $T$ with $e_{1},\ldots,e_{n}\in ET^{+}$ is reduced, and only if, $\mathcal{L}(\mathfrak{p})$ is reduced in $(\Delta,\iota)$. Moreover, by (3.4), for every $v=(x_{1},\ldots,x_{n})\in VT$ there is a unique reduced path from $v_{0}$ to $v$ in $T$, namely $\mathcal{O}_{v_{0}}$ if $v=v_{0}$, and $(v_{0},e_{1},v_{1},\ldots,e_{n},v_{n}=v)$ with $v_{i}=(x_{1},\ldots,x_{i})$ for every $1\leq i\leq n$ otherwise. In particular, $T$ is connected.
More precisely, $T$ is a tree. In fact, if $T$ admits a $n$-cycle $\gamma$ for some $n\geq 1$, there is a vertex $v$ of $\gamma$ such that the reduced path $\mathfrak{p}$ from $v_{0}$ to $v$ shares no edges with $\gamma$. Up to an edge-relabelling, we may assume that $\gamma=(e_{1},\ldots,e_{n})$ and $o(e_{1})=v$. Then $\mathfrak{p}$ and $\mathfrak{p}\cdot\gamma$ are two distinct reduced paths in $T$ from $v_{0}$ to $v$, impossible.
Being $T$ a tree, the map $\mathcal{L}$ in (3.5) restricts to a bijection
(3.6) $\mathcal{L}^{v_{0}}\colon\mathrm{Geod}_{T}(v_{0}\to T)\longrightarrow\mathcal{P}_{(\Delta,\iota)}(X_{c_{0}}\to X).$
Remark 3.4.3.
By (3.5), every edge lying in a geodesic from $v_{0}$ in $T$ belongs $ET^{+}$. Indeed, if $(e_{1},\ldots,e_{n})\in\mathrm{Geod}_{T}(v_{0}\to T)$ and $\mathcal{L}(e_{1},\ldots,e_{n})=(x_{1},\ldots,x_{n})$, then $e_{1}=(O_{c_{0}},x_{1})$ and $e_{i}=\big{(}(x_{1},\ldots,x_{i-1}),(x_{1},\ldots,x_{i})\big{)}$ for every $2\leq i\leq n$.
Moreover, if $e_{1}\in ET^{+}$, then for every $e_{2}\in o^{-1}(t(e_{1}))\setminus\{\bar{e}_{1}\}$ the path $[v_{0},e_{1}]\cdot e_{2}$ is a geodesic from $v_{0}$ and thus $e_{2}\in ET^{+}$.
We define the graph epimorphism $\pi\colon T\to\Gamma$ by putting $\pi(O_{c_{0}})=c_{0}$ and, provided $v=(x_{1},\ldots,x_{n})\in VT$ ($n\geq 1$) has underlying path $(a_{1},\ldots,a_{n})$ in $\Gamma$, by $\pi(v)=t(a_{n})$. The triple $(T,\pi,\mathcal{L})$ is a $\Delta$-tree that we call the standard $\Delta$-tree associated to $\iota$ and $c_{0}$.
Since the map in (3.6) is a bijection, we deduce the following.
Lemma 3.5.
Put $T=T(\Delta,\iota,c_{0})$ and consider $e\in ET^{+}$ with $\mathcal{L}(e)=x$. Then the map $\mathcal{L}$ in (3.5) restricts to the following bijection:
(3.7) $\mathcal{L}^{e}\colon\mathrm{Geod}_{T}(e\to T)\longrightarrow\mathcal{P}_{(\Delta,\iota)}(x\to X).$
If in particular $o(e)=v_{0}$, the map $\mathcal{L}$ in (3.5) restricts also to the following bijection:
(3.8) $\mathcal{L}^{\bar{e}}\colon\mathrm{Geod}_{T}(\bar{e}\to T)\longrightarrow\iota(x)\cdot\mathcal{P}_{(\Delta,\iota)}(X_{c_{0}}\setminus\{x\}\to X).$
Proof.
Write $e=(v,w)$, where $w=(x_{1},\ldots,x_{m},x)\in\mathcal{P}_{(\Delta,\iota)}(X_{c_{0}}\to X)$. Note that $\mathcal{L}([v_{0},v])=(x_{1},\ldots,x_{m})$. The bijection $\mathcal{L}^{v_{0}}$ in (3.6) restricts to a $1$-to-$1$ map
$[v_{0},v]\cdot\mathrm{Geod}_{T}(e\to T)\longrightarrow(x_{1},\ldots,x_{m})\cdot\mathcal{P}_{(\Delta,\iota)}(x\to X).$
This implies that $\mathcal{L}^{e}$ is a bijection. Moreover, we observe that
(3.9) $\mathrm{Geod}_{T}(\bar{e}\to T)=\bigsqcup_{f\in o^{-1}(v),\,f\neq e}\bar{e}\cdot\mathrm{Geod}_{T}(f\to T)$
and
(3.10) $\iota(x)\cdot\mathcal{P}_{(\Delta,\iota)}(X_{\pi(v)}\setminus\{x\}\to X)=\bigsqcup_{y\in X_{\pi(v)},\,y\neq x}\iota(x)\cdot\mathcal{P}_{(\Delta,\iota)}(y\to X).$
If $v=v_{0}$, then $o^{-1}(v)\subseteq ET^{+}$. From (3.9), (3.10) and the first part of the statement we conclude that $\mathcal{L}^{\bar{e}}$ is bijective. ∎
Remark 3.5.1.
Although it is not necessary for the discussion, for every $e=(v,w)\in ET^{+}$ one may restrict the map $\mathcal{L}$ in (3.5) to a bijection $\mathcal{L}^{\bar{e}}$ from $\mathrm{Geod}_{T}(\bar{e}\to T)$ to a suitable set of paths in $(\Delta,\iota)$. One may proceed inductively on $\ell([v_{0},v])=l\geq 0$. The case $l=0$ is done by Lemma 3.5. If $l\geq 1$, one assumes the claim true for $l-1$ and observes that $o^{-1}(v)\setminus\{e\}$ has exactly one edge which does not belong to $ET^{+}$: it is the edge whose reverse $f_{0}$ is the last edge of $[v_{0},v]$. By Lemma 3.5, for all $f\in o^{-1}(v)\setminus\{e,\bar{f}_{0}\}$, the map $\mathcal{L}^{f}$ as in (3.7) is bijective. Moreover, $\ell([v_{0},o(f_{0})])=\ell([v_{0},v])-1$ and by induction one has $\mathcal{L}^{\bar{f}_{0}}$ is bijective. Using the decomposition in (3.9), one determines the image of $\mathrm{Geod}_{T}(\bar{e}\to T)$ via $\mathcal{L}$ and thus the bijection $\mathcal{L}^{\bar{e}}$.
3.4.2. The universal group $U(\Delta,\mathbb{T})$
Let $\Delta$ be a local action diagram with a $\Delta$-tree $\mathbb{T}=(T,\pi,\mathcal{L})$. For all $g\in\mathrm{Aut}_{\pi}(T)$ and $v\in VT$, there is an induced permutation $\sigma(g,v)\colon X_{\pi(v)}\to X_{\pi(v)}$ given by
$\sigma(g,v)(x):=\Big{(}\mathcal{L}_{g\cdot v}\circ g\circ(\mathcal{L}_{v})^{-1}\Big{)}(x).$
Definition 3.6 ([17, Definition 3.8]).
The universal group associated to $\Delta$ and $\mathbb{T}$ is
$U(\Delta,\mathbb{T}):=\{g\in\mathrm{Aut}_{\pi}(T)\mid\forall\,v\in VT,\,\sigma(g,v)\in G(\pi(v))\}.$
In other words, $U(\Delta,\mathbb{T})$ collects all elements of $\mathrm{Aut}_{\pi}(T)$ acting on $o^{-1}(v)$ as a permutation in $G(\pi(v))$, for every $v\in VT$. If $\mathbb{T}$ is the standard $\Delta$-tree associated to $\iota$ and $c_{0}$ (cf. Section 3.4.1), we write $U(\Delta,\iota,c_{0})$ instead of $U(\Delta,\mathbb{T})$. If $T$ is locally finite, the group $U(\Delta,\mathbb{T})$ is a t.d.l.c. group with respect to the subspace topology from $\mathrm{Aut}(T)$. Indeed, $U(\Delta,\mathbbm{T})$ is a closed subgroup of the t.d.l.c. group $\mathrm{Aut}(T)$ (cf. [17, §6]). If additionally $\Gamma$ is finite, then $U(\Delta,\mathbbm{T})$ is a compactly generated t.d.l.c. group (cf. [17, Proposition 6.5]).
The group $U(\Delta,\mathbb{T})$ simultaneously generalises the three following notable examples.
Example 3.7.
Let $\Delta=(\Gamma,(X_{a}),(G(c)))$ be a local action diagram and $\mathbb{T}=(T,\pi,\mathcal{L})$ a $\Delta$-tree.
  • (i)
    For every $c\in V\Gamma$, let
    $G(c):=\{\sigma\in\mathrm{Sym}(X_{c})\mid\forall\,a\in o^{-1}(c),\,\sigma(X_{a})=X_{a}\}.$
    Then $U(\Delta,\mathbb{T})=\mathrm{Aut}_{\pi}(T)$.
  • (ii)
    Let $\Gamma$ be a $1$-segment with $E\Gamma=\{a,\bar{a}\}$, $c=o(a)$ and $d=t(a)$. Let $G(c)$ and $G(d)$ act transitively on $X_{c}$ and $X_{d}$, respectively.
    • (iia)
      Set $F:=G(c)$ and assume that $G(d)=C_{2}$, $|X_{c}|=k\geq 2$ and $|X_{d}|=2$. Adapting [18, Example 11] to actions on trees without edge inversion, the action $(U(\Delta,\mathbb{T}),T)$ is permutational isomorphic to the action of the Burger–Mozes universal group $U(F)$ on the barycentric subdivision $T_{k}^{\prime}$ of the $k$-regular tree $T_{k}$ (cf. [5, §3.2]). Here we consider $T_{k}^{\prime}$, and not $T_{k}$, because $U(F)$ acts vertex-transitively (and thus with edge inversions) on $T_{k}$.
    • (iib)
      Set $F_{1}:=G(c)$ and $F_{2}:=G(d)$. Following [18, Example 12], the action $U(\Delta,\mathbb{T})$-action on $T$ is permutational isomorphic to $(U(F_{1},F_{2}),T)$, where $U(F_{1},F_{2})$ is the group introduced by S. Smith in [22].
The following fact collects some key results of the work of C. Reid and S. Smith [17]. It motivates why, throughout the paper, we focus on actions of the form $(U(\Delta,\iota,c_{0}),T(\Delta,\iota,c_{0}))$ while considering (P)-closed actions on trees with associated local action diagram $\Delta$.
Theorem 3.8.
  • (i)
    ([17, Lemma 3.5, Theorem 3.12]) Let $\Delta$ be a local action diagram. For all $\Delta$-trees $\mathbb{T}=(T,\pi,\mathcal{L})$ and $\mathbb{T}^{\prime}=(T^{\prime},\pi^{\prime},\mathcal{L}^{\prime})$ and $v\in VT$, $v^{\prime}\in VT^{\prime}$ with $\pi(v)=\pi^{\prime}(v^{\prime})$, there is a graph isomorphism $\phi\colon T\to T^{\prime}$ such that $\phi(v)=v^{\prime}$, $\pi=\pi^{\prime}\circ\phi$ and $\phi U(\Delta,\mathbbm{T})\phi^{-1}=U(\Delta,\mathbb{T}^{\prime})$.
  • (ii)
    ([17, Theorem 3.9]) Let $\Delta$ be a local action diagram and $\mathbb{T}=(T,\pi,\mathcal{L})$ be a $\Delta$-tree. Then the local action diagram associated to $(U(\Delta,\mathbb{T}),T)$ is isomorphic to $\Delta$ (in the sense of [17, Definition 3.2]).
  • (iii)
    ([17, Theorem 3.10]) Let $(G,T)$ be a (P)-closed action on a tree with associated local action diagram $\Delta$. Then $(G,T)$ is permutational isomorphic to $(U(\Delta,\iota,c_{0}),T(\Delta,\iota,c_{0}))$, for every inversion $\iota$ in $\Delta$ and every $c_{0}\in V\Gamma$.
3.5. Weakly locally $\infty$-transitive actions on trees
Definition 3.9.
Let $(G,T)$ be a group action on a tree with quotient map $\pi\colon T\to\Gamma=G\backslash T$. Then $(G,T)$ is said to be weakly locally $\infty$-transitive at $v\in VT$ if, for every path $\mathfrak{p}$ in $\Gamma$ starting at $\pi(v)$, the stabiliser $G_{v}$ acts transitively on the set
(3.11) $\{\tilde{\mathfrak{p}}\in\mathrm{Geod}_{T}(v\to T)\mid\pi(\tilde{\mathfrak{p}})=\mathfrak{p}\}.$
Moreover, $(G,T)$ is said to be weakly locally $\infty$-transitive if it is weakly locally $\infty$-transitive at every $v\in VT$.
Remark 3.9.1.
  • (i)
    The action $(G,T)$ is weakly locally $\infty$-transitive at $v$ if, and only if, it is weakly locally $\infty$-transitive at every $u\in G\cdot v$.
  • (ii)
    Let $(G,T)$ be weakly locally $\infty$-transitive at $v$, and consider $\tilde{\mathfrak{p}}=(e_{1},\ldots,e_{n})\in\mathrm{Geod}_{T}(v\to T)$ with $\pi(\tilde{\mathfrak{p}})=\mathfrak{p}$. Then, for every $i<n$ the group $G_{(e_{1},\ldots,e_{i})}$ acts transitively on
    $\{\widetilde{\mathfrak{q}}=(f_{1},\ldots,f_{n})\in\mathrm{Geod}_{T}(v\to T)\mid\pi(\widetilde{\mathfrak{q}})=\mathfrak{p}\text{ and }\forall\,j\leq i,f_{j}=e_{j}\}.$
As the name suggests, weakly locally $\infty$-transitive actions generalise locally $\infty$-transitive ones as follows.
Lemma 3.10.
Let $(G,T)$ be a group action on a tree with quotient graph $\Gamma$. Then $(G,T)$ is locally $\infty$-transitive if, and only if, it is weakly locally $\infty$-transitive and $\Gamma$ is a $1$-segment.
Proof.
Let $\pi\colon T\to\Gamma=G\backslash T$ be the quotient map of $(G,T)$. Given $v\in VT$ with $\pi(v)=c$ and for every $d\geq 0$, we observe that
(3.12) $\{\tilde{\mathfrak{p}}\in\mathrm{Geod}_{T}(v\to T)\mid\ell(\tilde{\mathfrak{p}})=d\}=\bigsqcup_{\mathfrak{p}\in\mathcal{P}_{\Gamma}(c\to\Gamma),\atop\ell(\mathfrak{p})=d}\{\tilde{\mathfrak{p}}\in\mathrm{Geod}_{T}(v\to T)\mid\pi(\tilde{\mathfrak{p}})=\mathfrak{p}\}.$
Note that $\Gamma$ is a $1$-segment if, and only if, |$|\{\mathfrak{p}\in\mathcal{P}_{\Gamma}(c\to\Gamma)\,:\,\ell(\mathfrak{p})=d\}|=1$ for all $d\geq 0$ and $c\in V\Gamma$. If $(G,T)$ is weakly locally $\infty$-transitive and $\Gamma$ is a $1$-segment, then $G_{v}$ acts transitively on $\{\tilde{\mathfrak{p}}\in\mathrm{Geod}_{T}(v\to T)\mid\ell(\tilde{\mathfrak{p}})=d\}$, for all $v\in VT$ and $d\geq 0$. Conversely, if $(G,T)$ is locally $\infty$-transitive, then $(G,T)$ is edge-transitive. More precisely, since $G_{v}$ acts transitively on $o^{-1}(v)$ for every $v\in VT$, we deduce that $\Gamma$ is a $1$-segment. Indeed, if $\Gamma$ is a $1$-loop, for every $v\in VT$ the group $G_{v}$ has two orbits on $o^{-1}(v)$. Moreover, for all $v\in VT$ and $d\geq 0$, there is a unique path $\mathfrak{p}$ in $\Gamma$ starting at $\pi(v)$ of length $d$. By (3.12), we conclude that $(G,T)$ is weakly locally $\infty$-transitive. ∎
The next lemma provides a local characterisation of weakly locally $\infty$-transitive (P)-closed actions on trees. An analogous result has already been proved for Burger–Mozes universal groups [5, Lemma 3.1.1 and the lines before it].
Proposition 3.11.
Let $(G,T)$ be a group action on a tree with associated local action diagram $\Delta=(\Gamma,(X_{a}),(G(c)))$, and let $v\in VT$ with $\pi(v)=c$. If $(G,T)$ is weakly locally $\infty$-transitive at $v$, then
  • ($\diamondsuit$)
    for all $a,b\in o^{-1}(c)$ and $x\in X_{a}$, the group $G(c)_{x}$ acts transitively on $X_{b}\setminus\{x\}$.
Conversely, if $(G,T)$ is (P)-closed, condition ($\diamondsuit$) implies that $(G,T)$ is weakly locally $\infty$-transitive at $v$.
In ($\diamondsuit$), note that $X_{b}\setminus\{x\}=X_{b}$ unless $a=b$. Moreover, if $a=b$, condition ($\diamondsuit$) is equivalent to say that $G(c)$ acts $2$-transitively on $X_{a}$.
Proof.
To prove the first part of the assertion, let $a,b\in o^{-1}(c)$. By essential uniqueness of the local action diagram associated to $(G,T)$ (cf. Section 3.4), we may assume that $X_{a},X_{b}\subseteq o^{-1}(v)$. Since $(G,T)$ is weakly locally $\infty$-transitive at $v$, for every $e\in X_{a}$ the stabiliser $G_{e}$ acts transitively on
$\{\mathfrak{p}\in\mathrm{Geod}_{T}(\bar{e}\to T)\mid\pi(\mathfrak{p})=(\bar{a},b)\}=\{(\bar{e},f)\mid f\in X_{b}\setminus\{e\}\}$
and thus on $X_{b}\setminus\{e\}$.
Let now $(G,T)$ be (P)-closed and suppose that ($\diamondsuit$) holds. Without loss of generality, set $G=U(\Delta,\iota,c)$ and $T=T(\Delta,\iota,c)$. Consider two geodesics in $T$, say $[v,w_{1}]=(e_{1},\dots,e_{n})$ and $[v,w_{2}]=(f_{1},\dots,f_{n})$ for some $n\geq 1$ and $w_{1},w_{2}\in VT$, with the same image in $\Gamma$. For every $1\leq i\leq n$, since $\pi(e_{i})=\pi(f_{i})$, there exists $g_{i}\in G$ such that $f_{i}=g_{i}\cdot e_{i}$. Note that $o(g_{i}\cdot e_{i+1})=g_{i}\cdot t(e_{i})=t(f_{i})=o(f_{i+1})$ and $\pi(g_{i}\cdot e_{i+1})=\pi(e_{i+1})=\pi(f_{i+1})$. Hence, both $\mathcal{L}(g_{i}\cdot e_{i+1})$ and $\mathcal{L}(f_{i+1})$ belong to $X_{\pi(f_{i+1})}$. By ($\diamondsuit$) applied to the vertex $c_{i}=o(\pi(f_{i+1}))$, for every $i<n$ there is $h_{i}\in G_{\bar{f}_{i}}$ such that $f_{i+1}=h_{i}g_{i}\cdot e_{i+1}$. For every $i<n$, set $k_{i}:=h_{i}g_{i}$ and observe that
(3.13) $(k_{i}\cdot e_{i},k_{i}\cdot e_{i+1})=(f_{i},f_{i+1}).$
In particular, $k_{1}\in G_{v}$ and $k_{i}\cdot e_{i+1}=f_{i+1}=k_{i+1}\cdot e_{i+1}$ for every $i<n-1$. For every $i<n-1$ write $u_{i+1}:=k_{i}^{-1}k_{i+1}\in G_{e_{i+1}}$, and let $u_{i+1}^{-}\in G_{T_{\geq\bar{e}_{i+1}}}$ and $u_{i+1}^{+}\in G_{T_{\geq e_{i+1}}}$ be such that $u_{i+1}=u_{i+1}^{-}u_{i+1}^{+}$ (recall that $(G,T)$ is (P)-closed). Set $\tilde{k}_{1}:=k_{1}$ and, for every $12\leq i\leq n-$ 1, define inductively $\tilde{k}_{i}:=\tilde{k}_{i-1}u_{i}^{-}$. For each $12\leq i\leq n-$ 1, note that $k_{i}=\tilde{k}_{i}u_{i}^{+}$ and, since $u_{i}^{+}$ fixes both $e_{i}$ and $e_{i+1}$, from (3.13) we have
(3.14) $(\tilde{k}_{i}\cdot e_{i},\tilde{k}_{i}\cdot e_{i+1})=(k_{i}\cdot e_{i},k_{i}\cdot e_{i+1})=(f_{i},f_{i+1}).$
Moreover, we claim that $\tilde{k}_{i}$ fixes $v$ for every $11\leq i\leq n-$ 1. If $i=1$, this is clear. For $i\geq 2$, assuming inductively that $\tilde{k}_{i-1}$ fixes $v$, the fact that $u_{i}^{-}$ fixes $T_{\geq\bar{e}_{i}}$ pointwise implies that $\tilde{k}_{i}=\tilde{k}_{i-1}u_{i}^{-}$ fixes $v$. In particular, $\tilde{k}_{n-1}\in G_{v}$ and (3.14) yields
$\tilde{k}_{n-1}\cdot w_{1}=\tilde{k}_{n-1}\cdot t(e_{n})=t(\tilde{k}_{n-1}\cdot e_{n})=t(f_{n})=w_{2}.$
Being $T$ a tree, we conclude that $[v,w_{2}]=[v,\tilde{k}_{n-1}\cdot w_{1}]=\tilde{k}_{n}\cdot[v,w_{1}]$. ∎
Proposition 3.11 gives a recipe for constructing (P)-closed actions that are weakly locally $\infty$-transitive. We collect some explicit examples below.
Example 3.12.
  • (i)
    Let $(T,\pi)$ be the universal cover of a connected edge-weighted graph $(\Gamma,\omega)$ as in Example 1.1. According to Example 3.7(i), one checks that ($\diamondsuit$) is satisfied for every $c\in V\Gamma$. Hence, by Proposition 3.11, $(\mathrm{Aut}_{\pi}(T),T)$ is weakly locally $\infty$-transitive. This gives an alternative proof of [7, Theorem 3.1 and the comment thereafter].
  • (ii)
    Let $U(F)$ be the Burger–Mozes universal group associated to a transitive group $F\leq\mathrm{Sym}(\{1,\ldots,k\})$, $k\geq 2$, acting on the barycentric subdivision $T_{k}^{\prime}$ of the $k$-regular tree. According to Example 3.7(iia), condition ($\diamondsuit$) of Proposition 3.11 is satisfied for every $c\in V\Gamma$ if, and only if, $F$ is $2$-transitive. Moreover, $U(F)\backslash T_{k}^{\prime}$ is a $1$-segment. Hence, $(U(F),T_{k}^{\prime})$ is weakly locally $\infty$-transitive if, and only if, it is locally $\infty$-transitive (cf. Lemma 3.10). By Proposition 3.11, we deduce that $(U(F),T_{k}^{\prime})$ is locally $\infty$-transitive if, and only if, $F$ is $2$-transitive. This gives an alternative proof of what was observed in the first lines of [5, §3].
  • (iii)
    Let $U(F_{1},F_{2})$ be the Smith’s group associated to two transitive groups $F_{1}\leq\mathrm{Sym}(\{1,\ldots,k_{1}\})$ and $F_{2}\leq\mathrm{Sym}(\{1,\ldots,k_{2}\})$, $k_{1},k_{2}\geq 2$. According to Example 3.7(iib), condition ()($\diamondsuit$)( ♢ ) is satisfied for every $c\in V\Gamma$ if, and only if, both $F_{1}$ and $F_{2}$ are $2$-transitive. Moreover, the $U(F_{1},F_{2})$-action on the $(k_{1},k_{2})$-biregular tree $T_{k_{1},k_{2}}$ has quotient graph a $1$-segment. Hence, as in (ii), $(U(F_{1},F_{2}),T_{k_{1},k_{2}})$ is weakly locally $\infty$-transitive if, and only if, it is locally $\infty$-transitive. By Proposition 3.11 we conclude that $(U(F_{1},F_{2}),T_{k_{1},k_{2}})$ is locally $\infty$-transitive if, and only if, both $F_{1}$ and $F_{2}$ are $2$-transitive.
4. Cosets and double-cosets for groups acting on trees
Given a group acting on a tree, there is a standard geometric characterisation of coset and double-coset spaces (and of double-coset sizes) with respect to vertex and edge stabilisers. We expose it in Section 4.1. In view of Section 5, we rephrase this characterisation in local terms in case that the action is weakly locally $\infty$-transitive (cf. Section 4.2) or (P)-closed (cf. Section 4.3). For weakly locally $\infty$-transitive actions, we can conveniently enumerate double-cosets and compute their sizes in terms of paths in the quotient graph and a suitable weight on them (cf. Propositions 4.3 and 4.6). In the (P)-closed case, we bypass the problem of enumerating double-cosets in local terms and focus on the associated coset spaces. Indeed, the latter ones can be conveniently described in terms of paths in the local action diagram (cf. Proposition 4.7). This also allows us to express the relevant double-coset sizes in local terms (cf. Proposition 4.11). We will see in Section 5.1 that counting cosets instead of double-cosets has only little impact on the double-coset zeta functions.
4.1. Cosets and geodesics
Since a tree is uniquely geodesic and by the orbit-stabiliser theorem, we observe what follows.
Fact 4.1.
Let $(G,T)$ be a group action on a tree and let $v\in VT$, $e\in ET$. For every $t\in VT$, there are $G$-equivariant bijections $\varphi_{v,t}\colon G/G_{t}\to\mathrm{Geod}_{T}(v\to G\cdot t)$ and $\varphi_{e,t}\colon G/G_{t}\to\mathrm{Geod}_{T}(\{e,\bar{e}\}\to G\cdot t)$ defined as follows:
$\varphi_{v,t}(gG_{t}):=[v,g\cdot t]\quad\text{and}\quad\varphi_{e,t}(gG_{t}):=\left\{\begin{array}[]{cc}[e,g\cdot t],&\text{if }{g\cdot t\in T_{\geq e}};\\ {[\bar{e},g\cdot t]},&\text{if }g\cdot t\in T_{\geq\bar{e}}.\end{array}\right.$
Similarly, for every $t\in ET$ there are $G$-equivariant bijections $\varphi_{v,t}\colon G/G_{t}\to\mathrm{Geod}_{T}(v\to G\cdot\{t,\bar{t}\})$ and $\varphi_{e,t}\colon G/G_{t}\to\mathrm{Geod}_{T}(\{e,\bar{e}\}\to G\cdot\{t,\bar{t}\})$ defined as follows:
$\begin{split}\varphi_{v,t}(gG_{t})&:=\left\{\begin{array}[]{cc}[v,g\cdot t],&\text{if }[v,g\cdot t]\text{ exists in }T;\\ {[v,g\cdot\bar{t}]},&\text{if }[v,g\cdot\bar{t}]\text{ exists in }T;\end{array}\right.\quad\text{and}\\ \varphi_{e,t}(gG_{t})&:=\left\{\begin{array}[]{cc}{[e,g\cdot t]},&\text{if }g\cdot t\in T_{\geq e};\\ {[e,g\cdot\bar{t}]},&\text{if }g\cdot\bar{t}\in T_{\geq e};\\ {[\bar{e},g\cdot t]},&\text{if }g\cdot t\in T_{\geq\bar{e}};\\ {[\bar{e},g\cdot\bar{t}]},&\text{if }g\cdot\bar{t}\in T_{\geq\bar{e}.}\end{array}\right.\end{split}$
Lemma 4.2.
Let $(G,T)$ be a group action on a tree. According to Fact 4.1, for all $t_{1},t_{2}\in T$ we have
$|G_{t_{1}}gG_{t_{2}}/G_{t_{2}}|=|G_{t_{1}}\cdot\varphi_{t_{1},t_{2}}(gG_{t_{2}})|.$
Proof.
For every $g\in G$, we observe that
$|G_{t_{1}}gG_{t_{2}}/G_{t_{2}}|=|G_{t_{1}}:G_{t_{1}}\cap gG_{t_{2}}g^{-1}|.$
Since $gG_{t_{2}}g^{-1}=G_{g\cdot t_{2}}$ and $T$ is uniquely geodesic, the group $G_{t_{1}}\cap gG_{t_{2}}g^{-1}$ is the pointwise stabiliser of the geodesic $\varphi_{t_{1},t_{2}}(gG_{t_{2}})$. Then the orbit-stabiliser theorem yields the claim. ∎
4.2. The case of weakly locally $\infty$-transitive actions on trees
Let $(G,T)$ be a weakly locally $\infty$-transitive group action on a tree. Denote by $\pi\colon T\to\Gamma=G\backslash T$ the quotient map of $(G,T)$, and let $\omega$ be the standard edge weight on $\Gamma$. Given $u_{1},u_{2}\in\Gamma$ and $t_{1}\in\pi^{-1}(u_{1})$, let $\mathcal{P}^{\mathrm{lift}}_{\Gamma,t_{1}}(u_{1}\to u_{2})$ be the set of all paths $\mathfrak{p}\in\mathcal{P}_{\Gamma}(u_{1}\to u_{2})$ that can be lifted via $\pi$ to a geodesic in $T$ from $t_{1}$.
Proposition 4.3.
In the hypotheses before, let $v\in VT$ with $\pi(v)=c$ and $e\in ET$ with $\pi(e)=a$. According to Fact 4.1, we have the following bijections for every $t\in VT$ with $\pi(t)=u$:
$\begin{array}[]{rcll}\Psi_{v,t}\colon G_{v}\backslash G/G_{t}&\longrightarrow&\mathcal{P}^{\mathrm{lift}}_{\Gamma,v}(c\to U),&\Psi_{v,t}(G_{v}gG_{t})=\pi(\varphi_{v,t}(gG_{t}));\\ \Psi_{e,t}\colon G_{e}\backslash G/G_{t}&\longrightarrow&\mathcal{P}^{\mathrm{lift}}_{\Gamma,e}(A\to U),&\Psi_{e,t}(G_{e}gG_{t})=\pi(\varphi_{e,t}(gG_{t})).\end{array}$
Here the sets $A$ and $U$ are according to Notation 2.1.
Proof.
We only prove that $\Psi_{v,t}$ is bijective for $t\in VT$, as for the remaining cases one may argue analogously. Let $\pi_{v,t}\colon\mathrm{Geod}_{T}(v\to G\cdot t)\longrightarrow\mathcal{P}^{\mathrm{lift}}_{\Gamma,v}(c\to u)$ be the map defined as
$\pi_{v,t}([v,g\cdot t]):=\pi([v,g\cdot t]).$
Clearly, $\pi_{v,t}$ is $G$-equivariant and surjective. Moreover, since the $G$-action on $T$ is weakly locally $\infty$-transitive, for every $g\in G$ the $\pi_{v,t}$-fibre of $\pi([v,g\cdot t])$ is $G_{v}\cdot[v,g\cdot t].$ Thus $\pi_{v,t}$ induces a $1$–to–$1$ map
$\tilde{\Psi}_{v,t}\colon G_{v}\backslash\mathrm{Geod}_{T}(v\to G\cdot t)\longrightarrow\mathcal{P}^{\mathrm{lift}}_{\Gamma,v}(c\to u).$
Composing the bijection $G_{v}\backslash G/G_{t}\to G_{v}\backslash\mathrm{Geod}_{T}(v\to G\cdot t)$ induced by $\varphi_{v,t}$ with $\tilde{\Psi}_{v,t}$, we obtain $\Psi_{v,t}$. ∎
In view of Proposition 4.6, we introduce a weight on the paths of $G\backslash T$ which extends the standard edge weight $\omega$ and such that $|G_{t_{1}}gG_{t_{2}}/G_{t_{2}}|$ coincides with the weight of $\Psi_{t_{1},t_{2}}(G_{t_{1}}gG_{t_{2}})$, for all $t_{1},t_{2}\in T$. Such a weight can be defined even in the following more general framework.
Definition 4.4.
Let $\Gamma$ be a graph with a function $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$ (called edge weight). Define two functions $N_{\mathrm{edg}}=N_{\mathrm{edg}}^{\omega},N_{\mathrm{vert}}=N_{\mathrm{vert}}^{\omega}\colon\mathcal{P}_{\Gamma}\to\mathbb{Z}_{\geq 0}$ as follows. For every path $\mathfrak{p}$ in $\Gamma$, let
(4.1) $N_{\mathrm{edg}}(\mathfrak{p}):=\left\{\begin{array}[]{cl}\ell(\mathfrak{p}),&\text{if }\ell(\mathfrak{p})\leq 1\\ \displaystyle{\prod_{i=1}^{n-1}\Big{(}\omega(a_{i+1})-\mathbbm{1}_{\{\bar{a}_{i}\}}(a_{i+1})\Big{)},}&\text{if }\mathfrak{p}=(a_{1},\ldots,a_{n}),\,n\geq 2\end{array}\right.$
and
(4.2) $N_{\mathrm{vert}}(\mathfrak{p}):=\left\{\begin{array}[]{cl}1,&\text{if }\ell(\mathfrak{p})=0\\ \omega(a_{1})\cdot N_{\mathrm{edg}}(\mathfrak{p}),&\text{if }\mathfrak{p}=(a_{1},\ldots,a_{n}),\,n\geq 1.\end{array}\right.$
Notation 4.5.
For $(a_{1},\ldots,a_{n})\in\mathcal{P}_{\Gamma}$, we write $N_{\mathrm{edg}}(a_{1},\ldots,a_{n})$ and $N_{\mathrm{vert}}(a_{1},\ldots,a_{n})$ in place of $N_{\mathrm{edg}}((a_{1},\ldots,a_{n}))$ and $N_{\mathrm{vert}}((a_{1},\ldots,a_{n}))$, respectively.
Remark 4.5.1.
Let $\Gamma$ be the quotient graph of a group action on a tree $(G,T)$, and denote by $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$ its standard edge weight. For every path $\mathfrak{p}=(a_{1},\ldots,a_{n})$ in $\Gamma$ of positive length and for every $e\in ET$ with $\pi(e)=a_{1}$, one checks that
$N_{\mathrm{edg}}^{\omega}(\mathfrak{p})=|\{\widetilde{\mathfrak{p}}\in\mathrm{Geod}_{T}(e\to T)\,:\,\pi(\widetilde{\mathfrak{p}})=\mathfrak{p}\}|.$
Similarly, for every path $\mathfrak{p}$ in $\Gamma$ starting at $c\in V\Gamma$ and for every $v\in VT$ with $\pi(v)=c$, we have
$N_{\mathrm{vert}}^{\omega}(\mathfrak{p})=|\{\widetilde{\mathfrak{p}}\in\mathrm{Geod}_{T}(v\to T)\,:\,\pi(\widetilde{\mathfrak{p}})=\mathfrak{p}\}|.$
These observations apply if in particular $\Gamma$ is the underlying graph of a local action diagram $\Delta=(\Gamma,(X_{a}),(G(c)))$, and if $T=T(\Delta,\iota,c_{0})$ is a standard $\Delta$-tree (cf. Section 3.4.1). In this case, for path $\mathfrak{p}=(a_{1},\ldots,a_{n})$ in $\Gamma$ of positive length and every $e\in ET^{+}$ with $\pi(e)=a_{1}$, we have
(4.3) $N_{\mathrm{edg}}^{\omega}(\mathfrak{p})=|\{\xi\in\mathcal{P}_{(\Delta,\iota)}(\mathcal{L}(e)\to X)\,:\,\mathfrak{p}_{\xi}=\mathfrak{p}\}|,$
(cf. Lemma 3.5). Thus a reduced path $\xi$ in $(\Delta,\iota)$ can be lifted to a geodesic in $T$ if, and only if, $N_{\mathrm{edg}}^{\omega}(\mathfrak{p}_{\xi})\geq 1$.
Remark 4.5.2.
Under the hypotheses of Proposition 4.3, suppose that $\omega(a)\geq 2$ for every $a\in E\Gamma$. Then every path $\mathfrak{p}$ can be lifted to a geodesic in $T$ via $\pi$ (cf. Remark 4.5.1). In particular, the bijections in Proposition 4.3 are onto $\mathcal{P}_{\Gamma}(c\to U)$ and $\mathcal{P}_{\Gamma}(A\to U)$, respectively.
Proposition 4.6.
Let $(G,T)$ be a weakly locally $\infty$-transitive group action on a tree with quotient graph $\Gamma$. Assume that the standard edge weight $\omega$ on $\Gamma$ takes finite values. Then, for all $g\in G$ and $t_{1},t_{2}\in T$, we have
$|G_{t_{1}}gG_{t_{2}}/G_{t_{2}}|=|G_{t_{1}}:G_{\varphi_{t_{1},t_{2}}(gG_{t_{2}})}|=\left\{\begin{array}[]{ll}N_{\mathrm{vert}}^{\omega}(\Psi_{t_{1},t_{2}}(G_{t_{1}}gG_{t_{2}})),&\text{if }t_{1}\in VT;\\ N_{\mathrm{edg}}^{\omega}(\Psi_{t_{1},t_{2}}(G_{t_{1}}gG_{t_{2}})),&\text{if }t_{1}\in ET.\end{array}\right.$
Here the map $\Psi_{t_{1},t_{2}}$ is as in Proposition 4.3.
Proof.
We may assume that $\varphi_{t_{1},t_{2}}(gG_{t_{2}})=[t_{1},g\cdot t_{2}]$. Indeed, in the other cases the argument is analogous. Let $\mathfrak{p}:=\Psi_{t_{1},t_{2}}(G_{t_{1}}gG_{t_{2}})=\pi([t_{1},g\cdot t_{2}])$, where $\pi$ denotes the quotient map of $(G,T)$ (extended entrywise to all paths). By Lemma 4.2 and by weak local $\infty$-transitivity, we deduce that
$|G_{t_{1}}gG_{t_{2}}/G_{t_{2}}|=|G_{t_{1}}\cdot[t_{1},g\cdot t_{2}]|=|\{\widetilde{\mathfrak{p}}\in\mathrm{Geod}_{T}(t_{1}\to T)\,:\,\pi(\widetilde{\mathfrak{p}})=\mathfrak{p}\}|.$
Now Remark 4.5.1 applies. ∎
4.3. The case of (P)-closed actions on trees
Let $\Delta$ be a local action diagram and $(T=T(\Delta,\iota,c_{0}),\pi,\mathcal{L})$ be the standard $\Delta$-tree associated to an inversion map $\iota$ in $\Delta$ and a chosen $c_{0}\in V\Gamma$ (cf. Section 3.4.1). Let $G\leq U(\Delta,\iota,c_{0})$ be a subgroup acting on $T$ with local action diagram $\Delta$. For the definition of $U(\Delta,\iota,c_{0})$, see Section 3.4.2.
The following proposition rephrases Fact 4.1 in the language of local action diagrams.
Proposition 4.7.
Let $G\leq U(\Delta,\iota,c_{0})$ and $T=T(\Delta,\iota,c_{0})$ as before, and denote by $v_{0}$ the root of $T$. Consider also $e\in o^{-1}(v_{0})$ with $\mathcal{L}(e)=x$ and $t\in T$ with $\pi(t)=u$. Let $\varphi_{u,t},\varphi_{e,t}$ be the maps introduced in Fact 4.1, and denote by $\mathcal{L}$ the map defined in (3.5). Then the following two maps are bijective:
$\begin{array}[]{ccl}\mathcal{L}\circ\varphi_{v_{0},t}\colon G/G_{t}&\longrightarrow&\mathcal{P}_{(\Delta,\iota)}(X_{c_{0}}\to X_{U});\\ \mathcal{L}\circ\varphi_{e,t}\colon G/G_{t}&\longrightarrow&\mathcal{P}_{(\Delta,\iota)}(x\to X_{U})\sqcup\iota(x)\cdot\mathcal{P}_{(\Delta,\iota)}(X_{c_{0}}\setminus\{x\}\to X_{U}).\end{array}$
For $X_{U}$ see Notation 3.2.
Proof.
It is a direct consequence of Fact 4.1 and Lemma 3.5. ∎
Following a similar strategy to the one in Section 4.2, we define a weight on paths in $\Delta$ as follows.
Definition 4.8.
Let $\Delta=(\Gamma,(X_{a}),(G(c)))$ be a local action diagram and recall that $X=\bigsqcup_{a\in E\Gamma}X_{a}$. The standard weight on $\Delta$ is the function $\mathcal{W}\colon X\times X\longrightarrow\mathbb{Z}_{\geq 0}\cup\{\infty\}$ defined, for all $x\in X_{a}$ and $y\in X_{b}$ with $a,b\in E\Gamma$, as follows:
$\mathcal{W}(x,y):=\left\{\begin{array}[]{cl}|{G(t(a))}_{\iota(x)}\cdot y|,&\text{if }t(a)=o(b);\\ 0,&\text{otherwise.}\end{array}\right.$
Define also $\mathcal{W}_{\mathrm{rev}}\colon X\times X\longrightarrow\mathbb{Z}_{\geq 0}\cup\{\infty\}$ as follows, for all $x\in X_{a}$, $y\in X_{b}$ with $a,b\in E\Gamma$:
$\mathcal{W}_{\mathrm{rev}}(x,y):=\left\{\begin{array}[]{cl}|G(o(a))_{x}\cdot y|,&\text{if }o(a)=o(b);\\ 0,&\text{otherwise.}\end{array}\right.$
Moreover, for every sequence $\xi=(x_{1},\ldots,x_{n})$ of elements of $X$ of length $n\geq 0$, define
(4.4) $\mathcal{W}(\xi):=\left\{\begin{array}[]{cl}1,&\text{if }n\leq 1;\\ \displaystyle{\prod_{i=1}^{n-1}\mathcal{W}(x_{i},x_{i+1}),}&\text{if }n\geq 2.\end{array}\right.$
Remark 4.8.1.
For $x,y\in X$ note that $\mathcal{W}(x,y)\neq 0$ if, and only if, $(x,y)$ is a path in $\Delta$. More generally, given a sequence $\xi$ of elements of $X$ we have $\mathcal{W}(\xi)\neq 0$ if, and only if, $\xi$ is a path in $\Delta$.
Notation 4.9.
For a sequence $\xi=(x_{1},\ldots,x_{n})$, we write $\mathcal{W}(x_{1},\ldots,x_{n})$ in place of $\mathcal{W}((x_{1},\ldots,x_{n}))$.
Remark 4.9.1.
Assume the hypotheses of the section, set $G=U(\Delta,\iota,c_{0})$ and denote by $\omega$ and $\mathcal{W}$ be the standard edge weights on $\Gamma$ and $\Delta$, respectively. Hence, for every $e\in ET$ we have
$|G_{o(e)}\cdot e|=\omega(\pi(e)).$
Moreover, let $e,f\in ET$ with $t(e)=v=o(f)$. If $e\in ET^{+}$, then $\mathcal{L}(\bar{e})=\iota(\mathcal{L}(e))$ (cf. Remark 3.4.1) and
$|G_{e}\cdot f|=|G_{\bar{e}}\cdot f|=|G(\pi(v))_{\iota(\mathcal{L}(e))}\cdot\mathcal{L}(f)|=\mathcal{W}(\mathcal{L}(e),\mathcal{L}(f)).$
Moreover, if $e\in ET\setminus ET^{+}$ we have
$|G_{e}\cdot f|=|G_{\bar{e}}\cdot f|=|G(\pi(v))_{\mathcal{L}(\bar{e})}\cdot f|=\mathcal{W}_{\mathrm{rev}}(\mathcal{L}(\bar{e}),\mathcal{L}(f)).$
Example 4.10.
Let $\Delta=(\Gamma,(X_{a}),(G(c)))$ be a local action diagram and consider $a,b\in E\Gamma$ with $t(a)=c=o(b)$. Assume that, for every $x\in X_{a}$, the group ${G(c)}_{\iota(x)}$ acts transitively on $X_{b}\setminus\{\iota(x)\}$. Then, for all $x\in X_{a}$ and $y\in X_{b}\setminus\{\iota(x)\}$ we have
$\mathcal{W}(x,y)=|G(c)_{\iota(x)}\cdot y|=|X_{b}\setminus\{\iota(x)\}|=w(b)-\mathbbm{1}_{\{\bar{a}\}}(b).$
The transitivity condition before leads back to Proposition 3.11. For explicit examples satisfying it, see Example 3.12.
Proposition 4.11.
Let $G=U(\Delta,\iota,c_{0})$. Consider a geodesic $\mathfrak{p}=(e_{1},\ldots,e_{n})$ in $T=T(\Delta,\iota,c_{0})$ with $n\geq 1$, $e_{1},\ldots,e_{n}\in ET^{+}$ and $\mathcal{L}(\mathfrak{p})=(x_{1},\ldots,x_{n})$. Then,
$|G_{o(e_{1})}:G_{\mathfrak{p}}|=\omega(\pi(e_{1}))\cdot\mathcal{W}(\xi)\quad\text{and}\quad|G_{e_{1}}:G_{\mathfrak{p}}|=\mathcal{W}(\xi).$
Proof.
By the orbit stabiliser theorem,
(4.5) $\begin{split}|G_{o(e_{1})}:G_{\mathfrak{p}}|&=|G_{o(e_{1})}:G_{e_{1}}|\cdot\prod_{i=1}^{n-1}|G_{(e_{1},\dots,e_{i})}:G_{(e_{1},\dots,e_{i+1})}|\\ &=|G_{o(e_{1})}\cdot e_{1}|\cdot\prod_{i=1}^{n-1}|G_{(e_{1},\dots,e_{i})}\cdot e_{i+1}|.\end{split}$
Similarly,
(4.6) $|G_{e_{1}}:G_{\mathfrak{p}}|=\prod_{i=1}^{n-1}|G_{(e_{1},\dots,e_{i})}\cdot e_{i+1}|.$
By Proposition 3.1, for every $11\leq i\leq n-$ 1 we have
(4.7) $G_{(e_{1},\dots,e_{i})}\cdot e_{i+1}=G_{e_{i}}\cdot e_{i+1}.$
Combining Remark 4.9.1, (4.5), (4.6) and (4.7), we conclude the claim. ∎
Corollary 4.12.
Let $G=U(\Delta,\iota,c_{0})$ and denote by $v_{0}$ be the root of $T=T(\Delta,\iota,c_{0})$. Let $e\in ET\setminus ET^{+}$ with $t(e)=v_{0}$, and consider $t\in T$ such that $[v_{0},t]$ is defined. Set $\mathfrak{p}_{t}:=e\cdot[v_{0},t]=(e_{1}=e,e_{2},\ldots,e_{n})$, $\mathcal{L}(\bar{e})=x_{1}$ and $\mathcal{L}(\mathfrak{p}_{t})=(\iota(x_{1}),x_{2},\ldots,x_{n})$. Then
$|G_{e}:G_{\mathfrak{p}_{t}}|=\left\{\begin{array}[]{cl}1,&\text{if }n=1;\\ \mathcal{W}_{\mathrm{rev}}(x_{1},x_{2})\cdot\mathcal{W}(x_{2},\ldots,x_{n}),&\text{otherwise}.\end{array}\right.$
Proof.
First, note that
$|G_{e}:G_{\mathfrak{p}_{t}}|=|G_{e}:G_{(e_{1},e_{2})}|\cdot|G_{(e_{1},e_{2})}:G_{\mathfrak{p}_{t}}|.$
By Remark 4.9.1, we deduce that
$|G_{e}:G_{(e_{1},e_{2})}|=|G_{e_{1}}\cdot e_{2}|=\mathcal{W}_{\mathrm{rev}}(x_{1},x_{2}).$
From Remark 3.4.3 observe that $e_{2},\ldots,e_{n}\in ET^{+}$. Hence, Proposition 3.1 and Remark 4.9.1 yield
$\begin{split}|G_{(e_{1},e_{2})}:G_{\mathfrak{p}_{t}}|&=\prod_{i=2}^{n-1}|G_{(e_{1},\ldots,e_{i})}:G_{(e_{1},\ldots,e_{i+1})}|=\prod_{i=2}^{n-1}|G_{e_{i}}\cdot e_{i+1}|\\ &=\prod_{i=2}^{n-1}\mathcal{W}(x_{i},x_{i+1})=\mathcal{W}(x_{2},\ldots,x_{n}).\qed\end{split}$
Corollary 4.13.
Let $G=U(\Delta,\iota,c_{0})$, $T=T(\Delta,\iota,c_{0})$ and assume that $|X_{a}|\geq 2$ for every $a\in E\Gamma$. Then, for each geodesic $\mathfrak{p}=(e_{1},\ldots,e_{n})$ in $T$ of length $n\geq 1$, there is $g\in G_{o(e_{1})}$ such that $g\cdot e_{i}\in ET^{+}$ for every $i\leq n$ and
$|G_{e_{1}}:G_{\mathfrak{p}}|=\mathcal{W}(\mathcal{L}(g\cdot\mathfrak{p})).$
Proof.
If $e_{1}\in ET^{+}$, by Remark 3.4.3 we may take $g=1$ and Proposition 4.11 applies. Assume now that $e_{1}\in ET\setminus ET^{+}$ and let $(f_{1},\ldots,f_{r}=\bar{e}_{1})$ be the geodesic from the root $v_{0}$ of $T$ to $\bar{e}_{1}$. Set $\mathcal{L}(f_{i})=x_{i}$ for every $1\leq i\leq r$. In particular, $\mathcal{L}(\bar{e}_{1})=x_{r}$ and then $\mathcal{L}(e_{1})=\iota(x_{r})$ (cf. Remark 3.4.1). Since $|X_{\pi(e_{1})}|\geq 2$, there is $y\in X_{\pi(e_{1})}\setminus\{\iota(x_{r})\}$ such that $(x_{1},\ldots,x_{r},y)$ is a reduced path in $(\Delta,\iota)$. Then there is $f\in ET^{+}$ such that $o(f)=o(e_{1})$ and $t(f)$ corresponds to $(x_{1},\ldots,x_{r},y)$. Since $\mathcal{L}(f)=y,\mathcal{L}(e_{1})\in X_{\pi(e_{1})}$, there is $g\in G_{o(e_{1})}$ such that $f=g\cdot e_{1}$. By Remark 3.4.3, every edge of $g\cdot\mathfrak{p}=(g\cdot e_{1},\ldots,g\cdot e_{n})$ belongs to $ET^{+}$. Moreover,
$|G_{e_{1}}\cdot\mathfrak{p}|=|gG_{e_{1}}\cdot\mathfrak{p}|=|G_{g\cdot e_{1}}g\cdot\mathfrak{p}|$
and Proposition 4.11 applies. ∎
5. Double-coset zeta functions for groups acting on trees
This section deals with the convergence and explicit formulae for the relevant double-coset zeta functions for groups acting on trees. We introduce a family of properties (labelled with positive integers) on group actions on trees (cf. Section 5.2). In Proposition 5.5, we exploit that one of these properties is satisfied to deduce that the group has polynomial double-coset growth with respect to vertex or edge stabilisers. The latter result can be refined to a characterisation in case that the action is weakly locally $\infty$-transitive (cf. Theorem 5.6) or (P)-closed (cf. Theorem 5.7). In these two cases, we also provide explicit formulae for the relevant double-coset zeta functions in terms of the local data introduced in Section 4 (cf. Theorems 5.12 and 5.19).
5.1. From double-cosets to cosets
Let $G$ be a group and $H,K\leq G$ be subgroups such that $|HgK/K|<\infty$ for every $g\in G$. For each $n\geq 1$, consider $a_{n}(G,H,K)$ as in (1.1) and define
(5.1) $b_{n}(G,H,K):=|\{gK\in G/K\,:\,|HgK/K|=n\}|.$
We claim that $b_{n}(G,H,K)<\infty$ if, and only if, $a_{n}(G,H,K)<\infty$. Moreover, if $b_{n}(G,H,K)<\infty$ then
(5.2) $b_{n}(G,H,K)=n\cdot a_{n}(G,H,K).$
To see this, consider the map $\varphi\colon gK\in G/K\longmapsto HgK\in H\backslash G/K$. For every $g\in G$ we have $\varphi^{-1}(HgK)=\{hgK\in G/K\mid h\in H\}$ and $|\varphi^{-1}(HgK)|=|H:H\cap gKg^{-1}|=|HgK/K|$.
In particular, if $(G,H,K)$ has polynomial double-coset growth then
(5.3) $\zeta_{G,H,K}(s)=\sum_{n=1}^{\infty}b_{n}(G,H,K)\cdot n^{-s-1}=\sum_{gK\in G/K}|HgK/K|^{-s-1}.$
5.2. The property ($\ast_{k}$)
Let $k\geq 1$. A group action on a tree $(G,T)$ has property $(\ast_{k})$ if, for every geodesic $(e_{1},\ldots,e_{l+k})$ in $T$ with $l\geq 1$,
(5.4) $|G_{(e_{1},\ldots,e_{l})}\cdot(e_{l+1},\ldots,e_{l+k})|\geq 2.$
One checks that property ($\ast_{k}$) implies property ($\ast_{k+1}$), for every $k\geq 1$.
Remark 5.0.1.
Let $T$ be a tree and $G\leq\mathrm{Aut}(T)$ be a subgroup with the subspace topology induced by $\mathrm{Aut}(T)$. If $(G,T)$ has property ($\ast_{k}$) for some $k\geq 1$, then $G$ is non-discrete.
In detail, assume that $(G,T)$ has property $(\ast_{k})$. If we prove that all vertex-stabilisers in $G$ are infinite, then [2, Lemma 2.1] yields the claim. Let $v\in VT$. Since $T$ has no leaves, there is a ray $(e_{i})_{i\in\mathbb{Z}_{\geq 1}}$ in $T$ with $o(e_{1})=v$. For every $h\in\mathbb{Z}_{\geq 1}$, set $\mathfrak{p}_{h}=(e_{i})_{1\leq i\leq hk}$ and note that $|G_{\mathfrak{p}_{h}}\cdot(e_{hk+1},\ldots,e_{hk+k})|=|G_{\mathfrak{p}_{h}}:G_{\mathfrak{p}_{h+1}}|\geq 2$. Therefore,
$|G_{v}:G_{\mathfrak{p}_{h}}|=|G_{v}:G_{\mathfrak{p}_{1}}|\cdot\prod_{i=1}^{h-1}|G_{\mathfrak{p}_{i}}:G_{\mathfrak{p}_{i+1}}|\geq 2^{h-1}$
for every $h\geq 1$. Thus, $G_{v}$ is infinite.
Proposition 5.1.
Let $(G,T)$ be a weakly locally $\infty$-transitive action on a locally finite tree. Denote by $\omega$ the standard edge weight on $\Gamma=G\backslash T$, consider $N_{\mathrm{edg}}=N_{\mathrm{edg}}^{\omega}$ as in Definition 4.4, and let $k\geq 1$. Then $(G,T)$ has property ($\ast_{k}$) if, and only if, $N_{\mathrm{edg}}(\rho)\geq 2$ for every path $\rho$ in $\Gamma$ of length $k+1$ which can be lifted to a geodesic in $T$.
Proof.
Let $k\geq 1$ and $\mathfrak{p}=(e_{1},\ldots,e_{l+k})$ be a geodesic in $T$ with $l\geq 1$. Denote by $\pi\colon T\to\Gamma$ the quotient map of $(G,T)$ and set $\pi(e_{i})=a_{i}$ for every $1\leq i\leq l+k$. By Remark 3.9.1(ii), $G_{(e_{1},\ldots,e_{l})}$ acts transitively on $\{\mathfrak{q}=(f_{1},\ldots,f_{l+k})\in\mathrm{Geod}_{T}(e_{1}\to T)\,:\,\pi(\mathfrak{q})=\pi(\mathfrak{p})\text{ and }\forall\,i\leq l,\,f_{i}=e_{i}\}$. Hence, by Remark 4.5.1,
(5.8) $\begin{array}[]{rl}&|G_{(e_{1},\ldots,e_{l})}\cdot(e_{l+1},\ldots,e_{l+k})|=|G_{(e_{1},\ldots,e_{l})}\cdot\mathfrak{p}|=\\ =&\displaystyle{\Big{|}\{\mathfrak{q}=(f_{1},\ldots,f_{l+k})\in\mathrm{Geod}_{T}(e_{1}\to T)\,:\,\pi(\mathfrak{q})=\pi(\mathfrak{p}),\forall\,i\leq l,\,f_{i}=e_{i}\}\Big{|}}\\ =&\displaystyle{\prod_{i=l}^{l+k-1}N_{\mathrm{edg}}(a_{i},a_{i+1})=N_{\mathrm{edg}}(a_{l},\ldots,a_{l+k}).}\end{array}$
This yields the “if” part of the statement. For the “only if” part, let $\rho=(a_{1},\ldots,a_{k+1})$ be an arbitrary path in $\Gamma$ which can be lifted to a geodesic $\mathfrak{p}=(e_{1},\ldots,e_{k+1})$ in $T$. Remark 4.5.1 now yields $|G_{e_{1}}\cdot\mathfrak{p}|=N_{\mathrm{edg}}(\rho)\geq 2$. ∎
Corollary 5.2.
Let $(G,T)$ be a weakly locally $\infty$-transitive group action on a tree with quotient graph $\Gamma$ and standard edge weight $\omega$. Assume that $\omega(E\Gamma)\subseteq\mathbb{Z}_{\geq 2}$. Then the following are equivalent:
  • (i)
    $(G,T)$ has property $(\ast_{k})$ for some $k\geq 2$;
  • (ii)
    $(G,T)$ has property $(\ast_{2})$;
  • (iii)
    $\omega(a)\geq 3$ or $\omega(\bar{a})\geq 3$ for every $a\in E\Gamma$.
Moreover $(G,T)$ has property $(\ast_{1})$ if, and only if, $\omega(E\Gamma)\subseteq\mathbb{Z}_{\geq 3}$.
By Remark 4.5.1, the hypothesis that $\omega(E\Gamma)\subseteq\mathbb{Z}_{\geq 2}$ guarantees that all paths in $\Gamma$ can be lifted to a geodesic.
Proof.
For the first part of the statement, by Proposition 5.1 it suffices to prove that, given an arbitrary $k\geq 2$, condition (iii) is equivalent to have $N_{\mathrm{edg}}(\rho)\geq 2$ for all paths $\rho$ in $\Gamma$ of length $k+1$. But, given a path $(a_{1},\ldots,a_{k+1})$ in $\Gamma$ with $k\geq 2$, we have $N_{\mathrm{edg}}(a_{1},\ldots,a_{k+1})=1$ if, and only if, $a_{i+1}=\bar{a}_{i}$ and $\omega(a_{i+1})=2$ for every $1\leq i\leq k$.
For the second part of the statement, one proceeds analogously. Namely, it suffices to prove that $(G,T)$ does not have property $(\ast_{1})$ if, and only if, there is a length-$2$ path $(a_{1},a_{2})$ in $\Gamma$ such that $N_{\mathrm{edg}}(a_{1},a_{2})=1$. In turn, $N_{\mathrm{edg}}(a_{1},a_{2})=1$ is equivalent to have $a_{2}=\bar{a}_{1}$ and $\omega(a_{2})=2$. ∎
Proposition 5.3.
Let $(G,T)$ be a (P)-closed action on a tree with associated local action diagram $\Delta$, and let $\iota$ be an inversion on $\Delta$. Assume that the standard weight $\mathcal{W}$ on $\Delta$ takes values in $\mathbb{Z}_{\geq 1}$, and that $|X_{a}|\geq 2$ for every $a\in E\Gamma$. Let also $k\geq 1$. Then $(G,T)$ has property ($\ast_{k}$) if, and only if, every reduced path $\xi$ in $(\Delta,\iota)$ of length $k+1$ has $\mathcal{W}(\xi)\geq 2$.
Proof.
Let $k\geq 1$ and consider a geodesic $\mathfrak{p}=(e_{1},\ldots,e_{l+k})$ in $T$ with $l\geq 1$. By Proposition 3.1 and Corollary 4.13,
$\begin{split}|G_{(e_{1},\ldots,e_{l})}\cdot(e_{l+1},\ldots,e_{l+k})|&=|G_{e_{l}}\cdot(e_{l+1},\ldots,e_{l+k})|\\ &=\mathcal{W}(g\cdot(e_{l},\ldots,e_{l+k})),\end{split}$
for some $g\in G_{o(e_{l})}$. This yields the “if” part of the statement. For the “only if” part, let $\xi=(x_{1},\ldots,x_{k+1})\in\mathcal{P}_{(\Delta,\iota)}$ and $e\in ET^{+}$ such that $\mathcal{L}(e)=x_{1}$. By Lemma 3.5, there is a geodesic $\mathfrak{p}=(e_{1},\ldots,e_{k+1})$ in $T$ with $e_{1}=e$ such that $\mathcal{L}(\mathfrak{p})=\xi$. By Proposition 4.11 we conclude that
$|G_{e_{1}}\cdot(e_{2},\ldots,e_{k})|=\mathcal{W}(\xi)\geq 2.\qed$
5.3. Convergence properties
The main goal of what follows is to study the double-coset property and the polynomial double-coset growth of triples $(G,G_{t_{1}},G_{t_{2}})$, where $G$ is a group acting on a tree $T$ and $t_{1},t_{2}\in T$.
Lemma 5.4.
Let $(G,T)$ be a group action on a tree.
  • (i)
    Assume that $C:=\sup_{e,f\in ET\,:\,t(e)=o(f)}|G_{e}\cdot f|$ is finite. Then, for every geodesic $\mathfrak{p}=(e_{1},\ldots,e_{l})$ in $T$ of length $l\geq 1$, we have $|G_{e_{1}}:G_{\mathfrak{p}}|\leq C^{l-1}$.
  • (ii)
    Suppose that $(G,T)$ have property ($\ast_{k}$) for some $k\geq 1$. Then, for every geodesic $\mathfrak{p}=(e_{1},\ldots,e_{l})$ in $T$ of length $l\geq 1$, we have $|G_{e_{1}}:G_{\mathfrak{p}}|\geq 2^{\frac{l-k}{k}}$.
Proof.
Let $\mathfrak{p}=(e_{1},\ldots,e_{l})$ be a geodesic in $T$ of length $l\geq 1$. Arguing as for (4.6) we have
(5.9) $|G_{e_{1}}:G_{\mathfrak{p}}|=\prod_{i=1}^{l-1}|G_{(e_{1},\ldots,e_{i})}\cdot e_{i+1}|.$
Since $|G_{(e_{1},\ldots,e_{i})}\cdot e_{i+1}|\leq|G_{e_{i}}\cdot e_{i+1}|$ for every $11\leq i\leq l-$ 1, we obtain (i).
To prove (ii), we may assume that $l\geq k+1$. We claim that
[…]
To prove the latter inequality in (5.10), set $A_{j}=|G_{(e_{1},\ldots,e_{j})}\cdot e_{j+1}|$ for every $11\leq j\leq l-$ 1. Then the product on the right-hand side of (5.10) becomes
$\prod_{i=1}^{l-k}\prod_{j=i}^{k+i-1}A_{j}=\prod_{j=1}^{l-1}\prod_{i=\max\{1,j-k+1\}}^{\min\{j,l-k\}}A_{j}=\prod_{j=1}^{l-1}A_{j}^{\alpha_{j}},$
where, for every $11\leq j\leq l-$ 1, we put
$\alpha_{j}:=|\{i\,:\,\max\{1,j-k+1\}\leq i\leq\min\{j,l-k\}\}|.$
It remains to show that $\alpha_{j}\leq k$ for every $11\leq j\leq l-$ 1. If $j\leq k-1$, then $\alpha_{j}\leq j<k$. If $k\leq j\leq l-k$, then $\alpha_{j}=|\{i\,:\,j-k+1\leq i\leq j\}|=k$. Finally, if $j\geq k$ and $j\geq l-k$ then $\alpha_{j}=|\{i\,:\,j-k+1\leq j\leq l-k\}|=l-j\leq k$. Hence (5.10) holds. Combining (5.10), the orbit-stabiliser theorem and the fact that $(G,T)$ has property $(\ast_{k})$, we conclude that
(5.11) $\begin{split}|G_{e_{1}}:G_{\mathfrak{p}}|^{k}&\geq\prod_{i=1}^{l-k}\prod_{j=i}^{k+i-1}|G_{(e_{1},\ldots,e_{j})}:G_{(e_{1},\ldots,e_{j+1})}|=\\ &=\prod_{i=1}^{l-k}|G_{(e_{1},\ldots,e_{i})}\cdot(e_{i+1},\ldots,e_{i+k})|\geq 2^{l-k}.\qed\end{split}$
Let $(G,T)$ be a group action on a tree and consider $t_{1},t_{2}\in T$ such that $|G_{t_{1}}gG_{t_{2}}/G_{t_{2}}|<\infty$ for every $g\in G$. For $i\in\{1,2\}$, set $T_{i}=\{t_{i}\}$ if $t_{i}\in VT$ and $T_{i}=\{t_{i},\bar{t}_{i}\}$ if $t_{i}\in ET$. By Fact 4.1 and Lemma 4.2, we have
(5.12) $b_{n}(G,G_{t_{1}},G_{t_{2}})=\Big{|}\Big{\{}\mathfrak{p}\in\mathrm{Geod}_{T}(T_{1}\to G\cdot T_{2})\,:\,|G_{t_{1}}:G_{\mathfrak{p}}|=n\Big{\}}\Big{|},\quad\forall\,n\geq 1.$
Proposition 5.5.
Let $(G,T)$ be a group action on a locally finite tree with finite quotient graph, and set $M:=\sup_{v\in VT}|o^{-1}(v)|$. If $(G,T)$ has property $(\ast_{k})$ for some $k\geq 1$, then $3\leq M<\infty$ and, for all $t_{1},t_{2}\in T$ and $n\geq 1$,
$a_{n}(G,G_{t_{1}},G_{t_{2}})=O(n^{k\cdot\log(M-1)-1}).$
In particular, for all $t_{1},t_{2}\in~{}T$ the triple $(G,G_{t_{1}},G_{t_{2}})$ has polynomial double-coset growth.
Proof.
Let $t_{1},t_{2}\in T$ and consider $T_{1}$ and $T_{2}$ as defined before (5.12). We first observe that $3\leq M\leq\infty$. Clearly, $|o^{-1}(v)|=|o^{-1}(g\cdot v)|$ for all $v\in VT$ and $g\in G$. Since $T$ is locally finite and $G\backslash VT$ is finite, we have $M<\infty$. Moreover, $M\geq 2$ because $T$ is assumed to have no leaves (cf. Section 3.1). Since $(G,T)$ has property $(\ast_{k})$, then $T$ cannot be a bi-infinite line and $M\geq 3$.
Given $l\geq 0$ and $t\in T$, notice that the number of geodesics $\mathfrak{p}$ from $t$ in $T$ with $\ell(\mathfrak{p})=l$ is $\leq 1$ if $l=0$, and it is $\leq M(M-1)^{l-1}$ otherwise. Hence, by (5.12) and Lemma 5.4(ii), the following holds for every $n\geq 1$:
$\begin{split}b_{n}(G,G_{t_{1}},G_{t_{2}})&=|\{\mathfrak{p}\in\mathrm{Geod}_{T}(T_{1}\to T)\,:\,|G_{t_{1}}:G_{\mathfrak{p}}|=n\}|\\ &\leq|\{\mathfrak{p}\in\mathrm{Geod}_{T}(T_{1}\to T)\,:\,\ell(\mathfrak{p})\leq\lfloor k\cdot\log_{2}n\rfloor+k\}|\\ &=\sum_{t\in T_{1}}|\{\mathfrak{p}\in\mathrm{Geod}_{T}(t\to T)\,:\,\ell(\mathfrak{p})\leq\lfloor k\cdot\log_{2}n\rfloor+k\}|\\ &\leq\sum_{t\in T_{1}}\Bigg{(}1+\sum_{l=1}^{\lfloor k\cdot\log_{2}n\rfloor+k}M(M-1)^{l-1}\Bigg{)}\\ &=|T_{1}|\cdot\frac{M(M-1)^{\lfloor k\cdot\log_{2}n\rfloor+k}-2}{M-2}.\end{split}$
Hence,
$b_{n}(G,G_{t_{1}},G_{t_{2}})=O((M-1)^{k\cdot\log n})=O(n^{k\cdot\log(M-1)}).$
The latter claim of the statement now follows from (5.2). ∎
Theorem 5.6.
Let $(G,T)$ be a weakly locally $\infty$-transitive group action on a tree with finite quotient graph $\Gamma$. Assume that the standard edge weight $\omega$ on $\Gamma$ takes values in $\mathbb{Z}_{\geq 2}$. Then the following are equivalent for all $t_{1},t_{2}\in T$:
  • (i)
    $(G,G_{t_{1}},G_{t_{2}})$ has the double-coset property;
  • (ii)
    $(G,G_{t_{1}},G_{t_{2}})$ has polynomial double-coset growth;
  • (iii)
    $(G,T)$ has property $(\ast_{k})$ for some $k\geq 1$;
  • (iv)
    $\omega(a)\geq 3$ or $\omega(\bar{a})\geq 3$ for every $a\in E\Gamma$.
Proof.
Clearly, (ii)$\Rightarrow$(i). Moreover, Proposition 5.5 and Corollary 5.2 imply (iii)$\Rightarrow$(ii) and (iii)$\Leftrightarrow$(iv), respectively. It remains to prove (i)$\Rightarrow$(iv).
Assume that there is $a\in E\Gamma$ such that $\omega(a)=\omega(\bar{a})=2$, and let $\mathfrak{p}=(a,\bar{a})$. Then $N_{\mathrm{edg}}(\mathfrak{p}^{d})=(\omega(\bar{a})-1)^{d}(\omega(a)-1)^{d-1}=1$ for every $d\geq 1$. Consider two paths $\mathfrak{q}_{1}=(a_{1},\ldots,a_{h})$ and $\mathfrak{q}_{2}=(b_{1},\ldots,b_{k})$ of positive length in $\Gamma$ from $\pi(t_{1})$ to $o(a)$ and from $o(a)$ to $\pi(t_{2})$, respectively. Then, for every $d\geq 1$ we have
$\begin{split}N_{\mathrm{edg}}(\mathfrak{q}_{1}\cdot\mathfrak{p}^{d}\cdot\mathfrak{q}_{2})&=N_{\mathrm{edg}}(\mathfrak{q}_{1})N_{\mathrm{edg}}(a_{h},a)N_{\mathrm{edg}}(\mathfrak{p}^{d})N_{\mathrm{edg}}(\bar{a},b_{1})N_{\mathrm{edg}}(\mathfrak{q}_{2})\\ &=N_{\mathrm{edg}}(\mathfrak{q}_{1})N_{\mathrm{edg}}(a_{h},a)N_{\mathrm{edg}}(\bar{a},b_{1})N_{\mathrm{edg}}(\mathfrak{q}_{2})=:N\geq 1.\end{split}$
Since $\omega$ takes values in $\mathbb{Z}_{\geq 2}$, for every $d\geq 1$ there is $\tilde{\mathfrak{q}}_{d}\in\mathrm{Geod}_{T}(T_{1}\to G\cdot T_{2})$ satisfying $\pi(\tilde{\mathfrak{q}}_{d})=\mathfrak{q}_{1}\cdot\mathfrak{p}^{d}\cdot\mathfrak{q}_{2}$ (cf. Remark 4.5.2). By Proposition 4.6, for every $d\geq 1$ we have
$|G_{t_{1}}:G_{\tilde{\mathfrak{q}}_{d}}|=\left\{\begin{array}[]{cl}N_{\mathrm{edg}}(\mathfrak{q}_{1}\cdot\mathfrak{p}^{d}\cdot\mathfrak{q}_{2})=N,&\text{if }t_{1}\in ET;\\ N_{\mathrm{vert}}(\mathfrak{q}_{1}\cdot\mathfrak{p}^{d}\cdot\mathfrak{q}_{2})=\omega(a_{1})\cdot N=:N^{\prime},&\text{if }t_{1}\in VT.\end{array}\right.$
Since $\tilde{\mathfrak{q}}_{d}\neq\tilde{\mathfrak{q}}_{d^{\prime}}$ for all $d\neq d^{\prime}$, by (5.12) we conclude that $b_{N}(G,G_{t_{1}},G_{t_{2}})=\infty$ if $t_{1}\in ET$, and $b_{N^{\prime}}(G,G_{t_{1}},G_{t_{2}})=\infty$ if $t_{1}\in VT$. ∎
Theorem 5.7.
Let $(G,T)$ be a (P)-closed action on a locally finite tree. Assume that the quotient graph is finite and its standard edge weight takes values in $\mathbb{Z}_{\geq 2}$. Then the following are equivalent for all $t_{1},t_{2}\in T$:
  • (i)
    $(G,G_{t_{1}},G_{t_{2}})$ has the double-coset property;
  • (ii)
    $(G,G_{t_{1}},G_{t_{2}})$ has polynomial double-coset growth;
  • (iii)
    $(G,T)$ has property $(\ast_{k})$ for some $k\geq 1$.
Proof.
The implication (ii)$\Rightarrow$(i) is immediate, and (iii)$\Rightarrow$(ii) follows from Proposition 5.5. It remains to prove (i)$\Rightarrow$(iii). Assume that $(G,T)$ does not have property $(\ast_{k})$ for every $k\geq 1$. Let $\Delta$ be the local action diagram associated to $(G,T)$ and consider an inversion $\iota$ in $\Delta$. Without loss of generality, $G=U(\Delta,\iota,c_{0})$ and $T=T(\Delta,\iota,c_{0})$ for some $c_{0}\in V\Gamma$ (cf. Theorem 3.8). By Proposition 5.3, there is a reduced path $(x_{1},\ldots,x_{k})$ in $(\Delta,\iota)$ of length $k\geq|X|^{2}+2$ such that $\mathcal{W}(x_{1},\ldots,x_{k})=1$, i.e., $\mathcal{W}(x_{i},x_{i+1})=1$ for every $11\leq i\leq k-$ 1. Since $k\geq|X|^{2}+2$, we have $(x_{i},x_{i+1})=(x_{i+l},x_{i+1+l})$ for some $i,l\geq 1$. Set $\eta:=(x_{j})_{i\leq j\leq i+l-1}$. Hence, for every $d\geq 1$ the power $\eta^{d}$ is a reduced path in $(\Delta,\iota)$ satisfying
$\mathcal{W}(\eta^{d})=\mathcal{W}(\eta)^{d}\cdot\mathcal{W}(x_{i+l-1},x_{i})^{d-1}=\mathcal{W}(\eta)^{d}\cdot\mathcal{W}(x_{i+l-1},x_{i+l})^{d-1}=1.$
Set $y\in X_{\pi(t_{1})}$, $z\in X_{\pi(t_{2})}$ and choose arbitrary reduced paths of positive length in $(\Delta,\iota)$, say $\xi_{1}=(y_{1},\ldots,y_{h})$ and $\xi_{2}=(z_{1},\ldots,z_{r})$, such that $\xi_{1}\cdot x_{1}\in\mathcal{P}_{(\Delta,\iota)}(y\to x_{1})$ and $x_{i+l-1}\cdot\xi_{2}\in\mathcal{P}_{(\Delta,\iota)}(x_{i+l-1}\to z)$. Such reduced paths exist because the standard edge weight of $G\backslash T$ takes values in $\mathbb{Z}_{\geq 2}$ (cf. Remark 4.5.1). For every $d\geq 1$, the path $\xi_{1}\cdot\eta^{d}\cdot\xi_{2}$ is reduced in $(\Delta,\iota)$ and
$\begin{split}\mathcal{W}(\xi_{1}\cdot\eta^{d}\cdot\xi_{2})&=\mathcal{W}(\xi_{1})\mathcal{W}(y_{h},x_{i})\mathcal{W}(\eta^{d})\mathcal{W}(x_{i+l-1},z_{1})\mathcal{W}(\xi_{2})\\ &=\mathcal{W}(\xi_{1})\mathcal{W}(y_{h},x_{i})\mathcal{W}(x_{i+l-1},z_{1})\mathcal{W}(\xi_{2})=:N.\end{split}$
By Lemma 3.5, for every $d\geq 1$ there is $\tilde{\mathfrak{q}}_{d}\in\mathrm{Geod}_{T}(T_{1}\to G\cdot T_{2})$ such that $\mathcal{L}(\tilde{\mathfrak{q}}_{d})=\xi_{1}\cdot\eta^{d}\cdot\xi_{2}$. By Corollary 4.13, we may assume that all edges of $\tilde{\mathfrak{q}}_{d}$ are in $ET^{+}$ for every $n\geq 1$. Note that each $\tilde{\mathfrak{q}}_{d}$ has the same first edge, say $e_{1}$. By Proposition 4.11, for every $d\geq 1$ we have
$|G_{t_{1}}:G_{\tilde{\mathfrak{q}}_{d}}|=\left\{\begin{array}[]{cl}\mathcal{W}(\mathcal{L}(\tilde{\mathfrak{q}}_{d}))=N,&\text{if }t_{1}\in ET;\\ \omega(\pi(e_{1}))\cdot\mathcal{W}(\mathcal{L}(\tilde{\mathfrak{q}}_{d}))=:N^{\prime},&\text{if }t_{1}\in VT.\end{array}\right.$
Since $\tilde{\mathfrak{q}}_{d}\neq\tilde{\mathfrak{q}}_{d^{\prime}}$ for all $d\neq d^{\prime}$, from (5.12) we conclude that $b_{N}(G,G_{t_{1}},G_{t_{2}})=\infty$ if $t_{1}\in ET$, and $b_{N^{\prime}}(G,G_{t_{1}},G_{t_{2}})=\infty$ if $t_{1}\in VT$. ∎
5.4. Explicit formulae: the weakly locally $\infty$-transitive case
Setting [WLIT].
Let $(G,T)$ be a weakly locally $\infty$-transitive group action on a locally finite tree with quotient map $\pi\colon T\to\Gamma=G\backslash T$. Assume that $\Gamma$ is finite, and that its standard edge weight $\omega$ takes values in $\mathbb{Z}_{\geq 2}$ and satisfies $\omega(a)\geq 3$ or $\omega(\bar{a})\geq 3$ for every $a\in E\Gamma$. Let also $N_{\mathrm{edg}}=N_{\mathrm{edg}}^{\omega}$ and $N_{\mathrm{vert}}=N_{\mathrm{vert}}^{\omega}$ be as in Definition 4.4.
Setting [WLIT] guarantees that the series defining $\zeta_{G,G_{t_{1}},G_{t_{2}}}(s)$ converges at some $s\in\mathbb{C}$, for all $t_{1},t_{2}\in T$ (cf. Theorem 5.6).
Proposition 5.8.
Suppose Setting [WLIT], and let $t\in T$ with $\pi(t)=u$. Then, for every $v\in VT$ with $\pi(v)=c$, we have
$\zeta_{G,G_{v},G_{t}}(s)=\sum_{\mathfrak{p}\in\mathcal{P}_{\Gamma}(c\to U)}N_{\mathrm{vert}}(\mathfrak{p})^{-s}.$
Moreover, for every $e\in ET$ with $\pi(e)=a$, we have
$\zeta_{G,G_{e},G_{t}}(s)=\varepsilon_{a}(u)+\sum_{\mathfrak{p}\in\mathcal{P}_{\Gamma}(A\to U),\atop\ell(\mathfrak{p})\geq 2}N_{\mathrm{edg}}(\mathfrak{p})^{-s},$
where $\varepsilon_{a}(u)=\mathbbm{1}_{\{o(a),t(a)\}}(u)$ if $u\in V\Gamma$ and $\varepsilon_{a}(u)=\mathbbm{1}_{A}(u)$ if $u\in E\Gamma$.
Proof.
It is a direct consequence of Proposition 4.3 (recalling Remark 4.5.2) and Proposition 4.6. ∎
Proposition 5.8 suggests the following generalisation.
Definition 5.9.
Let $\Gamma$ be a non-empty graph with an edge weight $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 2}$, and let $c\in V\Gamma$, $a\in E\Gamma$ and $u\in\Gamma$. Define the following formal Dirichlet series:
$\begin{split}\mathcal{Z}_{\Gamma,c\to u}(s)&:=\sum_{\mathfrak{p}\in\mathcal{P}_{\Gamma}(c\to U)}N_{\mathrm{vert}}(\mathfrak{p})^{-s},\\ \mathcal{Z}_{\Gamma,a\to u}(s)&:=\varepsilon_{a}(u)+\sum_{\mathfrak{p}\in\mathcal{P}_{\Gamma}(A\to u),\atop\ell(\mathfrak{p})\geq 2}N_{\mathrm{edg}}(\mathfrak{p})^{-s},\end{split}$
where $\varepsilon_{a}(u)=\mathbbm{1}_{\{o(a),t(a)\}}(u)$ if $u\in V\Gamma$ and $\varepsilon_{a}(u)=\mathbbm{1}_{A}(u)$ if $u\in E\Gamma$.
Remark 5.9.1.
In Definition 5.9, we may assume that $\Gamma$ is connected. Indeed, given $u_{1},u_{2}\in\Gamma$, if there is a connected component $\Lambda$ of $\Gamma$ containing both $u_{1}$ ad $u_{2}$, then $\mathcal{Z}_{\Gamma,u_{1}\to u_{2}}(s)=\mathcal{Z}_{\Lambda,u_{1}\to u_{2}}(s)$. If such a connected component does not exist, the function $\mathcal{Z}_{\Gamma,u_{1}\to u_{2}}(s)$ is identically zero.
Remark 5.9.2.
By Proposition 5.8, for all $t_{1},t_{2}\in T$ with $\pi(t_{1})=u_{1}$ and $\pi(t_{2})=u_{2}$ we have
$\zeta_{G,G_{t_{1}},G_{t_{2}}}(s)=\mathcal{Z}_{\Gamma,u_{1}\to u_{2}}(s).$
In view of an explicit formula for $\mathcal{Z}_{\Gamma,u_{1}\to u_{2}}(s)$, we introduce the following linear operator.
Definition 5.10.
Let $\Gamma$ be a non-empty graph with an edge weight $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 2}$. Let $\mathbb{C}\llbracket E\Gamma\rrbracket$ be the complex vector space of all formal sums $\sum_{a\in E\Gamma}\gamma_{a}a$, where $\gamma_{a}\in\mathbb{C}$ for every $a\in E\Gamma$. For each $u\in\Gamma$, define $e_{u}\in\mathbb{C}\llbracket E\Gamma\rrbracket$ as follows:
(5.13) $e_{u}:=\left\{\begin{array}[]{cl}u,&\text{if }u\in E\Gamma;\\ \sum_{a\in t^{-1}(u)}a,&\text{if }u\in V\Gamma.\end{array}\right.$
For every $s\in\mathbb{C}$, the Bass operator $\mathcal{E}(s)=\mathcal{E}^{(\Gamma,\omega)}(s)\colon\mathbb{C}\llbracket E\Gamma\rrbracket\to\mathbb{C}\llbracket E\Gamma\rrbracket$ of $\Gamma$ at $s\in\mathbb{C}$ is the linear extension of the following assignment:
(5.14) $\mathcal{E}(s)(a):=\sum_{b\in E\Gamma}\mathcal{E}(s)(a,b)b,\quad\forall\,a\in E\Gamma$
where, for all $a,b\in E\Gamma$,
(5.15) $\mathcal{E}(s)(a,b):=\left\{\begin{array}[]{cl}N_{\mathrm{edg}}(a,b)^{-s},&\text{if }t(a)=o(b);\\ 0,&\text{otherwise.}\end{array}\right.$
Notation 5.11.
  • (i)
    We will usually write $\mathcal{E}(s)$ instead of $\mathcal{E}^{(\Gamma,\omega)}(s)$. If we want to specify $\Gamma$ (but $\omega$ is clear from the context), we write $\mathcal{E}^{\Gamma}(s)$ instead of $\mathcal{E}(s)$ or $\mathcal{E}^{(\Gamma,\omega)}(s)$, and $e^{\Gamma}_{u}$ instead of $e_{u}$, for all $u\in\Gamma$.
  • (ii)
    We implicitly set a total order on $E\Gamma$. Thus, provided $|E\Gamma|<\infty$, we can regard $\mathcal{E}(s)$ as a $|E\Gamma|$-dimensional matrix $[\mathcal{E}(s)(a,b)]_{a,b\in E\Gamma}$ with complex entries, and the $e_{u}$’s in (5.13) as row vectors in $\mathbb{C}^{|E\Gamma|}$. For all $a,b\in E\Gamma$, note that $e_{a}\mathcal{E}(s)e_{b}^{t}=\mathcal{E}(s)(a,b)$.
The term “Bass operator” is taken after [9, Definition 3.10]. The reader is referred to Section 7.3 for further connections with [9].
Remark 5.11.1.
Let $\mathcal{E}(s)$ be as in Definition 5.10. For every $n\geq 1$, let $\mathcal{E}(s)^{n}$ be the $n$-th power of $\mathcal{E}(s)$, and $\mathcal{E}(s)^{0}$ be the identity operator on $\mathbb{C}\llbracket E\Gamma\rrbracket$. Then, for all $n\geq 0$ and $a,b\in E\Gamma$, we observe that
$e_{a}\mathcal{E}(s)^{n}e_{b}^{t}=\mathcal{E}(s)^{n}(a,b)=\sum_{\mathfrak{p}\in\mathcal{P}_{\Gamma}(a\to b)\atop\ell(\mathfrak{p})=n+1}N_{\mathrm{edg}}(\mathfrak{p})^{-s}.$
If $n\leq 1$, it is clear. For every $n\geq 2$, it suffices to observe that
$\mathcal{E}(s)^{n}(a,b)=\sum_{a_{2},\ldots,\,a_{n}\in E\Gamma}\mathcal{E}(s)(a,a_{2})\cdot\ldots\cdot\mathcal{E}(s)(a_{n},b).$
As we did in Setting [WLIT], we fix a setting which ensures that the series $\mathcal{Z}_{\Gamma,u_{1}\to u_{2}}(s)$ as in Definition 5.9 converges at some $s\in\mathbb{C}$, for all $u_{1},u_{2}\in\Gamma$.
Setting [$\Gamma$].
Let $\Gamma$ be a finite connected non-empty graph with an edge weight $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 2}$ satisfying $\omega(a)\geq 3$ or $\omega(\bar{a})\geq 3$, for every $a\in E\Gamma$.
Convention: Every subgraph $\Gamma^{\prime}\subseteq\Gamma$ is endowed “by default” with the restricted edge weight from $\omega$.
Theorem 5.12.
Let $(\Gamma,\omega)$ be an edge-weighted graph satisfying Setting [$\Gamma$]. Then, for all $u,w\in\Gamma$,
(5.16) $\mathcal{Z}_{\Gamma,u\to w}(s)=\frac{\det(I-\mathcal{E}(s)+\mathcal{U}_{u,w}(s))}{\det(I-\mathcal{E}(s))}+\epsilon_{u}(w),$
where $I$ be the identity matrix in $\mathrm{Mat}_{|E\Gamma|}(\mathbb{C})$ and
$\mathcal{U}_{u,w}(s):=\left\{\begin{array}[]{ll}\sum_{a\in o^{-1}(u)}\omega(a)^{-s}e_{w}^{t}\cdot e_{a},&\text{if }u,w\in V\Gamma;\\ \sum_{a\in o^{-1}(u)}\omega(a)^{-s}(e_{w}+e_{\bar{w}})^{t}\cdot e_{a}\mathcal{E}(s),&\text{if }u\in V\Gamma,\,w\in E\Gamma;\\ e_{w}^{t}\cdot(e_{u}+e_{\bar{u}})\mathcal{E}(s),&\text{if }u\in E\Gamma,\,w\in V\Gamma;\\ (e_{w}+e_{\bar{w}})^{t}\cdot(e_{u}+e_{\bar{u}})\mathcal{E}(s),&\text{if }u,w\in E\Gamma;\end{array}\right.$
$\epsilon_{u}(w):=\left\{\begin{array}[]{ll}\mathbbm{1}_{\{u\}}(w)-1,&\text{if }u,w\in V\Gamma;\\ \mathbbm{1}_{o^{-1}(u)}(w)\cdot\omega(w)-1,&\text{if }u\in V\Gamma,\,w\in E\Gamma;\\ \mathbbm{1}_{\{o(u),t(u)\}}(w)-1,&\text{if }u\in E\Gamma,\,w\in V\Gamma;\\ \mathbbm{1}_{\{u,\bar{u}\}}(w)-1,&\text{if }u,w\in E\Gamma.\end{array}\right.$
In particular, for all $u,w\in\Gamma$ the function $\mathcal{Z}_{\Gamma,u\to w}(s)$ is a meromorphic function over $\mathbb{C}$.
The proof of Theorem 5.12 makes use of the following fact.
Fact 5.13 (Matrix Determinant Lemma, cf. [15]).
Consider $A\in\mathrm{GL}_{n}(\mathbb{C})$ with adjugate matrix $\mathrm{adj}(A)$, and let $u,v\in\mathbb{C}^{n}$ be row vectors. Then,
$\frac{\det(A+u^{t}\cdot v)}{\det(A)}=1+vA^{-1}u^{t}.$
Proof of Theorem 5.12.
Let $s\in\mathbb{C}$ be such that $\sum_{n=0}^{\infty}\mathcal{E}(s)$ converges. Recall that $\sum_{n=0}^{\infty}\mathcal{E}(s)=(I-\mathcal{E}(s))^{-1}$. By Remark 5.11.1, if $u\in V\Gamma$ then
$\begin{split}&\mathcal{Z}_{\Gamma,u\to w}(s)=\\ &=\left\{\begin{array}[]{ll}\displaystyle{\mathbbm{1}_{\{u\}}(w)+\sum_{n=0}^{\infty}\sum_{a\in o^{-1}(u)}\omega(a)^{-s}e_{a}\mathcal{E}(s)^{n}e_{w}^{t}},&\text{if }w\in V\Gamma;\\ \displaystyle{\mathbbm{1}_{o^{-1}(u)}(w)\cdot\omega(w)^{-s}+\sum_{n=1}^{\infty}\sum_{a\in o^{-1}(u)}\omega(a)^{-s}e_{a}\mathcal{E}(s)^{n}(e_{w}+e_{\bar{w}})^{t}},&\text{if }w\in E\Gamma.\end{array}\right.\end{split}$
Similarly, if $u\in E\Gamma$ then
$\begin{split}\mathcal{Z}_{\Gamma,u\to w}(s)=\left\{\begin{array}[]{ll}\displaystyle{\mathbbm{1}_{\{o(u),t(u)\}}(w)+\sum_{n=1}^{\infty}(e_{u}+e_{\bar{u}})\mathcal{E}(s)^{n}e_{w}^{t}},&\text{if }w\in V\Gamma;\\ \displaystyle{\mathbbm{1}_{\{u,\bar{u}\}}(w)+\sum_{n=1}^{\infty}(e_{u}+e_{\bar{u}})\mathcal{E}(s)^{n}(e_{w}^{t}+e_{\bar{w}}^{t})},&\text{if }w\in E\Gamma.\end{array}\right.\end{split}$
We now focus on the case in which $u,w\in V\Gamma$, as the other cases are analogous. Namely, if $u,w\in V\Gamma$ then
$\begin{split}\mathcal{Z}_{\Gamma,u\to w}(s)&=\mathbbm{1}_{\{u\}}(w)+\sum_{a\in o^{-1}(u)}\omega(a)^{-s}e_{a}\Bigg{(}\sum_{n=0}^{\infty}\mathcal{E}(s)^{n}\Bigg{)}e_{w}^{t}\\ &=\mathbbm{1}_{\{u\}}(w)+\Bigg{(}\sum_{a\in o^{-1}(u)}\omega(a)^{-s}e_{a}\Bigg{)}(I-\mathcal{E}(s))^{-1}e_{w}^{t}\end{split}$
and Fact 5.13 applies. ∎
In view of Section 7, we provide some explicit formulae for $\mathcal{Z}_{\Gamma,u\to u}(s)$ in case that $\Gamma$ has one edge-pair.
Example 5.14.
Let $\Gamma$ be a $1$-segment with $E\Gamma=\{a,\bar{a}\}$, $c=o(a)$ and $d=t(a)$. Set $\omega(a):=\alpha+1$ and $\omega(\bar{a}):=\beta+1$, where $\alpha,\beta\in\mathbb{Z}_{\geq 1}$ with $\alpha\geq 2$ or $\beta\geq 2$. Set the order $\leq$ on $E\Gamma$ such that $a<\bar{a}$, and identify $\mathbb{C}\llbracket E\Gamma\rrbracket$ with $\mathbb{C}^{2}$, $e_{a}=e_{d}$ with $(1,0)$ and $e_{\bar{a}}=e_{c}=(0,1)$. Then, for every $s\in\mathbb{C}$,
$\begin{split}\mathcal{E}(s)&=\begin{bmatrix}\mathcal{E}(s)(a,a)&\mathcal{E}(s)(a,\bar{a})\\ \mathcal{E}(s)(\bar{a},a)&\mathcal{E}(s)(\bar{a},\bar{a})\end{bmatrix}=\begin{bmatrix}0&\beta^{-s}\\ \alpha^{-s}&0\end{bmatrix};\\ \mathcal{U}_{c,c}(s)&=(\alpha+1)^{-s}e_{c}^{t}\cdot e_{a}=\begin{bmatrix}0&0\\ (\alpha+1)^{-s}&0\end{bmatrix};\\ \mathcal{U}_{a,a}(s)&=(e_{a}+e_{\bar{a}})^{t}\cdot(e_{a}+e_{\bar{a}})\cdot\mathcal{E}(s)=\begin{bmatrix}\alpha^{-s}&\beta^{-s}\\ \alpha^{-s}&\beta^{-s}\end{bmatrix}.\end{split}$
Let $I$ be the identity matrix in $\mathrm{Mat}_{2}(\mathbb{C})$. By Theorem 5.12,
(5.17) $\begin{split}\mathcal{Z}_{\Gamma,c\to c}(s)&=\frac{1+((\alpha+1)^{-s}-\alpha^{-s})\cdot\beta^{-s}}{1-\alpha^{-s}\beta^{-s}};\\ \mathcal{Z}_{\Gamma,a\to a}(s)&=\frac{(1+\alpha^{-s})(1+\beta^{-s})}{1-\alpha^{-s}\beta^{-s}}.\end{split}$
In particular, let $(G,T)$ be a locally $\infty$-transitive action on a locally finite tree with quotient graph $\Gamma$ and standard edge weight $\omega$ as before. By Remark 5.9.2 and (5.17), we have explicit formulae for $\zeta_{G,G_{v},G_{v}}(s)$ and $\zeta_{G,G_{e},G_{e}}(s)$ for all $v\in VT$ with $G\cdot v=c$ and $e\in ET$ with $G\cdot e=a$. For instance, one may take $G=\mathrm{SL}_{2}(\mathbb{Q}_{p})$ and $T$ the Bruhat–Tits tree of $G$ (cf. Example 3.3(ii)). In this case $\alpha=\beta=p$. Let $v\in VT$ be the vertex with $G_{v}=\mathrm{SL}_{2}(\mathbb{Z}_{p})$, and $e\in ET$ be the edge whose pointwise stabiliser is the standard Iwahori subgroup. Then,
$\zeta_{G,G_{v},G_{v}}(s)=\frac{1+((p+1)^{-s}-p^{-s})\cdot p^{-s}}{1-p^{-2s}}\quad\text{and}\quad\zeta_{G,G_{e},G_{e}}(s)=\frac{1+p^{-s}}{1-p^{-s}}.$
Other examples can be obtained from Example 3.12. Note that the formulae before agree with [8, Example 1.7] in case that $G$ is the group of automorphisms of a bi-coloured tree $T$.
Example 5.15.
Let $\Gamma$ be a $1$-bouquet of loops with $E\Gamma=\{a,\bar{a}\}$ and $c=o(a)=t(a)$. Set $\omega(a):=\alpha+1$ and $\omega(\bar{a}):=\beta+1$, for some $\alpha,\beta\in\mathbb{Z}_{\geq 1}$ with $\alpha\geq 2$ or $\beta\geq 2$. Consider the order $\leq$ on $E\Gamma$ such that $a<\bar{a}$, and identify $\mathbb{C}\llbracket E\Gamma\rrbracket$ with $\mathbb{C}^{2}$, $e_{a}$ with the vector $(1,0)$, $e_{\bar{a}}$ with $(0,1)$ and $e_{c}=e_{a}+e_{\bar{a}}$ with $(1,1)$. Then, for every $s\in\mathbb{C}$,
$\begin{split}\mathcal{E}(s)&=\begin{bmatrix}\mathcal{E}(s)(a,a)&\mathcal{E}(s)(a,\bar{a})\\ \mathcal{E}(s)(\bar{a},a)&\mathcal{E}(s)(\bar{a},\bar{a})\end{bmatrix}=\begin{bmatrix}(\alpha+1)^{-s}&\beta^{-s}\\ \alpha^{-s}&(\beta+1)^{-s}\end{bmatrix};\\ \mathcal{U}_{c,c}(s)&=(\alpha+1)^{-s}e_{c}^{t}\cdot e_{a}+(\beta+1)^{-s}e_{c}^{t}\cdot e_{\bar{a}}=\begin{bmatrix}(\alpha+1)^{-s}&(\beta+1)^{-s}\\ (\alpha+1)^{-s}&(\beta+1)^{-s}\end{bmatrix};\\ \mathcal{U}_{a,a}(s)&=(e_{a}+e_{\bar{a}})^{t}\cdot(e_{a}+e_{\bar{a}})\mathcal{E}(s)=\begin{bmatrix}(\alpha+1)^{-s}+\alpha^{-s}&(\beta+1)^{-s}+\beta^{-s}\\ (\alpha+1)^{-s}+\alpha^{-s}&(\beta+1)^{-s}+\beta^{-s}\end{bmatrix}.\end{split}$
Let $I$ be the identity matrix in $\mathrm{Mat}_{2}(\mathbb{C})$. By Theorem 5.12, we have
(5.18) $\begin{split}\mathcal{Z}_{\Gamma,c\to c}(s)&=\frac{1-\Big{(}(\alpha+1)^{-s}-\alpha^{-s}\Big{)}\cdot\Big{(}(\beta+1)^{-s}-\beta^{-s}\Big{)}}{\Big{(}1-(\alpha+1)^{-s}\Big{)}\cdot\Big{(}1-(\beta+1)^{-s}\Big{)}-\alpha^{-s}\beta^{-s}};\\ \mathcal{Z}_{\Gamma,a\to a}(s)&=\frac{(\alpha^{-s}+1)(\beta^{-s}+1)-(\alpha+1)^{-s}(\beta+1)^{-s}}{\Big{(}1-(\alpha+1)^{-s}\Big{)}\cdot\Big{(}1-(\beta+1)^{-s}\Big{)}-\alpha^{-s}\beta^{-s}}.\end{split}$
If $\alpha=\beta$, after basic algebraic manipulations, the formulae in (5.18) become
(5.19) $\mathcal{Z}_{\Gamma,c\to c}(s)=\frac{1-\alpha^{-s}+(\alpha+1)^{-s}}{1-\alpha^{-s}-(\alpha+1)^{-s}}\quad\text{and}\quad\mathcal{Z}_{\Gamma,a\to a}(s)=\frac{1+\alpha^{-s}+(\alpha+1)^{-s}}{1-\alpha^{-s}-(\alpha+1)^{-s}}.$
Consider a weakly locally $\infty$-transitive group action on a locally finite tree $(G,T)$ with quotient graph $\Gamma$ and standard edge weight $\omega$. For explicit examples, see Example 3.12(i). By Remark 5.9.2, the computations in (5.18) provide explicit formulae for $\zeta_{G,G_{v},G_{v}}(s)$ and $\zeta_{G,G_{e},G_{e}}(s)$ whenever $v\in VT$ and $e\in ET$ satisfy $G\cdot v=c$ and $G\cdot e=a$ (or $G\cdot e=\bar{a}$), respectively.
5.5. Explicit formulae: the (P)-closed case
Setting [(P)-cl].
Let $\Delta=(\Gamma,(X_{a}),(G(c)))$ be a local action diagram based on a non-empty finite connected graph $\Gamma$. Choose an inversion $\iota$ in $\Delta$ and $c_{0}\in V\Gamma$. Denote by $(T=T(\Delta,\iota,c_{0}),\pi,\mathcal{L})$ the standard $\Delta$-tree associated to $\iota$ and $c_{0}$, let $v_{0}$ be the root of $T$ (cf. Section 3.4.1), and set $G=U(\Delta,\iota,c_{0})$. Assume that the standard edge weights on $\Gamma$ and $\Delta$, denoted by $\omega$ and $\mathcal{W}$ respectively, take values in $\mathbb{Z}_{\geq 2}$ and $\mathbb{Z}_{\geq 0}$, respectively. Finally, assume that $(G,T)$ has property $(\ast_{k})$ for some $k\geq 1$.
Setting [(P)-cl] guarantees that the series defining $\zeta_{G,G_{t_{1}},G_{t_{2}}}(s)$ converges at some $s\in\mathbb{C}$, for all $t_{1},t_{2}\in T$ (cf. Theorem 5.7). Thanks to the following remark, from now on we may focus only on the case in which $t_{1}\in\{v_{0}\}\cup o^{-1}(v_{0})$ while studying $\zeta_{G,G_{t_{1}},G_{t_{2}}}(s)$.
Remark 5.15.1.
Let $\Delta$ be a local action diagram and $\mathbb{T}=(T,\pi,\mathcal{L})$ be $\Delta$-tree. Consider an inversion $\iota$ in $\Delta$ and denote by $(T(\Delta,\iota,c_{0}),\pi_{0},\mathcal{L}_{0})$ the standard $\Delta$-tree associated to $\iota$ and some $c_{0}\in V\Gamma$. By Theorem 3.8(i), there is a graph isomorphism $\phi\colon T\to T(\Delta,\iota,c_{0})$ such that $\pi_{0}=\pi\circ\phi$ and $U(\Delta,\iota,c_{0})=\phi U(\Delta,\mathbb{T})\phi^{-1}$. Set $G=U(\Delta,\mathbb{T})$ and $H=U(\Delta,\iota,c_{0})$. Then, for all $t\in T$ the following map is bijective:
$G/G_{t}\longrightarrow H/H_{\phi(t)},\quad gG_{t}\longmapsto\phi g\phi^{-1}H_{\phi(t)}.$
Moreover, for all $t_{1},t_{2}\in T$ and $g\in G$, provided $h:=\phi g\phi^{-1}$ we have
$|G_{t_{1}}\cap G_{t_{1}}\cap gG_{t_{2}}g^{-1}|=|H_{\phi(t_{1})}:H_{\phi(t_{1})}\cap hH_{\phi(t_{2})}h^{-1}|$
and then
$\zeta_{G,G_{t_{1}},G_{t_{2}}}(s)=\zeta_{H,H_{\phi(t_{1})},H_{\phi(t_{2})}}(s),$
whenever one series before is defined.
In particular, by Theorem 3.8(i), given $v\in VT$ with $\pi(v)=:c_{0}$ one may take $\phi$ so that $\phi(v)$ is the root $v_{0}$ of $T(\Delta,\iota,c_{0})$. Then, for all $t\in T$,
$\zeta_{G,G_{v},G_{t}}(s)=\zeta_{H,H_{v_{0}},H_{\phi(t)}}(s).$
Moreover, for all $e\in o^{-1}(v)$ we have $\phi(e)\in o^{-1}(v_{0})$ and, for all $t\in T$,
$\zeta_{G,G_{e},G_{t}}(s)=\zeta_{H,H_{\phi(e)},H_{\phi(t)}}(s).$
The analogue of Proposition 5.8 for (P)-closed actions is the following.
Proposition 5.16.
Let $(G,T)$ be as in Setting [(P)-cl], and let $t\in T$ with $\pi(t)=u$. Then
(5.20) $\zeta_{G,G_{v_{0}},G_{t}}(s)=\left\{\begin{array}[]{ll}\displaystyle{\mathbbm{1}_{\{c_{0}\}}(u)+\sum_{a\in o^{-1}(c_{0}),\atop\xi\in\mathcal{P}_{(\Delta,\iota)}(X_{a}\to X_{U})}\omega(a)^{-s-1}\mathcal{W}(\xi)^{-s-1},}&\text{if }u\in V\Gamma;\\ \displaystyle{\sum_{a\in o^{-1}(c_{0}),\atop\xi\in\mathcal{P}_{(\Delta,\iota)}(X_{a}\to X_{U})}\omega(a)^{-s-1}\mathcal{W}(\xi)^{-s-1},}&\text{if }u\in E\Gamma.\end{array}\right.$
Moreover, for every $e\in ET$ with $\pi(e)=a$ and $\mathcal{L}(e)=x\in X_{c_{0}}$, we have
(5.21) $\zeta_{G,G_{e},G_{t}}(s)=\eta_{a}(u)+\sum_{\xi\in\mathcal{P}_{(\Delta,\iota)}(x\to X_{U}),\atop\ell(\xi)\geq 2}\mathcal{W}(\xi)^{-s-1}+\sum_{y\in X_{c_{0}}\setminus\{x\},\atop\xi\in\mathcal{P}_{(\Delta,\iota)}(y\to X_{U}),\,\ell(\xi)\geq 1}\mathcal{W}_{\mathrm{rev}}(x,y)^{-s-1}\mathcal{W}(\xi)^{-s-1},$
where $\eta_{a}(u)=\mathbbm{1}_{\{o(a),t(a)\}}(u)$ if $u\in V\Gamma$ and $\eta_{a}(u)=\mathbbm{1}_{A}(u)$ if $u\in E\Gamma$.
Proof.
The statement is a direct consequence of (5.3), Proposition 4.7, Proposition 4.11 and Corollary 4.12. ∎
As in Section 5.4, we introduce a linear operator to express the series defining $\zeta_{G,G_{t_{1}},G_{t_{2}}}(s)$ within Setting [(P)-cl].
Definition 5.17.
Let $\Delta=(\Gamma,(X_{a}),(G(c)))$ be a local action diagram together with a function $\mathcal{W}\colon X\times X\to\mathbb{Z}_{\geq 0}$ (recall that $X=\bigsqcup_{a\in E\Gamma}X_{a}$). Let $\mathbb{C}\llbracket X\rrbracket$ be the complex vector space of all formal sums $\sum_{x\in X}\gamma_{x}x$, where $\gamma_{x}\in\mathbb{C}$ for every $x\in X$. For every non-empty set $S\subseteq X$, let
$f_{S}=\sum_{x\in S}x\in\mathbb{C}\llbracket X\rrbracket.$
Given $s\in\mathbb{C}$ and for all $x\in X_{a},y\in X_{b}$ with $a,b\in E\Gamma$, define
$\mathcal{F}(s)(x,y):=\left\{\begin{array}[]{cl}\begin{array}[]{cl}\mathcal{W}(x,y)^{-s},&\text{if }t(a)=o(b)\text{ and }y\neq\iota(x);\\ 0,&\text{otherwise.}\end{array}\end{array}\right.$
The Bass operator $\mathcal{F}(s)\colon\mathbb{C}\llbracket X\rrbracket\longrightarrow\mathbb{C}\llbracket X\rrbracket$ of $(\Delta,\mathcal{W})$ at $s\in\mathbb{C}$ is defined by linearly extending the following assignment, for all $x\in X$:
$\mathcal{F}(x):=\sum_{y\in X}\mathcal{F}(s)(x,y)y.$
Notation 5.18.
Technically, $\mathcal{F}(s)$ depends on $\Delta$ and $\mathcal{W}$. In our case, since $\Delta$ and $\mathcal{W}$ will be always clear from the context (in particular, $\mathcal{W}$ will be always the standard weight on $\Delta$), we avoid underlying this dependence.
In what follows, we implicitly fix a total order on $X$. Thus, we regard $\mathcal{F}(s)$ and the $f_{x}$’s as a $|X|$-dimensional matrix $[\mathcal{F}(s)(x,y)]_{x,y\in X}$ and as $|X|$-dimensional row vectors with complex entries, respectively. For all $x,y\in X$, note that $f_{x}\mathcal{F}(s)f_{y}^{t}=\mathcal{F}(s)(x,y)$.
Continuing the analogy with Section 4.2, we observe the following.
Remark 5.18.1.
Let $\mathcal{F}(s)$ be as in Definition 5.17. For every $n\geq 1$, let $\mathcal{F}(s)^{n}$ be the $n$-th power of $\mathcal{F}(s)$, and $\mathcal{F}(s)^{0}$ be the identity operator on $\mathbb{C}\llbracket X\rrbracket$. For $n\geq 0$ and for all $x,y\in X$, we claim that
(5.22) $f_{x}\cdot\mathcal{F}(s)^{n}\cdot f_{y}^{t}=\mathcal{F}(s)^{n}(x,y)=\sum_{\xi\in\mathcal{P}_{(\Delta,\iota)}(x\to y):\,\ell(\xi)=n+1}\mathcal{W}(\xi)^{-s},$
where $\mathcal{W}(\xi)=1$ if $\ell(\xi)=1$, and $\mathcal{W}(\xi)=\prod_{i=1}^{l-1}\mathcal{W}(x_{i},x_{i+1})$ if $\xi=(x_{1},\ldots,x_{l})$ for some $l\geq 2$. Indeed, (5.22) is immediate if $n\leq 1$. For $n\geq 2$, one argues as in Remark 5.11.1.
Theorem 5.19.
Let $G=U(\Delta,\iota,c_{0})$ and $T=T(\Delta,\iota,c_{0})$ be as in Setting [(P)-cl]. Let $t\in T$ with $\pi(t)=u$ and $e\in o^{-1}(v_{0})$ with $\mathcal{L}(e)=x$. Then, for every $r\in\{v_{0},e\}$, we have
$\zeta_{G,G_{r},G_{t}}(s)=\frac{\det(I-\mathcal{F}(s+1)+\mathcal{Y}_{\pi(r),u}(s+1))}{\det(I-\mathcal{F}(s+1))}+\kappa_{\pi(r)}(u),$
where $I$ is the identity matrix in $\mathrm{Mat}_{|X|}(\mathbb{C})$,
$\mathcal{Y}_{\pi(r),u}(s)=\left\{\begin{array}[]{ll}\sum_{a\in o^{-1}(c_{0})}\omega(a)^{-s}f_{X_{U}}^{t}f_{X_{a}},&\text{if }r=v_{0};\\ f_{X_{U}}^{t}\Big{(}f_{x}\mathcal{F}(s)+\sum_{y\in X_{c_{0}}\setminus\{x\}}\mathcal{W}_{\mathrm{rev}}(x,y)^{-s}f_{y}\Big{)},&\text{if }r=e;\end{array}\right.$
and
$\kappa_{\pi(r)}(u)=\left\{\begin{array}[]{cl}\mathbbm{1}_{\{c_{0}\}}(u)-1,&\text{if }r=v_{0};\\ \mathbbm{1}_{X_{A}}(x)-1,&\text{if }r=e.\end{array}\right.$
Proof.
One proceeds analogously as in the proof of Theorem 5.12. Let $s\in\mathbb{C}$ such that $\sum_{n=0}^{\infty}\mathcal{F}(s+1)$ converges. By Proposition 5.16 and Remark 5.18.1, we deduce what follows:
$\begin{split}&\zeta_{G,G_{v},G_{t}}(s)=\\ &=\left\{\begin{array}[]{ll}\displaystyle{\mathbbm{1}_{\{c_{0}\}}(u)+\sum_{a\in o^{-1}(c_{0})}\omega(a)^{-s-1}f_{X_{a}}\Bigg{(}\sum_{n=0}^{\infty}\mathcal{F}(s+1)^{n}\Bigg{)}f_{X_{U}}^{t},}&\text{if }u\in V\Gamma;\\ \displaystyle{\sum_{a\in o^{-1}(c_{0})}\omega(a)^{-s-1}f_{X_{a}}\Bigg{(}\sum_{n=0}^{\infty}\mathcal{F}(s+1)^{n}\Bigg{)}f_{X_{U}}^{t}},&\text{if }u\in E\Gamma;\end{array}\right.\\ &\zeta_{G,G_{e},G_{t}}(s)=\eta_{a}(u)+f_{x}\Bigg{(}\sum_{n=1}^{\infty}\mathcal{F}(s+1)^{n}\Bigg{)}f_{X_{U}}^{t}+\\ &\hskip 62.59596pt+\sum_{y\in X_{c_{0}}\setminus\{x\}}\mathcal{W}_{\mathrm{rev}}(x,y)^{-s-1}f_{y}\Bigg{(}\sum_{n=0}^{\infty}\mathcal{F}(s+1)^{n}\Bigg{)}f_{X_{U}}^{t}=\\ &=\mathbbm{1}_{X_{U}}(x)+\Big{(}f_{x}\mathcal{F}(s+1)+\sum_{y\in X_{c_{0}}\setminus\{x\}}\mathcal{W}_{\mathrm{rev}}(x,y)^{-s-1}f_{y}\Big{)}\Bigg{(}\sum_{n=0}^{\infty}\mathcal{F}(s+1)^{n}\Bigg{)}f_{X_{U}}^{t}.\end{split}$
Since $\sum_{n=0}^{\infty}\mathcal{F}(s+1)^{n}=(I-\mathcal{F}(s+1))^{-1}$, Fact 5.13 yields the claim. ∎
6. The reciprocal of $\mathcal{Z}_{\Gamma,u\to u}(s)$
In view of Section 7.2, we present some formulae involving the reciprocal of the function $\mathcal{Z}_{\Gamma,u\to u}(s)$, for $u\in\Gamma$, introduced in Definition 5.9. Recall that this function is a generalisation to weighted graphs of $\zeta_{G,G_{t},G_{t}}(s)$, where $(G,T)$ is a weakly locally $\infty$-transitive group action on a locally finite tree and $t\in T$ (cf. Remark 5.9.2).
Definition 6.1.
Let $\Gamma$ be a non-empty graph with an edge weight $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$. Consider $\mathcal{E}(s)$, for $s\in\mathbb{C}$, and $\{e_{u}\}_{u\in\Gamma}$ as in Definition 5.10. For all $c\in V\Gamma$ and $a\in E\Gamma$, define
$\mathcal{G}_{c}(s):=\mathcal{E}(s)-\mathcal{U}_{c,c}(s)\quad\text{and}\quad\mathcal{G}_{a}(s):=\mathcal{E}(s)-\mathcal{U}_{a,a}(s),$
where $\mathcal{U}_{c,c}(s)=\sum_{a\in o^{-1}(c)}\omega(a)^{-s}e_{c}^{t}e_{a}$ and $\mathcal{U}_{a,a}(s)=(e_{a}+e_{\bar{a}})^{t}\cdot(e_{a}+e_{\bar{a}})\mathcal{E}(s)$ (cf. Theorem 5.12).
Notation 6.2.
If necessary, we write $\mathcal{G}_{\bullet}^{\Gamma}(s)$, $\mathcal{E}^{\Gamma}(s)$, $\mathcal{U}_{\bullet,\bullet}^{\Gamma}(s)$, $e^{\Gamma}_{\bullet}$ and $I^{\Gamma}$ instead of $\mathcal{G}_{\bullet}(s)$, $\mathcal{E}(s)$, $\mathcal{U}_{\bullet,\bullet}(s)$, $e_{\bullet}$ and the identity matrix in $\mathrm{Mat}_{|E\Gamma|}(\mathbb{C})$, respectively.
Lemma 6.3.
Let $(\Gamma,\omega)$ satisfy Setting [$\Gamma$], and denote by $I$ the identity matrix on $\mathrm{Mat}_{|E\Gamma|}(\mathbb{C})$. Then, for all $c\in V\Gamma$ and $s\in\mathbb{C}$ such that $I-\mathcal{G}_{c}(s)$ is invertible, we have
$\mathcal{Z}_{\Gamma,c\to c}(s)^{-1}=1-\sum_{a\in o^{-1}(c)}\omega(a)^{-s}e_{a}(I-\mathcal{G}_{c}(s))^{-1}e_{c}^{t}.$
Moreover, for all $a\in E\Gamma$ and $s\in\mathbb{C}$ such that $I-\mathcal{G}_{a}(s)$ is invertible, we have
$\mathcal{Z}_{\Gamma,a\to a}(s)^{-1}=1-(e_{a}+e_{\bar{a}})\mathcal{E}(s)(I-\mathcal{G}_{a}(s))^{-1}(e_{a}+e_{\bar{a}})^{t}.$
Proof.
By Theorem 5.12, we deduce that
$\begin{array}[]{l}\displaystyle{\mathcal{Z}_{\Gamma,c\to c}(s)^{-1}=\frac{\det(I-\mathcal{G}_{c}(s)-\mathcal{U}_{c,c}(s))}{\det(I-\mathcal{G}_{c}(s))};}\\ \displaystyle{\mathcal{Z}_{\Gamma,a\to a}(s)^{-1}=\frac{\det(I-\mathcal{G}_{a}(s)-\mathcal{U}_{a,a}(s))}{\det(I-\mathcal{G}_{a}(s))}}.\end{array}$
The statements now follow from Fact 5.13. ∎
Proposition 6.4.
Let $(\Gamma,\omega)$ satisfy Setting [$\Gamma$]. Consider two subgraphs $\Gamma_{1}$ and $\Gamma_{2}$ of $\Gamma$ such that $\Gamma=\Gamma_{1}\cup\Gamma_{2}$ and $\Gamma_{1}\cap\Gamma_{2}=\{c\}$, for some $c\in V\Gamma$. Then,
$\mathcal{Z}_{\Gamma,c\to c}(s)^{-1}=\mathcal{Z}_{\Gamma_{1},c\to c}(s)^{-1}+\mathcal{Z}_{\Gamma_{2},c\to c}(s)^{-1}-1.$
Proof.
Let $s\in\mathbb{C}$ such that $I^{\Gamma}-\mathcal{G}_{c}^{\Gamma}(s)$ is invertible, and set $\Gamma_{3}=\Gamma_{1}\cap\Gamma_{2}$. By Definition 6.1, for all $a,b\in E\Gamma$ we have
(6.1) $\begin{split}(I^{\Gamma}-\mathcal{G}_{c}^{\Gamma}(s))(a,b)&=e_{a}^{\Gamma}\cdot\big{(}I^{\Gamma}-\mathcal{G}_{c}^{\Gamma}(s)\big{)}\cdot(e_{b}^{\Gamma})^{t}\\ &=(I^{\Gamma}-\mathcal{E}^{\Gamma}(s))(a,b)+\sum_{a^{\prime}\in o^{-1}(c)}\omega(a^{\prime})^{-s}(e_{a}^{\Gamma}(e_{c}^{\Gamma})^{t})\cdot(e_{a^{\prime}}^{\Gamma}(e_{b}^{\Gamma})^{t})\\ &=(I^{\Gamma}-\mathcal{E}^{\Gamma}(s))(a,b)+\mathbbm{1}_{t^{-1}(c)}(a)\mathbbm{1}_{o^{-1}(c)}(b)\omega(b)^{-s}.\end{split}$
Similarly, for $1\leq i\leq 3$ and for all $a,b\in E\Gamma_{i}$ we have
(6.2) $(I^{\Gamma_{i}}-\mathcal{G}_{c}^{\Gamma_{i}}(s))(a,b)=(I^{\Gamma_{i}}-\mathcal{E}^{\Gamma_{i}}(s))(a,b)+\mathbbm{1}_{t^{-1}(c)\cap E\Gamma_{i}}(a)\mathbbm{1}_{o^{-1}(c)\cap E\Gamma_{i}}(b)\omega(b)^{-s}.$
Combining (6.1) and (6.2), for every $1\leq i\leq 3$ we deduce that
(6.3) $(I^{\Gamma}-\mathcal{G}_{c}^{\Gamma}(s))(a,b)=(I^{\Gamma_{i}}-\mathcal{G}_{c}^{\Gamma_{i}})(a,b),\quad\forall\,a,b\in E\Gamma_{i}$
and Lemma 6.3 implies that
(6.4) $\mathcal{Z}_{\Gamma_{i},c\to c}(s)^{-1}=1-\sum_{a\in o^{-1}(c)\cap E\Gamma_{i},\atop b\in t^{-1}(c)\cap E\Gamma_{i}}\omega(a)^{-s}(I_{\Gamma}-\mathcal{G}_{c}^{\Gamma}(s))^{-1}(a,b).$
Moreover, for all $a,b\in E\Gamma$ with $t(a)=c=o(b)$ we have $(I^{\Gamma}-\mathcal{G}_{c}^{\Gamma}(s))(a,b)=\mathbbm{1}_{\{a\}}(b)-(\omega(b)-\mathbbm{1}_{\{\bar{a}\}}(b))^{-s}+\omega(b)^{-s}$ and then
(6.5) $(I^{\Gamma}-\mathcal{G}_{c}^{\Gamma}(s))(a,b)=0,\quad\forall\,b\in o^{-1}(c)\setminus\{a,\bar{a}\}.$
We claim that
$(I^{\Gamma}-\mathcal{G}_{c}^{\Gamma}(s))(a,b)=0,\quad\forall\,(a,b)\in(E\Gamma_{1}\times E\Gamma_{2})\cup(E\Gamma_{2}\times E\Gamma_{1}).$
Indeed, recall that $E\Gamma_{1}\cap E\Gamma_{2}=\emptyset$ and $V\Gamma_{1}\cap V\Gamma_{2}=\{c\}$. Hence, for all $a\in E\Gamma_{1}$ and $b\in E\Gamma_{2}$, we have $b\not\in\{a,\bar{a}\}$ and either $t(a)\neq o(b)$ or $t(a)=c=o(b)$. Now (6.1) and (6.5) apply. A similar argument holds for all $a\in E\Gamma_{2}$ and $b\in E\Gamma_{1}$.
Therefore, once fixed a total order $\leq$ on $E\Gamma$ so that $a<b$ for all $a\in E\Gamma_{1}$ and $b\in E\Gamma_{2}$, we have the following decomposition in diagonal blocks:
(6.6) $I^{\Gamma}-\mathcal{G}_{c}^{\Gamma}(s)=\begin{bmatrix}I^{\Gamma_{1}}-\mathcal{G}_{c}^{\Gamma_{1}}(s)&0\\ 0&I^{\Gamma_{2}}-\mathcal{G}_{c}^{\Gamma_{2}}(s)\end{bmatrix}.$
Since
$o^{-1}(c)=(o^{-1}(c)\cap E\Gamma_{1})\sqcup(o^{-1}(c)\cap E\Gamma_{2}),$
by Lemma 6.3, (6.6) and then  (6.4), we conclude that
$\begin{array}[]{ccl}\mathcal{Z}_{\Gamma,c\to c}(s)^{-1}&=&\displaystyle{1-\sum_{a,b\in o^{-1}(c)}\omega(a)^{-s}(I^{\Gamma}-\mathcal{G}_{c}^{\Gamma}(s))^{-1}(a,\bar{b})}\\ &=&\displaystyle{1-\sum_{a,b\in o^{-1}(c)\cap E\Gamma_{1}}\omega(a)^{-s}(I^{\Gamma_{1}}-\mathcal{G}_{c}^{\Gamma_{1}}(s))^{-1}(a,\bar{b})+}\\ &&\displaystyle{-\sum_{a,b\in o^{-1}(c)\cap E\Gamma_{2}}\omega(a)^{-s}(I^{\Gamma_{2}}-\mathcal{G}_{c}^{\Gamma_{2}}(s))^{-1}(a,\bar{b})}\\ &=&\mathcal{Z}_{\Gamma_{1},c\to c}(s)^{-1}+\mathcal{Z}_{\Gamma_{2},c\to c}(s)^{-1}-1.\qed\end{array}$
Corollary 6.5.
Let $(\Gamma,\omega)$ satisfy Setting [$\Gamma$]. Assume that there are subgraphs $\Lambda_{1}$ and $\Lambda_{2}$ of $\Gamma$ such that $\Gamma=\Lambda_{1}\cup\Lambda_{2}$ and $\Lambda_{1}\cap\Lambda_{2}=\{c\}$, for some vertex $c\in V\Gamma$. Then, for all subgraphs $\Gamma_{1}$ and $\Gamma_{2}$ of $\Gamma$ satisfying $\Gamma_{i}\supseteq\Lambda_{i}$ for every $i\in\{1,2\}$, we have
$\mathcal{Z}_{\Gamma,c\to c}(s)^{-1}=\mathcal{Z}_{\Gamma_{1},c\to c}(s)^{-1}+\mathcal{Z}_{\Gamma_{2},c\to c}(s)^{-1}-\mathcal{Z}_{\Gamma_{1}\cap\Gamma_{2},c\to c}(s)^{-1}.$
Proof.
Let $\Gamma_{1}$ and $\Gamma_{2}$ be as in the statement, and set $\Gamma_{3}:=\Gamma_{1}\cap\Gamma_{2}$. Note that $\Lambda_{1}\cap\Gamma_{3}=\Lambda_{1}\cap\Gamma_{2}$ and $\Lambda_{2}\cap\Gamma_{3}=\Lambda_{2}\cap\Gamma_{1}$. Therefore,
(6.11) $\begin{array}[]{lcl}\Gamma=\Lambda_{1}\cup\Lambda_{2}&\text{and}&\Lambda_{1}\cap\Lambda_{2}=\{c\};\\ \Gamma_{1}=\Lambda_{1}\cup(\Lambda_{2}\cap\Gamma_{1})&\text{and}&\Lambda_{1}\cap(\Lambda_{2}\cap\Gamma_{1})=\{c\};\\ \Gamma_{2}=(\Lambda_{1}\cap\Gamma_{2})\cup\Lambda_{2}&\text{and}&(\Lambda_{1}\cap\Gamma_{2})\cap\Lambda_{2}=\{c\};\\ \Gamma_{3}=(\Lambda_{1}\cap\Gamma_{2})\cup(\Lambda_{2}\cap\Gamma_{1})&\text{and}&(\Lambda_{1}\cap\Gamma_{2})\cap(\Lambda_{2}\cap\Gamma_{1})=\{c\}.\end{array}$
Applying Proposition 6.4 to each decomposition in (6.11) yields the claim. ∎
The hypotheses of Proposition 6.4 are satisfied for every $c\in V\Gamma$ if, for instance, $\Gamma$ is a connected graph without $n$-cycles, for every $n\geq 2$.
Lemma 6.6.
Let $\Gamma$ be a connected graph without $n$-cycles, for every $n\geq 2$. For every $c\in V\Gamma$, there are connected subgraphs $\Gamma_{1}$ and $\Gamma_{2}$ of $\Gamma$ such that $\Gamma=\Gamma_{1}\cup\Gamma_{2}$ and $\Gamma_{1}\cap\Gamma_{2}=\{c\}$. If in particular $|o^{-1}(c)|\geq 2$, one may take $\Gamma_{1}$ and $\Gamma_{2}$ to be proper subgraphs of $\Gamma$.
Proof.
Let $\{\Xi_{i}\}_{i\in I}$ be the collection of all connected components of the graph $\Gamma\setminus(o^{-1}(c)\cup\overline{o^{-1}(c)}\cup\{c\})$. Recall that $\Gamma\setminus(o^{-1}(c)\cup\overline{o^{-1}(c)}\cup\{c\})=\bigcup_{i\in I}\Xi_{i}$ and the union is disjoint. For every $i\in I$, there is exactly one edge $a_{i}\in o^{-1}(c)$ such that $t(a_{i})\in\Xi_{i}$. In fact, assume that there are $a,b\in o^{-1}(c)$ with $a\neq b$ and $x=t(a),y=t(b)\in V\Xi_{i}$. Then the reduced path $[x,y]$ as in Remark 2.1.1 is contained in $\Xi_{i}$. Therefore, $a\cdot[x,y]\cdot\bar{b}$ is a cycle of length $\geq 2$ in $\Gamma$, impossible.
Consider subsets $\mathcal{E}_{1},\mathcal{E}_{2}$ of $o^{-1}(c)$ with the following properties: $\mathcal{E}_{1}\cap\mathcal{E}_{2}=\emptyset$, $o^{-1}(c)=\mathcal{E}_{1}\cup\mathcal{E}_{2}$ and, for every $k\in\{1,2\}$, every $1$-loop $a$ in $\mathcal{E}_{k}$ satisfies that $\bar{a}\in\mathcal{E}_{k}$. Provided ¯$\overline{\mathcal{E}}_{k}=\{\bar{a}\mid a\in\mathcal{E}_{k}\}$ for every $k\in\{1,2\}$, note that $(\mathcal{E}_{1}\cup\overline{\mathcal{E}}_{1})\cap(\mathcal{E}_{2}\cup\overline{\mathcal{E}}_{2})=\emptyset$. Moreover, if $|o^{-1}(c)|\geq 2$, take $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ so that $\mathcal{E}_{1}\neq\emptyset$ and $\mathcal{E}_{2}\neq\emptyset$. For $k\in\{1,2\}$, set also $I_{k}:=\{i\in I\mid a_{i}\in\mathcal{E}_{k}\}$. Note that $I_{1}\cap I_{2}=\emptyset$ and $I=I_{1}\cup I_{2}$. For every $k\in\{1,2\}$, define the following subgraph of $\Gamma$:
$\Gamma_{k}:=\{c\}\cup\mathcal{E}_{k}\cup\overline{\mathcal{E}}_{k}\cup\bigcup_{i\in I_{k}}\Xi_{i}.$
One checks that $\Gamma=\Gamma_{1}\cup\Gamma_{2}$ and $\Gamma_{1}\cap\Gamma_{2}=\{c\}$. If in particular $|o^{-1}(c)|\geq 2$, then $\Gamma_{1}\setminus\Gamma_{2}\supseteq\mathcal{E}_{1}\neq\emptyset$ and $\Gamma_{2}\setminus\Gamma_{1}\supseteq\mathcal{E}_{2}\neq\emptyset$. Therefore, both $\Gamma_{1}$ and $\Gamma_{2}$ are proper subgraphs of $\Gamma$. ∎
Proposition 6.7.
Let $(\Gamma,\omega)$ satisfy Setting [$\Gamma$]. Consider subgraphs $\Gamma_{1}$ and $\Gamma_{2}$ of $\Gamma$ satisfying $\Gamma=\Gamma_{1}\cup\Gamma_{2}$ and such that $\Gamma_{3}:=\Gamma_{1}\cap\Gamma_{2}$ is a $1$-segment with edge set $\{a,\bar{a}\}$. Then,
$\mathcal{Z}_{\Gamma,a\to a}(s)^{-1}=\mathcal{Z}_{\Gamma_{1},a\to a}(s)^{-1}+\mathcal{Z}_{\Gamma_{2},a\to a}(s)^{-1}-\mathcal{Z}_{\Gamma_{3},a\to a}(s)^{-1}.$
Proof.
Fix $s\in\mathbb{C}$ such that $I^{\Gamma}-\mathcal{G}_{a}^{\Gamma}(s)$ is invertible, and set $c=o(a)$, $d=t(a)$. By Lemma 6.3, we have
(6.12) $\begin{array}[]{c}\mathcal{Z}_{\Gamma,a\to a}(s)^{-1}=\\ \displaystyle{=1-\Bigg{(}\sum_{b\in E\Gamma,\atop o(b)=d}\mathcal{E}^{\Gamma}(s)(a,b)e_{b}^{\Gamma}+\sum_{b\in E\Gamma,\atop o(b)=c}\mathcal{E}^{\Gamma}(s)(\bar{a},b)e_{b}^{\Gamma}\Bigg{)}(I^{\Gamma}-\mathcal{G}_{a}^{\Gamma}(s))^{-1}(e_{a}^{\Gamma}+e_{\bar{a}}^{\Gamma})^{t}.}\end{array}$
By Definition 6.1, for all $b_{1},b_{2}\in E\Gamma$ we observe that
(6.17) $\begin{array}[]{c}(I^{\Gamma}-\mathcal{G}_{a}^{\Gamma}(s))(b_{1},b_{2})=\\ \begin{array}[]{rl}=&\mathbbm{1}_{\{b_{1}\}}(b_{2})-\mathcal{E}^{\Gamma}(s)(b_{1},b_{2})+e_{b_{1}}^{\Gamma}(e_{a}^{\Gamma}+e_{\bar{a}}^{\Gamma})^{t}(e_{a}^{\Gamma}+e_{\bar{a}}^{\Gamma})\mathcal{E}^{\Gamma}(s)(e_{b_{2}}^{\Gamma})^{t}\\ =&\mathbbm{1}_{\{b_{1}\}}(b_{2})-\mathcal{E}^{\Gamma}(s)(b_{1},b_{2})+\mathbbm{1}_{\{a,\bar{a}\}}(b_{1})\cdot\Big{(}\mathcal{E}^{\Gamma}(s)(a,b_{2})+\mathcal{E}^{\Gamma}(s)(\bar{a},b_{2})\Big{)}.\end{array}\end{array}$
Hence, for every $i\in\{1,2\}$,
(6.18) $\begin{split}(I^{\Gamma_{i}}-\mathcal{G}_{a}^{\Gamma_{i}}(s))(b_{1},b_{2})&=(I^{\Gamma}-\mathcal{G}_{a}^{\Gamma}(s))(b_{1},b_{2}),\quad\forall\,b_{1},b_{2}\in E\Gamma_{i};\\ (I^{\Gamma_{i}}-\mathcal{G}_{a}^{\Gamma_{i}}(s))(b_{1},b_{2})&=(I^{\Gamma_{3}}-\mathcal{G}_{a}^{\Gamma_{3}}(s))(b_{1},b_{2}),\quad\forall\,b_{1},b_{2}\in\{a,\bar{a}\}.\end{split}$
Let $\Lambda_{1}$ (resp. $\Lambda_{2}$) be the graph obtained from $\Gamma_{1}$ (resp. $\Gamma_{2}$) by removing $a,\bar{a}$ and $d$ (resp. $a,\bar{a}$ and $c$). The following picture sketches an example of $\Lambda_{1}$ (with wavy edges) and $\Lambda_{2}$ (with dashed edges).
[…]
Note that $c\in V\Lambda_{1}$, $d\in V\Lambda_{2}$ and $V\Lambda_{1}\cap V\Lambda_{2}=\emptyset$. In particular, no edges of $\Lambda_{1}$ end in a vertex of $\Lambda_{2}$. We claim that
(6.19) $\begin{array}[]{c}(I^{\Gamma}-\mathcal{G}_{a}^{\Gamma}(s))(b_{1},b_{2})=0,\quad\forall\,(b_{1},b_{2})\in(E\Lambda_{1}\sqcup\{a\})\times(E\Lambda_{2}\sqcup\{\bar{a}\});\\ (I^{\Gamma}-\mathcal{G}_{a}^{\Gamma}(s))(b_{1},b_{2})=0,\quad\forall\,(b_{1},b_{2})\in(E\Lambda_{2}\sqcup\{\bar{a}\})\times(E\Lambda_{1}\sqcup\{a\}).\end{array}$
Indeed, let $b_{1}\in E\Lambda_{1}\sqcup\{a\}$ and $b_{2}\in E\Lambda_{2}\sqcup\{\bar{a}\}$. If $b_{1}\in E\Lambda_{1}$, then $V\Lambda_{1}\ni t(b_{1})\neq o(b_{2})\in V\Lambda_{2}$ and (6.17) implies that $I^{\Gamma}-\mathcal{G}^{\Gamma}_{a}(s)=-\mathcal{E}^{\Gamma}(b_{1},b_{2})=0$. If $b_{1}=a$, then $c=t(\bar{a})\neq o(b_{2})\in E\Lambda_{2}$ and (6.17) implies that $I^{\Gamma}-\mathcal{G}_{a}^{\Gamma}(s)=\mathcal{E}^{\Gamma}(s)(\bar{a},b_{2})=0$. The second line of (6.19) can be proved analogously.
Fix a total order $\leq$ on $E\Gamma=E\Lambda_{1}\sqcup\{a,\bar{a}\}\sqcup E\Lambda_{2}$ so that $b_{1}<a<\bar{a}<b_{2}$ for all $b_{1}\in E\Lambda_{1}$ and $b_{2}\in E\Lambda_{2}$. Set also
$\begin{split}A&:=[(I^{\Gamma_{1}}-\mathcal{G}_{a}^{\Gamma_{1}}(s))(b_{1},b_{2})]_{b_{1},b_{2}\in E\Lambda_{1}\sqcup\{a\}};\\ B&:=[(I^{\Gamma_{2}}-\mathcal{G}_{a}^{\Gamma_{2}}(s))(b_{1},b_{2})]_{b_{1},b_{2}\in\{\bar{a}\}\sqcup E\Lambda_{2}};\\ \alpha&:=\omega(a)-1\quad\text{and}\quad\beta:=\omega(\bar{a})-1.\end{split}$
From (6.18) we observe that $A(a,a)=1+\alpha^{-s}$ and $B(\bar{a},\bar{a})=1+\beta^{-s}$. Moreover, by (6.18) and (6.19), we have the following decompositions in diagonal blocks:
(6.20) $\begin{split}I^{\Gamma}-\mathcal{G}_{a}^{\Gamma}(s)=\begin{bmatrix}A&0\\ 0&B\end{bmatrix}&\quad\text{and}\quad(I^{\Gamma}-\mathcal{G}_{a}^{\Gamma}(s))^{-1}=\begin{bmatrix}A^{-1}&0\\ 0&B^{-1}\end{bmatrix};\\ I^{\Gamma_{1}}-\mathcal{G}_{a}^{\Gamma_{1}}(s)=\begin{bmatrix}A&0\\ 0&B(\bar{a},\bar{a})\end{bmatrix}&\quad\text{and}\quad(I^{\Gamma_{1}}-\mathcal{G}_{a}^{\Gamma_{1}}(s))^{-1}=\begin{bmatrix}A^{-1}&0\\ 0&B(\bar{a},\bar{a})^{-1}\end{bmatrix};\\ I^{\Gamma_{2}}-\mathcal{G}_{a}^{\Gamma_{2}}(s)=\begin{bmatrix}A(a,a)&0\\ 0&B\end{bmatrix}&\quad\text{and}\quad(I^{\Gamma_{2}}-\mathcal{G}_{a}^{\Gamma_{2}}(s))^{-1}=\begin{bmatrix}A(a,a)^{-1}&0\\ 0&B^{-1}\end{bmatrix}.\end{split}$
Note that
(6.21) $o^{-1}(c)=\{a\}\sqcup(o^{-1}(c)\cap E\Lambda_{1})\,\text{ and }\,o^{-1}(d)=\{\bar{a}\}\sqcup(o^{-1}(d)\cap E\Lambda_{2}).$
Therefore, by (6.20) and (6.21), we rewrite (6.12) as follows:
(6.22) $\begin{split}\mathcal{Z}_{\Gamma,a\to a}(s)^{-1}&=1-\beta^{-s}B^{-1}(\bar{a},\bar{a})-\sum_{b\in o^{-1}(d)\cap E\Lambda_{2}}\mathcal{E}^{\Gamma_{2}}(s)(a,b)\cdot B^{-1}(b,\bar{a})+\\ &\quad-a^{-s}A^{-1}(a,a)-\sum_{b\in o^{-1}(c)\cap E\Lambda_{1}}\mathcal{E}^{\Gamma_{1}}(s)(\bar{a},b)\cdot A^{-1}(b,a).\end{split}$
A formula analogous to (6.12) holds for $\Gamma_{1}$, namely
$\mathcal{Z}_{\Gamma_{1},a\to a}(s)^{-1}=$
$1-\Bigg{(}\mathcal{E}^{\Gamma_{1}}(s)(a,\bar{a})e_{\bar{a}}^{\Gamma_{1}}+\mathcal{E}^{\Gamma_{1}}(s)(\bar{a},a)e_{a}^{\Gamma_{1}}+\sum_{b\in o^{-1}(c)\cap E\Lambda_{1}}\mathcal{E}^{\Gamma_{1}}(s)(\bar{a},b)e_{b}^{\Gamma_{1}}\Bigg{)}\cdot$
$\cdot\Big{(}I^{\Gamma_{1}}-\mathcal{G}_{a}^{\Gamma_{1}}(s)\Big{)}^{-1}\cdot(e_{a}^{\Gamma_{1}}+e_{\bar{a}}^{\Gamma_{1}})^{t}$
which, by (6.20), yields
(6.23) $\mathcal{Z}_{\Gamma_{1},a\to a}(s)^{-1}=1-\beta^{-s}B(\bar{a},\bar{a})^{-1}-\alpha^{-s}A^{-1}(a,a)-\sum_{b\in o^{-1}(c)\cap E\Lambda_{1}}\mathcal{E}^{\Gamma_{1}}(s)(\bar{a},b)\cdot A^{-1}(b,a).$
In a similar manner, we deduce that
(6.24) $\mathcal{Z}_{\Gamma_{2},a\to a}(s)^{-1}=1-\alpha^{-s}A(a,a)^{-1}-\beta^{-s}B^{-1}(\bar{a},\bar{a})-\sum_{b\in o^{-1}(d)\cap E\Lambda_{2}}\mathcal{E}^{\Gamma_{2}}(s)(a,b)\cdot B^{-1}(b,\bar{a}).$
By Example 5.14, one also checks that
(6.25) $\mathcal{Z}_{\Gamma_{3},a\to a}(s)^{-1}=\frac{1-\alpha^{-s}\beta^{-s}}{(1+\alpha^{-s})(1+\beta^{-s})}=1-\alpha^{-s}A(a,a)^{-1}-\beta^{-s}B(\bar{a},\bar{a})^{-1}.$
Combining (6.22), (6.23), (6.24) and (6.25), we conclude the claim. ∎
Note that the strategy to prove Proposition 6.7 strictly depends on the fact that $a$ has distinct endpoints (cf. (6.19)).
Corollary 6.8.
Let $(\Gamma,\omega)$ satisfy Setting [$\Gamma$]. Assume that there are subgraphs $\Lambda_{1}$ and $\Lambda_{2}$ of $\Gamma$ satisfying $\Gamma=\Lambda_{1}\cup\Lambda_{2}$ and such that $\Lambda_{1}\cap\Lambda_{2}$ is a $1$-segment graph with edge set $\{a,\bar{a}\}$. Then, for all subgraphs $\Gamma_{1}$ and $\Gamma_{2}$ of $\Gamma$ such that $\Gamma_{i}\supseteq\Lambda_{i}$ for every $i\in\{1,2\}$, we have
$\mathcal{Z}_{\Gamma,a\to a}(s)^{-1}=\mathcal{Z}_{\Gamma_{1},a\to a}(s)^{-1}+\mathcal{Z}_{\Gamma_{2},a\to a}(s)^{-1}-\mathcal{Z}_{\Gamma_{1}\cap\Gamma_{2},a\to a}(s)^{-1}.$
Proof.
Let $\Gamma_{1}$ and $\Gamma_{2}$ as in the statement. For simplicity, set $\Gamma_{3}=\Gamma_{1}\cap\Gamma_{2}$ and $\Gamma_{a}=\Lambda_{1}\cap\Lambda_{2}$. Since $\Lambda_{1}\subseteq\Gamma_{1}$ and $\Lambda_{2}\subseteq\Gamma_{2}$, we have $\Lambda_{1}\cap\Gamma_{2}=\Lambda_{1}\cap\Gamma_{3}$ and $\Lambda_{2}\cap\Gamma_{1}=\Lambda_{2}\cap\Gamma_{3}$. Therefore, the following decompositions hold:
(6.30) $\begin{array}[]{lcl}\Gamma=\Lambda_{1}\cup\Lambda_{2}&\text{and}&\Lambda_{1}\cap\Lambda_{2}=\Gamma_{a};\\ \Gamma_{1}=\Lambda_{1}\cup(\Lambda_{2}\cap\Gamma_{1})&\text{and}&\Lambda_{1}\cap(\Lambda_{2}\cap\Gamma_{1})=\Gamma_{a};\\ \Gamma_{2}=(\Lambda_{1}\cap\Gamma_{2})\cup\Lambda_{2}&\text{and}&(\Lambda_{1}\cap\Gamma_{2})\cap\Lambda_{2}=\Gamma_{a};\\ \Gamma_{3}=(\Lambda_{1}\cap\Gamma_{2})\cup(\Lambda_{2}\cap\Gamma_{1})&\text{and}&(\Lambda_{1}\cap\Gamma_{2})\cap(\Lambda_{2}\cap\Gamma_{1})=\Gamma_{a}.\end{array}$
Applying Proposition 6.7 to each decomposition in (6.30) yields the claim. ∎
For completeness, in analogy to Lemma 6.6 we observe the following.
Remark 6.8.1.
Let $\Gamma$ be a connected graph without $n$-cycles for every $n\geq 2$, and let $a\in E\Gamma$. By Lemma 6.6, there are connected subgraphs $\Lambda_{1}$ and $\Lambda_{2}$ of $\Gamma$ such that $\Gamma=\Lambda_{1}\cup\Lambda_{2}$ and $\Lambda_{1}\cap\Lambda_{2}=\{o(a)\}$. Denote by $\Gamma_{1}$ and $\Gamma_{2}$ the smallest subgraphs of $\Gamma$ containing $\Lambda_{1}\cup\{a,\bar{a}\}$ and $\Lambda_{2}\cup\{a,\bar{a}\}$, respectively. Then $\Gamma=\Gamma_{1}\cup\Gamma_{2}$ and $\Gamma_{1}\cap\Gamma_{2}$ is the subgraph of $\Gamma$ with edge set $\{a,\bar{a}\}$.
In view of the next proofs, it might be useful to recall the following well-known fact [24]. Given a $2\times 2$ block-matrix $M=[M_{ij}]_{1\leq i,j\leq 2}\in\mathrm{Mat}_{n}(\mathbb{C})$ with $M_{22}$ invertible, one has
(6.31) $\det(M)=\det(M_{22})\cdot\det(M_{11}-M_{12}M_{22}^{-1}M_{21}).$
Proposition 6.9.
Let $(\Gamma,\omega)$ satisfy Setting [$\Gamma$]. Consider $a\in E\Gamma$ with $o(a)=:c\neq d:=t(a)$, and assume that $c$ is a terminal vertex in $\Gamma$. Put $\omega(a)=\alpha+1$, $\omega(\bar{a})=\beta+1$, and denote by $\Lambda$ the graph obtained from $\Gamma$ by removing $a$, $\bar{a}$ and $c$. Then,
$\begin{split}\mathcal{Z}_{\Gamma,c\to c}(s)^{-1}&=\frac{(1+\alpha^{-s}\xi(\beta,s))\cdot\mathcal{Z}_{\Lambda,d\to d}(s)^{-1}-\alpha^{-s}(\beta+1)^{-s}}{(1-\xi(\alpha,s)\xi(\beta,s))\cdot\mathcal{Z}_{\Lambda,d\to d}(s)^{-1}+\xi(\alpha,s)(\beta+1)^{-s}};\\ \mathcal{Z}_{\Gamma,a\to a}(s)^{-1}&=\frac{(1+\alpha^{-s}\xi(\beta,s))\cdot\mathcal{Z}_{\Lambda,d\to d}(s)^{-1}-\alpha^{-s}(\beta+1)^{-s}}{(1+\alpha^{-s})\Big{(}(1-\xi(\beta,s))\cdot\mathcal{Z}_{\Lambda,d\to d}(s)^{-1}+(\beta+1)^{-s}\Big{)}},\end{split}$
where $\xi(\alpha,s)=(\alpha+1)^{-s}-\alpha^{-s}$ and $\xi(\beta,s)=(\beta+1)^{-s}-\beta^{-s}$.
The picture below sketches a possible setting for Proposition 6.9. The edges of $\Lambda$ are dashed.
[…]
Proof.
Let $s\in\mathbb{C}$ such that $\mathcal{Z}_{\Gamma,a\to a}(s)\neq 0$, and consider a total order $\leq$ on $E\Gamma$ such that $a<\bar{a}<b$ for all $b\in E\Lambda$. Then $\mathcal{E}^{\Gamma}(s)$ admits the following block decomposition:
(6.32) $\mathcal{E}^{\Gamma}(s)=\begin{bmatrix}A&B\\ C&D\end{bmatrix},$
where $A=[\mathcal{E}^{\Gamma}(s)(b_{1},b_{2})]_{b_{1},b_{2}\in\{a,\bar{a}\}}$, $B=[\mathcal{E}^{\Gamma}(s)(b_{1},b_{2})]_{b_{1}\in\{a,\bar{a}\},\,b_{2}\in E\Lambda}$, $C=[\mathcal{E}^{\Gamma}(s)(b_{1},b_{2})]_{b_{1}\in E\Lambda,\,b_{2}\in\{a,\bar{a}\}}$ and $D=[\mathcal{E}^{\Gamma}(s)(b_{1},b_{2})]_{b_{1},b_{2}\in E\Lambda}=\mathcal{E}^{\Lambda}(s)$. For $i\in\{1,2\}$, let $B_{i}$ and $C^{i}$ denote the $i$-th row of $B$ and the $i$-th column of $C$, respectively. One checks that
(6.33) $B_{1}=\sum_{b\in o^{-1}(d)\cap E\Lambda}\omega(b)^{-s}e_{b}^{\Lambda};\quad B_{2}=\underline{0};\quad C^{1}=\underline{0}^{t}\quad\text{and}\quad C^{2}=(\beta+1)^{-s}(e_{d}^{\Lambda})^{t},$
where $\underline{0}$ denotes the row zero vector in $\mathbb{C}^{|E\Lambda|}$. Moreover, denote by $E_{21}$ and $\underline{1}\in\mathrm{Mat}_{2}(\mathbb{C})$ the elementary matrix associated to $(2,1)$ and the matrix with all the entries equal to $1$, respectively. Hence,
$\begin{split}\mathcal{U}_{cc}^{\Gamma}(s)&=(\alpha+1)^{-1}(e_{\bar{a}}^{\Gamma})^{t}(e_{a}^{\Gamma})=\begin{bmatrix}(\alpha+1)^{-s}E_{21}&0_{2\times|E\Lambda|}\\ 0_{|E\Lambda|\times 2}&0_{|E\Lambda|\times|E\Lambda|}\end{bmatrix};\\ \mathcal{U}_{aa}^{\Gamma}(s)&=(e_{a}^{\Gamma}+e_{\bar{a}}^{\Gamma})^{t}(e_{a}^{\Gamma}+e_{\bar{a}}^{\Gamma})\mathcal{E}^{\Gamma}(s)=\begin{bmatrix}\underline{1}\cdot A&\underline{1}\cdot B\\ 0_{|E\Lambda|\times 2}&0_{|E\Lambda|\times|E\Lambda|}\end{bmatrix}.\end{split}$
Denoting by $I_{2}$ the identity matrix in $\mathrm{Mat}_{2}(\mathbb{C})$, the following holds:
(6.34) $\begin{split}I^{\Gamma}-\mathcal{E}^{\Gamma}(s)&=\begin{bmatrix}I_{2}-A&-B\\ -C&I_{\Lambda}-D\end{bmatrix};\\ I^{\Gamma}-\mathcal{E}^{\Gamma}(s)+\mathcal{U}_{c,c}^{\Gamma}(s)&=\begin{bmatrix}I_{2}-A+(\alpha+1)^{-s}E_{21}&-B\\ -C&I_{\Lambda}-D\end{bmatrix};\\ I^{\Gamma}-\mathcal{E}^{\Gamma}(s)+\mathcal{U}_{a,a}^{\Gamma}(s)&=\begin{bmatrix}I_{2}+(\underline{1}-I_{2})A&(\underline{1}-I_{2})B\\ -C&I_{\Lambda}-D\end{bmatrix};\\ I_{2}-A=\begin{bmatrix}1&-\beta^{-s}\\ -\alpha^{-s}&1\end{bmatrix};&\qquad I_{2}+(\underline{1}-I_{2})A=\begin{bmatrix}1+\alpha^{-s}&0\\ 0&1+\beta^{-s}\end{bmatrix}.\end{split}$
By Theorem 5.12, (6.31) and (6.34), we deduce that
(6.35) $\begin{split}\mathcal{Z}_{\Gamma,c\to c}(s)^{-1}&=\frac{\det(I^{\Gamma}-\mathcal{E}^{\Gamma}(s))}{\det(I-\mathcal{E}^{\Gamma}(s)+\mathcal{U}_{c,c}^{\Gamma}(s))}\\ &=\frac{\det\Big{(}I_{2}-A-\overbrace{B(I_{\Lambda}-D)^{-1}C}^{=:X}\Big{)}}{\det\Big{(}I_{2}-A+(\alpha+1)^{-s}E_{21}-\underbrace{B(I_{\Lambda}-D)^{-1}C}_{=X}\Big{)}}\\ &=\frac{(1-X(a,a))(1-X(\bar{a},\bar{a}))-(X(a,\bar{a})+\beta^{-s})(X(\bar{a},a)+\alpha^{-s})}{(1-X(a,a))(1-X(\bar{a},\bar{a}))+(\xi(\alpha,s)-X(\bar{a},a))(\beta^{-s}+X(a,\bar{a}))}.\end{split}$
Similarly,
(6.36) $\begin{split}\mathcal{Z}_{\Gamma,a\to a}(s)^{-1}&=\frac{\det(I^{\Gamma}-D)}{\det(I-D+\mathcal{U}_{a,a}^{\Gamma}(s))}\\ &=\frac{\det\Big{(}I_{2}-A-\overbrace{B(I_{\Lambda}-D)^{-1}C}^{=X}\Big{)}}{\det\Big{(}I_{2}+(\underline{1}-I_{2})A+\underbrace{(\underline{1}-I_{2})B(I_{\Lambda}-D)^{-1}C}_{=:Y}\Big{)}}\\ &=\frac{(1-X(a,a))(1-X(\bar{a},\bar{a}))-(X(a,\bar{a})+\beta^{-s})(X(\bar{a},a)+\alpha^{-s})}{(1+\alpha^{-s}+Y(a,a))(1+\beta^{-s}+Y(\bar{a},\bar{a}))-Y(a,\bar{a})Y(\bar{a},a)}.\end{split}$
It remains to study the entries of $X$ and $Y$. Observe that $(\underline{1}-I_{2})B$ is the matrix obtained from $B$ by interchanging its two rows. Since $B_{2}=\underline{0}$ and $C^{1}=\underline{0}^{t}$, we deduce the following:
$\begin{split}X(a,a)&=Y(\bar{a},a)=B_{1}(I_{\Lambda}-\mathcal{E}^{\Lambda}(s))^{-1}C^{1}=0;\\ X(\bar{a},a)&=Y(a,a)=B_{2}(I_{\Lambda}-\mathcal{E}^{\Lambda}(s))^{-1}C^{1}=0;\\ X(\bar{a},\bar{a})&=Y(a,\bar{a})=B_{2}(I_{\Lambda}-\mathcal{E}^{\Lambda}(s))^{-1}C^{2}=0.\end{split}$
Moreover, by Theorem 5.12, Fact 5.13 and since $\mathcal{E}^{\Lambda}(s)=D$,
$\begin{split}X(a,\bar{a})=Y(\bar{a},\bar{a})&=B_{1}(I_{\Lambda}-\mathcal{E}^{\Lambda}(s))^{-1}C^{2}\\ &=(\beta+1)^{-s}\cdot\sum_{b\in o^{-1}(d)\cap E\Lambda}\omega(b)^{-s}e_{b}^{\Lambda}(I_{\Lambda}-\mathcal{E}^{\Lambda}(s))^{-1}(e_{d}^{\Lambda})^{t}\\ &=(\beta+1)^{-s}(\mathcal{Z}_{\Lambda,d\to d}(s)-1).\qed\end{split}$
The claim now follows by substitution and elementary algebraic manipulations.
Proposition 6.10.
Let $(\Gamma,\omega)$ satisfy Setting [$\Gamma$], and consider $a\in E\Gamma$ with $o(a)=t(a)=c$. Set $\omega(a)=\alpha+1$, $\omega(\bar{a})=\beta+1$, and consider the subgraph of $\Gamma$ given by $\Lambda:=\Gamma\setminus\{a,\bar{a}\}$. Then,
$\mathcal{Z}_{\Gamma,a\to a}(s)^{-1}=\frac{\xi_{1}(\alpha,\beta)\cdot\mathcal{Z}_{\Lambda,c\to c}(s)^{-1}-\eta(\alpha,\beta)}{\xi_{2}(\alpha,\beta)\cdot\mathcal{Z}_{\Lambda,c\to c}(s)^{-1}+\eta(\alpha,\beta)},$
where
$\begin{split}\xi_{1}(\alpha,\beta)&=1-\Big{(}\alpha^{-s}-(\alpha+1)^{-s}\Big{)}\Big{(}\beta^{-s}-(\beta+1)^{-s}\Big{)};\\ \xi_{2}(\alpha,\beta)&=\Big{(}1+\alpha^{-s}-(\alpha+1)^{-s}\Big{)}\Big{(}1+\beta^{-s}-(\beta+1)^{-s}\Big{)};\\ \eta(\alpha,\beta)&=(\alpha^{-s}+1)(\beta+1)^{-s}+(\alpha+1)^{-s}(\beta^{-s}+1)-2(\alpha+1)^{-s}(\beta+1)^{-s}.\end{split}$
The picture below sketches the setting of Proposition 6.10. The edges of $\Lambda$ are dashed.
[…]
Remark 6.10.1.
In Proposition 6.10, if $\alpha=\beta$, after elementary manipulations the given formula becomes
$\mathcal{Z}_{\Gamma,a\to a}(s)^{-1}=\frac{(1-\alpha^{-s}+(\alpha+1)^{-s})\cdot\mathcal{Z}_{\Lambda,c\to c}(s)^{-1}-2(\alpha+1)^{-s}}{(1+\alpha^{-s}-(\alpha+1)^{-s})\cdot\mathcal{Z}_{\Lambda,c\to c}(s)^{-1}+2(\alpha+1)^{-s}}.$
Proof of Proposition 6.10.
The strategy of the proof is analogous to the one of Proposition 6.9. Thus, we keep the same notation and proof structure, and we only specify what needs to be changed. First, instead of (6.33), the rows $B_{1}$ and $B_{2}$ of $B$ and the columns $C^{1}$ and $C^{2}$ of $C$ are the following:
(6.37) $\begin{split}B_{1}&=B_{2}=\sum_{b\in o^{-1}(c)\cap E\Lambda}\omega(b)^{-s}e_{b}^{\Lambda};\\ C^{1}&=\sum_{b\in t^{-1}(c)\cap E\Lambda}(\alpha+1)^{-s}(e_{b}^{\Lambda})^{t}=(\alpha+1)^{-s}(e_{c}^{\Lambda})^{t};\\ C^{2}&=\sum_{b\in t^{-1}(c)\cap E\Lambda}(\beta+1)^{-s}(e_{b}^{\Lambda})^{t}=(\beta+1)^{-s}(e_{c}^{\Lambda})^{t}.\end{split}$
Moreover, in (6.34) the only matrices that change are the following:
(6.38) $\begin{split}I_{2}-A&=\begin{bmatrix}1-(\alpha+1)^{-s}&-\beta^{-s}\\ -\alpha^{-s}&1-(\beta+1)^{-s}\end{bmatrix};\\ I_{2}+(\underline{1}-I_{2})A&=\begin{bmatrix}1+\alpha^{-s}&(\beta+1)^{-s}\\ (\alpha+1)^{-s}&1+\beta^{-s}\end{bmatrix}.\end{split}$
Analogously to (6.36), we deduce that
(6.39) $\begin{split}\mathcal{Z}_{\Gamma,a\to a}(s)^{-1}&=\frac{\det\Big{(}I_{2}-A-\overbrace{B(I_{\Lambda}-D)^{-1}C}^{=X}\Big{)}}{\det\Big{(}I_{2}+(\underline{1}-I_{2})A+\underbrace{(\underline{1}-I_{2})B(I_{\Lambda}-D)^{-1}C}_{=Y}\Big{)}}.\end{split}$
In this case $X=Y$, because $B_{1}=B_{2}$ and then $(\underline{1}-I_{2})B=B$. Moreover, recalling Theorem 5.12 and Fact 5.13, we have
(6.40) $\begin{split}X(a,a)&=X(\bar{a},a)=B_{1}(I_{\Lambda}-D)^{-1}C^{1}\\ &=(\alpha+1)^{-s}\cdot\sum_{b\in o^{-1}(c)\cap E\Lambda}\omega(b)^{-s}e_{b}^{\Lambda}(I_{\Lambda}-\mathcal{E}^{\Lambda}(s))^{-1}(e_{c}^{\Lambda})^{t}\\ &=(\alpha+1)^{-s}(\mathcal{Z}_{\Lambda,c\to c}(s)-1).\end{split}$
Similarly,
(6.41) $X(a,\bar{a})=X(\bar{a},\bar{a})=B_{1}(I_{\Lambda}-D)^{-1}C^{2}=(\beta+1)^{-s}(\mathcal{Z}_{\Lambda,c\to c}(s)-1).$
The statement now follows by (6.40), (6.41) and elementary algebraic manipulations. ∎
7. The behaviour at $s=-1$
The main goal of this section is to prove that the Euler–Poincaré identity for a unimodular t.d.l.c. group $G$ having a weakly locally $\infty$-transitive or (P)-closed action on a tree as prescribed by Corollary F. To achieve this result, we first give a formula of the relevant Euler–Poincaré characteristic in terms of local data of the action (cf. Proposition 7.3) and then use the splitting formulae of Section 6 to prove Theorem E and Corollary F. Finally, given an edge-weighted graph $(\Gamma,\omega)$, we exploit the machinery introduced in this paper to relate the behaviour of $\mathcal{Z}_{\Gamma,u\to u}(-1)$ with the behaviour at $1$ of a suitable weighted Ihara zeta function associated to $\Gamma$ (cf. Section 7.3).
7.1. The Euler–Poincaré characteristic
According to [8, §5], every unimodular t.d.l.c. group $G$ of type $\mathrm{FP}$ (with respect to the category of discrete left $\mathbb{Q}[G]$-modules) admits an Euler–Poincaré characteristic $\widetilde{\chi}_{G}$. For every compact open subgroup $K\leq G$, this invariant is a determined rational multiple of the Haar measure $\mu_{K}$ on $G$ normalised with respect to $K$, written $\widetilde{\chi}_{G}=\chi(G,\mu_{K})\cdot\mu_{K}$. By the uniqueness of the Haar measure on $G$ up to a positive real rescaling, for every Haar measure $\mu$ on $G$ there is a unique real number $\chi(G,\mu)$ such that
$\widetilde{\chi}_{G}=\chi(G,\mu)\cdot\mu.$
Moreover, if $\mu$ and $\mu^{\prime}=c\cdot\mu$ ($c\in\mathbb{R}_{>0}$) are Haar measures on $G$, then
(7.1) $\chi(G,\mu)=c\cdot\chi(G,\mu^{\prime}).$
In the present note, we focus on the case of t.d.l.c. groups acting on a tree with compact open vertex stabilisers and finite quotient graph. For those groups, their unimodularity and their Euler–Poincaré characteristic can be characterised in terms of local data of the action as shown in Proposition 7.1 and Proposition 7.3, respectively.
Proposition 7.1 ([3, Propositions 1.2 and 3.6], [6, §3.6]).
Let $G$ be a t.d.l.c. group acting on a tree $T$ with compact open vertex stabilisers. Let $\Gamma$ be the quotient graph of $(G,T)$, and denote by $\omega$ its standard edge weight. Then $G$ is unimodular if, and only if, for every closed path $(a_{1},\ldots,a_{n})$ in $\Gamma$ we have
(7.2) $\prod_{i=1}^{n}\omega(a_{i})=\prod_{i=1}^{n}\omega(\bar{a}_{i}).$
Remark 7.1.1.
Let $\mathfrak{p}=(a_{1},\ldots,a_{m})$ and $\mathfrak{q}=(b_{1},\ldots,b_{n})$ be reduced paths in $\Gamma$ with $o(a_{1})=o(b_{1})$ and $t(a_{m})=t(b_{n})$. Hence $N_{\mathrm{vert}}(\mathfrak{p})=\prod_{i=1}^{m}\omega(a_{i})$, $N_{\mathrm{edg}}(\mathfrak{p})=1$ if $m=1$ and $N_{\mathrm{edg}}(\mathfrak{p})=\prod_{i=2}^{m}\omega(a_{i})$ if $m\geq 2$. Similar observations hold for $\overline{\mathfrak{p}}$, $\mathfrak{q}$ and $\overline{\mathfrak{q}}$. By Proposition 7.1, we deduce that
$N_{\mathrm{vert}}(\mathfrak{p})N_{\mathrm{vert}}(\overline{\mathfrak{q}})=N_{\mathrm{vert}}(\mathfrak{q})N_{\mathrm{vert}}(\overline{\mathfrak{p}}).$
If in particular $a_{m}=b_{n}$, we also have
$N_{\mathrm{vert}}(\mathfrak{p})N_{\mathrm{edg}}(\overline{\mathfrak{q}})=N_{\mathrm{vert}}(\mathfrak{q})N_{\mathrm{edg}}(\overline{\mathfrak{p}}).$
Moreover, if $a_{1}=b_{1}$ and $a_{m}=b_{n}$ then
$N_{\mathrm{edg}}(\mathfrak{p})N_{\mathrm{edg}}(\overline{\mathfrak{q}})=N_{\mathrm{edg}}(\mathfrak{q})N_{\mathrm{edg}}(\overline{\mathfrak{p}}).$
Theorem 7.2 ([8, Theorem 5.6]).
Let $G$ be a unimodular t.d.l.c. group acting on a tree $T$ with compact open vertex stabilisers and finite quotient graph. Let $\mathcal{V}\subseteq VT$ and $\mathcal{E}^{+}\subseteq ET$ be sets of representatives for the $G$-orbits on $VT$ and on a fixed orientation $ET^{+}$ in $T$. Then, for every Haar measure $\mu$ on $G$,
$\chi(G,\mu)=\sum_{v\in\mathcal{V}}\frac{1}{\mu(G_{v})}-\sum_{e\in\mathcal{E}^{+}}\frac{1}{\mu(G_{e})}.$
From the hypotheses of Theorem 7.2 one deduces that $T$ is locally finite. In particular, in this case the quantity $\chi(G,\mu)$ coincides with the Euler–Poincaré characteristic of $G$ with respect to $\mu$ as defined in [16, Definition 4.8].
Proposition 7.3.
Let $G$ be a unimodular t.d.l.c. group acting on a tree $T$ with compact open vertex stabilisers, finite quotient graph $\Gamma$, and such that $(G,T)$ is weakly locally $\infty$-transitive or (P)-closed. Let $\omega$ be the standard edge weight on $\Gamma$, and let $N_{\mathrm{vert}}=N_{\mathrm{vert}}^{\omega}$, $N_{\mathrm{edg}}=N_{\mathrm{edg}}^{\omega}$ be as in Definition 4.4. Let $c\in V\Gamma$ and $\Lambda\subseteq\Gamma$ be a maximal subtree, and consider an orientation $E\Lambda^{+}$ in $\Lambda$ such that the restricted origin map $o\colon E\Lambda^{+}\to V\Lambda\setminus\{c\}$ is a bijection. Let also $E\Gamma^{+}$ be an arbitrary orientation in $\Gamma$ such that $E\Gamma^{+}\cap\Lambda=E\Lambda^{+}$.
Then, for every $v\in VT$ with $\pi(v)=c$, we have
(7.3) $\chi(G,\mu_{G_{v}})=1+\sum_{a\in E\Gamma^{+}\cap E\Lambda}(1-\omega(a))\frac{N_{\mathrm{vert}}(\mathfrak{p}_{c,o(a)})}{N_{\mathrm{vert}}(\overline{\mathfrak{p}_{c,o(a)}})}-\sum_{b\in E\Gamma^{+}\setminus E\Lambda}\frac{N_{\mathrm{vert}}(\mathfrak{q}_{c,b})}{N_{\mathrm{edg}}(\overline{\mathfrak{q}_{c,b}})},$
for arbitrary reduced paths $\mathfrak{p}_{c,o(a)}\in\mathcal{P}^{\mathrm{lift}}_{\Gamma,v}(c\to o(a))$ and $\mathfrak{q}_{c,b}\in\mathcal{P}^{\mathrm{lift}}_{\Gamma,v}(c\to b)$, for all $a\in E\Gamma^{+}\cap E\Lambda$ and $b\in E\Gamma^{+}\setminus E\Lambda$.
Remark 7.3.1.
In the following, we comment on the choices made in the statement of Proposition 7.3.
  • (i)
    Let $\Gamma$ be a finite tree. For every $c\in V\Gamma$, there is an orientation $E\Gamma^{+}$ for which the origin map restricts to a bijection $o\colon E\Gamma^{+}\to V\Gamma\setminus\{c\}$.
    Indeed, let $\mathcal{E}^{+}$ be an arbitrary orientation in $\Gamma$ and set
    $E\Gamma^{+}:=\{a\mid a\in\mathcal{E}^{+},\,o(a)\neq c\}\sqcup\{\bar{a}\mid a\in\mathcal{E}^{+},\,o(a)=c\}.$
    Then $E\Gamma^{+}$ is an orientation. Moreover, since $o(a)\neq t(a)$ for every $a\in E\Gamma$, the origin map in $\Gamma$ restricts to a map $o\colon E\Gamma^{+}\to V\Gamma\setminus\{c\}$. By [20, §I.2, Proposition 12] we have $|E\Gamma^{+}|=|V\Gamma|-1$ and then $o\colon E\Gamma^{+}\to V\Gamma\setminus\{c\}$ is bijective.
  • (ii)
    By Remark 7.1.1, the right-hand side of (7.3) does not depend on the choice of specific reduced paths $\mathfrak{p}_{c,o(a)}$ and $\mathfrak{q}_{c,b}$ from $c$ to $o(a)$ and from $c$ to $b$, respectively.
  • (iii)
    A formula analogous to the one in (7.3) holds for $\chi(G,\mu_{G_{e}})$, $e\in ET$. Indeed, since $\mu_{G_{e}}=|G_{v}:G_{e}|\cdot\mu_{G_{v}}=\omega(\pi(e))\cdot\mu_{G_{v}}$, by (7.1) we have
    $\chi(G,\mu_{G_{e}})=\omega(\pi(e))^{-1}\chi(G,\mu_{G_{o(e)}}).$
    Then Proposition 7.3 applies.
Proof of Proposition 7.3.
Let $\pi\colon T\to\Gamma$ be the quotient map and consider a set of representatives $\mathcal{E}^{+}\subseteq ET$ for $E\Gamma^{+}$. Up to replace elements of $E\Gamma^{+}\setminus E\Lambda$ with their reverse, we may assume that for every $e\in\mathcal{E}^{+}$ with $\pi(e)\not\in E\Lambda$ the geodesic from $v$ to $e$ is defined in $T$. Since $o\colon E\Lambda^{+}\to V\Gamma\setminus\{c\}$ is bijective, notice that $\mathcal{V}:=\{v\}\sqcup\{o(e)\mid e\in\mathcal{E}^{+}\text{ and }\pi(e)\in E\Lambda\}$ is a set of representatives for $V\Gamma$. Moreover, $\mu_{G_{v}}(G_{v})=1$ and, for every $e\in\mathcal{E}^{+}$,
$\mu_{G_{v}}(G_{o(e)})=|G_{o(e)}:G_{e}|\cdot\mu_{G_{v}}(G_{e})=\omega(\pi(e))\cdot\mu_{G_{v}}(G_{e}).$
By Theorem 7.2, we have
$\begin{split}\chi(G,\mu_{G_{v}})&=1+\sum_{e\in\mathcal{E}^{+}:\atop\pi(e)\in E\Lambda^{+}}\Bigg{(}\frac{1}{\mu_{G_{v}}(G_{o(e)})}-\frac{1}{\mu_{G_{v}}(G_{e})}\Bigg{)}-\sum_{e\in\mathcal{E}^{+}:\atop\pi(e)\in E\Gamma^{+}\setminus\Lambda}\frac{1}{\mu_{G_{v}}(G_{e})}\\ &=1+\sum_{e\in\mathcal{E}^{+}:\atop\pi(e)\in E\Lambda^{+}}(1-\omega(\pi(e)))\frac{1}{\mu_{G_{v}}(G_{o(e)})}-\sum_{e\in\mathcal{E}^{+}:\atop\pi(e)\in E\Gamma^{+}\setminus\Lambda}\frac{1}{\mu_{G_{v}}(G_{e})}.\end{split}$
Let $e\in\mathcal{E}^{+}$ and set $\pi(e)=a$. For $t\in\{o(e),e\}$, consider the geodesic $[v,t]=(e_{1},\dots,e_{n})$ in $T$ lifting $\mathfrak{p}_{c,o(a)}$ if $t=o(e)$, and lifting $\mathfrak{q}_{c,a}$ if $t=e$. Then,
(7.4) $\begin{split}\frac{1}{\mu_{G_{v}}(G_{t})}&=\frac{\overbrace{|G_{v}:G_{v}\cap G_{t}|\cdot\mu_{G_{v}}(G_{v}\cap G_{t})}^{=\mu_{G_{v}}(G_{v})=1}}{\underbrace{|G_{t}:G_{v}\cap G_{t}|\cdot\mu_{G_{v}}(G_{v}\cap G_{t})}_{=\mu_{G_{v}}(G_{t})}}=\frac{|G_{v}:G_{[v,t]}|}{|G_{t}:G_{\overline{[v,t]}}|}\\ &=\frac{|G_{v}\cdot e_{1}|}{|G_{t}\cdot\bar{e}_{n}|}\prod_{k=1}^{n-1}\frac{|G_{(e_{1},\ldots,e_{k})}\cdot e_{k+1}|}{|G_{(\bar{e}_{n},\ldots,\bar{e}_{k+1})}\cdot\bar{e}_{k}|}.\end{split}$
For the latter equality in (7.4), see (4.5) and (4.6). Note that $|G_{v}\cdot e_{1}|=\omega(\pi(e_{1}))$. Moreover, $|G_{t}\cdot\bar{e}_{n}|=1$ if $t=e$ (because $e_{n}=t$) and $|G_{t}\cdot\bar{e}_{n}|=\omega(\pi(\bar{e}_{n}))$ if $t=o(e)$ (because $o(\bar{e}_{n})=t$). For $11\leq k\leq n-$ 1 we claim that
(7.5) $|G_{(e_{1},\ldots,e_{k})}\cdot e_{k+1}|=|G_{e_{k}}\cdot e_{k+1}|\quad\text{and}\quad|G_{(\bar{e}_{n},\ldots,\bar{e}_{k+1})}\cdot\bar{e}_{k}|=|G_{\bar{e}_{k+1}}\cdot\bar{e}_{k}|.$
If $(G,T)$ is (P)-closed, (7.5) follows from (4.7). If $(G,T)$ is weakly locally $\infty$-transitive, Remark 3.9.1(ii) and Proposition 4.6 yield
$|G_{(e_{1},\ldots,e_{k})}\cdot e_{k+1}|=N_{\mathrm{edg}}(\pi(e_{k}),\pi(e_{k+1}))=|G_{e_{k}}\cdot e_{k+1}|.$
A similar argument holds for $|G_{(\bar{e}_{n},\ldots,\bar{e}_{k+1})}\cdot\bar{e}_{k}|$.
For $11\leq k\leq n-$ 1, we now prove that
(7.6) $\frac{|G_{e_{k}}\cdot e_{k+1}|}{|G_{\bar{e}_{k+1}}\cdot\bar{e}_{k}|}=\frac{\omega(\pi(e_{k+1}))}{\omega(\pi(\bar{e}_{k}))}.$
To see this, set $v_{k}=t(e_{k})$ and $H=G_{e_{k}}\cap G_{e_{k+1}}$. Since $G_{e_{k}}=G_{\bar{e}_{k}}$ and $G_{e_{k+1}}=G_{\bar{e}_{k+1}}$, we have
$|G_{v_{k}}:H|=|G_{v_{k}}:G_{e_{k+1}}|\cdot|G_{\bar{e}_{k+1}}:G_{(\bar{e}_{k+1},\bar{e}_{k})}|=\omega(\pi(e_{k+1}))\cdot|G_{\bar{e}_{k+1}}\cdot\bar{e}_{k}|$
and, at the same time,
$|G_{v_{k}}:H|=|G_{v_{k}}:G_{\bar{e}_{k}}|\cdot|G_{e_{k}}:G_{(e_{k},e_{k+1})}|=\omega(\pi(\bar{e}_{k}))\cdot|G_{e_{k}}\cdot e_{k+1}|.$
Combining (7.4), (7.5) and (7.6), we deduce that
(7.7) $\frac{1}{\mu_{G_{v}}(G_{t})}=\frac{\omega(\pi(e_{1}))}{|G_{t}\cdot\bar{e}_{n}|}\prod_{i=1}^{n-1}\frac{\omega(\pi(e_{k+1}))}{\omega(\pi(\bar{e}_{k}))},$
where $|G_{t}\cdot\bar{e}_{n}|$ equals $1$ if $t=e$, and it equals $\omega(\pi(\bar{e}_{n}))$ if $t=o(e)$. By design, $\pi([v,t])=:(a_{1},\ldots,a_{n})$ is a reduced path in $\Gamma$. Then
(7.10) $N_{\mathrm{vert}}(\pi([v,t]))=\prod_{i=1}^{n}\omega(a_{i})\,\text{ and }\,N_{\mathrm{edg}}(\pi([v,t]))=\left\{\begin{array}[]{cl}1,&\text{if }n=1;\\ \prod_{i=2}^{n}\omega(a_{i}),&\text{if }n\geq 2.\end{array}\right.$
By (7.7) and (7.10), we conclude that
$\begin{split}\frac{1}{\mu_{G_{v}}(G_{o(e)})}&=\frac{N_{\mathrm{vert}}(\mathfrak{p}_{c,o(a)})}{N_{\mathrm{vert}}(\overline{\mathfrak{p}_{c,o(a)}})},\quad\forall\,e\in\mathcal{E}^{+}\text{ with }\pi(e)\in E\Lambda;\\ \frac{1}{\mu_{G_{v}}(G_{e})}&=\frac{N_{\mathrm{vert}}(\mathfrak{q}_{c,a})}{N_{\mathrm{edg}}(\overline{\mathfrak{q}_{c,a}})},\quad\forall\,e\in\mathcal{E}^{+}\text{ with }\pi(e)\not\in E\Lambda.\qed\end{split}$
Remark 7.3.2.
Let $(G_{1},T_{1})$ and $(G_{2},T_{2})$ be group actions on trees that satisfy the hypotheses of Proposition 7.3. Let $(\Gamma_{1},\omega_{1})$ and $(\Gamma_{2},\omega_{2})$ be the quotient graphs of $(G_{1},T_{1})$ and $(G_{2},T_{2})$ endowed with their standard edge weights, respectively. Assume there is a graph isomorphism $\varphi\colon\Gamma_{1}\to\Gamma_{2}$ such that $\omega_{2}(\varphi(a))=\omega_{1}(a)$ for every $a\in E\Gamma_{1}$. Let $v_{1}\in VT_{1}$ and $v_{2}\in VT_{2}$ be vertices satisfying $G_{1}\cdot v_{1}=c_{1}$ and $G_{2}\cdot v_{2}=\varphi(c_{1})$. By Proposition 7.3,
$\chi(G_{1},\mu_{(G_{1})_{v_{1}}})=\chi(G_{2},\mu_{(G_{2})_{v_{2}}}),$
where $\mu_{(G_{i})_{v_{i}}}$ is the Haar measure of $G_{i}$ normalised with respect to $(G_{i})_{v_{i}}$.
A notable consequence of Proposition 7.3 is that the value $\chi(G,\mu_{G_{v}})$ depends only on $(\Gamma,\omega)$. This suggests the following definition.
Definition 7.4.
Let $\Gamma$ be a finite connected non-empty graph with an edge weight $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$. Let $(T,\pi)$ be the universal cover of $(\Gamma,\omega)$, and set $G=\mathrm{Aut}_{\pi}(T)$ (cf. Example 1.1). The pair $(\Gamma,\omega)$ is said to be unimodular if $\mathrm{Aut}_{\pi}(T)$ is unimodular.
Let $(\Gamma,\omega)$ be unimodular. For all $c\in V\Gamma$ and $a\in E\Gamma$ and given arbitrary $v\in VT$ and $e\in ET$ satisfying $\pi(v)=c$ and $\pi(e)=a$, define
(7.11) $\chi(\Gamma,c):=\chi(G,\mu_{G_{v}})\quad\text{and}\quad\chi(\Gamma,a):=\chi(G,\mu_{G_{e}}).$
Since $G$ is unimodular, the assignments in (7.11) do not depend on the choice of $v\in\pi^{-1}(c)$ and of $e\in\pi^{-1}(a)$, respectively.
Remark 7.4.1.
Let $G$ be a unimodular t.d.l.c. group acting on a tree $T$ with compact open vertex-stabilisers and finite quotient graph $\Gamma$. Denote by $\omega$ the standard edge weight, and assume that $(G,T)$ is weakly locally $\infty$-transitive or (P)-closed. For every $t\in T$ with $G\cdot t=u$, from Proposition 7.3 we have
$\chi(\Gamma,u)=\chi(G,\mu_{G_{t}}).$
Example 7.5.
  • (i)
    Let $\Gamma$ be a $1$-segment with $E\Gamma=\{a,\bar{a}\}$. Since $\Gamma$ is a tree, $(\Gamma,\omega)$ is unimodular for every $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$ (cf. Proposition 7.1). Moreover, $\Gamma$ is its only maximal subtree. Consider an edge weight $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$. Set $c=o(a)$, $E\Gamma^{+}=\{\bar{a}\}$ and let $\mathfrak{p}_{c,o(\bar{a})}$ be the $1$-edge path $a$. Then Proposition 7.3 implies that
    $\chi(\Gamma,c)=1+(1-\omega(\bar{a}))\frac{\omega(a)}{\omega(\bar{a})}.$
    With a similar strategy one computes $\chi(\Gamma,t(a))$.
  • (ii)
    Let $\Gamma$ be a $n$-bouquet of loops based on the vertex $c$. Note that the $1$-point subgraph is the only maximal subtree of $\Gamma$. By Proposition 7.1, for an edge weight $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$ the pair $(\Gamma,\omega)$ is unimodular if, and only if, $\omega(a)=\omega(\bar{a})$ for every $a\in E\Gamma$. Provided $(\Gamma,\omega)$ is unimodular and $E\Gamma=\{a_{i},\bar{a}_{i}\mid 1\leq i\leq n\}$, from Proposition 7.3 we deduce that
    $\chi(\Gamma,c)=1-\sum_{i=1}^{n}\omega(a_{i}).$
Lemma 7.6.
Let $\Gamma$ be a finite connected non-empty graph, and let $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 2}$ be such that $(\Gamma,\omega)$ is unimodular. Then, for every $a\in E\Gamma$,
(7.12) $\chi(\Gamma,o(a))=\omega(a)\cdot\chi(\Gamma,a).$
Moreover, for all $c,d\in V\Gamma$,
(7.13) $\chi(\Gamma,c)=\frac{N_{\mathrm{vert}}(\mathfrak{p})}{N_{\mathrm{vert}}(\overline{\mathfrak{p}})}\chi(\Gamma,d),$
where $\mathfrak{p}$ is any reduced path in $\Gamma$ from $c$ to $d$. Similarly, for all $a,b\in E\Gamma$ for which there is a reduced path in $\Gamma$ from $a$ to $b$, we have
(7.14) $\chi(\Gamma,a)=\frac{N_{\mathrm{edg}}(\mathfrak{q})}{N_{\mathrm{edg}}(\overline{\mathfrak{q}})}\chi(\Gamma,b),$
where $\mathfrak{q}$ is any reduced path from $a$ to $b$ in $\Gamma$.
In Lemma 7.6, since $\Gamma$ is connected, replacing $a$ with $\bar{a}$ or $b$ with $\bar{b}$ if necessary, we can always find a reduced path from $a$ to $b$ in $\Gamma$.
Proof.
First, (7.12) follows from Remark 7.3.1. By Remark 7.1.1, the ratios in (7.13) and (7.14) do not depend on the choices of $\mathfrak{p}$ and $\mathfrak{q}$, respectively. Moreover, if we prove (7.13) and (7.14) for $\ell(\mathfrak{p})=1$ and $\ell(\mathfrak{q})=2$, the general statements follow iteratively. It remains to observe what follows. First, for every $a\in E\Gamma$ we have
(7.15) $\chi(\Gamma,o(a))=\omega(a)\cdot\chi(\Gamma,a)=\frac{\omega(a)}{\omega(\bar{a})}\chi(\Gamma,t(a)).$
Moreover, let $(a,b)$ is a length-$2$ reduced path in $\Gamma$ and set $t(a)=c=o(b)$. Then (7.12) and (7.15) imply that
$\chi(\Gamma,a)=\chi(\Gamma,\bar{a})=\frac{1}{\omega(\bar{a})}\chi(\Gamma,c)=\frac{\omega(b)}{\omega(\bar{a})}\chi(\Gamma,b).\qed$
Lemma 7.7.
Let $\Gamma$ be a finite connected non-empty graph with an edge weight $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$ such that $(\Gamma,\omega)$ is unimodular. Suppose that there are connected subgraphs $\Gamma_{1}$ and $\Gamma_{2}$ of $\Gamma$ such that $\Gamma=\Gamma_{1}\cup\Gamma_{2}$ and $\Gamma_{1}\cap\Gamma_{2}=\{c\}$, for some $c\in V\Gamma$. Then $(\Gamma_{i},\omega{|_{E\Gamma_{i}}})$ is unimodular for every $i\in\{1,2\}$, and
(7.16) $\chi(\Gamma,c)=\chi(\Gamma_{1},c)+\chi(\Gamma_{2},c)-1.$
Proof.
Let $\Lambda$ be a maximal subtree of $\Gamma$. We claim that $\Lambda_{i}:=\Gamma_{i}\cap\Lambda$ is a maximal subtree of $\Gamma_{i}$, for every $i\in\{1,2\}$. Clearly, both $\Lambda_{1}$ and $\Lambda_{2}$ are subtrees of $\Gamma_{1}$ and $\Gamma_{2}$, respectively. We prove the maximality for $i=1$, as for $i=2$ one may proceed analogously. For every subtree $\Xi_{1}\subseteq\Gamma_{1}$ with $\Xi_{1}\supseteq\Lambda_{1}$, we have $\Xi_{1}\cap\Lambda_{2}=\{c\}$ and thus $\Xi_{1}\cup\Lambda_{2}$ is a subtree of $\Gamma$ containing $\Lambda$. Hence $\Lambda=\Xi_{1}\cup\Lambda_{2}$ and
$\Lambda_{1}=\Gamma_{1}\cap\Lambda=(\Gamma_{1}\cap\Xi_{1})\cup(\Gamma_{1}\cap\Lambda_{2})=\Xi_{1}\cup\{c\}=\Xi_{1}.$
Consider an orientation $E\Gamma^{+}$ in $E\Gamma$ such that the origin map in $\Gamma$ restricts to a bijection $o\colon E\Gamma^{+}\cap E\Lambda\to V\Gamma\setminus\{c\}$ (cf. Remark 7.3.1). For every $i\in\{1,2\}$, the set $E\Gamma_{i}^{+}:=E\Gamma^{+}\cap E\Gamma_{i}$ is an orientation in $E\Gamma_{i}$ and the origin map in $\Gamma_{i}$ restricts to a bijection $o_{i}\colon E\Gamma_{i}^{+}\cap E\Lambda_{i}\to V\Gamma_{i}\setminus\{c\}$. By Proposition 7.3, we conclude that
$\begin{split}\chi(\Gamma,c)&=1+\sum_{i=1}^{2}\Bigg{(}\sum_{a\in E\Gamma_{i}^{+}\cap E\Lambda_{i}}(1-\omega(a))\frac{N_{\mathrm{vert}}(\mathfrak{p}_{c,o(a)})}{N_{\mathrm{vert}}(\overline{\mathfrak{p}_{c,o(a)}})}-\sum_{a\in E\Gamma_{i}^{+}\setminus E\Lambda_{i}}\frac{N_{\mathrm{vert}}(\mathfrak{q}_{c,a})}{N_{\mathrm{edg}}(\overline{\mathfrak{q}_{c,a}})}\Bigg{)}\\ &=\chi(\Gamma_{1},c)+\chi(\Gamma_{2},c)-1.\end{split}$
7.2. The evaluation at $s=-1$ and the Euler–Poincaré characteristic
The goal of what follows is to prove Theorem E and Corollary F. In view of Theorem E, we first formulate a version of Lemma 7.6 for $\mathcal{Z}_{\Gamma,u\to u}(-1)^{-1}$.
Lemma 7.8.
Let $(\Gamma,\omega)$ be a unimodular edge-weighted graph satisfying Setting [$\Gamma$] and such that $\Gamma$ has no cycles of length $\geq 2$. Then, for every $a\in E\Gamma$,
(7.17) $\mathcal{Z}_{\Gamma,o(a)\to o(a)}(-1)^{-1}=\omega(a)\cdot\mathcal{Z}_{\Gamma,a\to a}(-1)^{-1}.$
Moreover, for all $c,d\in V\Gamma$ and all $a,b\in E\Gamma$ such there is a reduced path from $a$ to $b$ in $\Gamma$, we have
(7.20) $\begin{array}[]{c}\displaystyle{\mathcal{Z}_{\Gamma,c\to c}(-1)^{-1}=\frac{N_{\mathrm{vert}}(\mathfrak{p})}{N_{\mathrm{vert}}(\overline{\mathfrak{p}})}\mathcal{Z}_{\Gamma,d\to d}(-1)^{-1},}\\ \displaystyle{\mathcal{Z}_{\Gamma,a\to a}(-1)^{-1}=\frac{N_{\mathrm{edg}}(\mathfrak{q})}{N_{\mathrm{edg}}(\overline{\mathfrak{q}})}\mathcal{Z}_{\Gamma,b\to b}(-1)^{-1},}\end{array}$
where $\mathfrak{p}$ and $\mathfrak{q}$ are arbitrary reduced paths in $\Gamma$ from $c$ to $d$ and from $a$ to $b$, respectively.
Proof.
Once proved (7.17) (which is analogous to (7.12)), arguing as in the proof of Lemma 7.6 one can deduce (7.20). We first prove (7.17) for every $1$-loop $a$. Namely, let $a\in E\Gamma$ with $o(a)=t(a)=c$. Since $(\Gamma,\omega)$ is unimodular, note that $\omega(a)=\omega(\bar{a})$. Let $\Lambda$ be the graph obtained from $\Gamma$ removing $a$ and $\bar{a}$, and let $L_{a}$ be the subgraph of $\Gamma$ with $VL_{a}=\{c\}$ and $EL_{a}=\{a,\bar{a}\}$. By Remark 6.10.1, we have
(7.21) $\mathcal{Z}_{\Gamma,a\to a}(-1)^{-1}=\omega(a)^{-1}\cdot\Big{(}\mathcal{Z}_{\Lambda,c\to c}(-1)^{-1}-\omega(a)\Big{)}.$
Moreover, Proposition 6.4 and (5.19) yield
(7.22) $\mathcal{Z}_{\Gamma,c\to c}(-1)^{-1}=\mathcal{Z}_{\Lambda,c\to c}(-1)^{-1}-\omega(a).$
Combining (7.21) and (7.22), we deduce (7.17).
For all edges $a$ with $o(a)\neq t(a)$, the relation in (7.17) is proved by induction on $|E\Gamma|/2=:k(\Gamma)\geq 1$. If $k(\Gamma)=1$, then $\Gamma$ is a $1$-segment and (7.17) follows from (5.17). Let $k(\Gamma)\geq 2$ and assume that the claim holds for every graph $\Gamma^{\prime}$ with $k(\Gamma^{\prime})<k(\Gamma)$. Let $a\in E\Gamma$ be such that $o(a)=:c\neq d:=t(a)$. If $o^{-1}(c)=\{a\}$, then Proposition 6.9 directly implies the claim. In case that $o^{-1}(d)=\{\bar{a}\}$, let $\Lambda$ be the graph obtained from $\Gamma$ by removing $a$ and $\bar{a}$. Then Proposition 6.9 yields
(7.23) $\mathcal{Z}_{\Gamma,a\to a}(-1)^{-1}=\mathcal{Z}_{\Gamma,\bar{a}\to\bar{a}}(-1)^{-1}=\frac{\mathcal{Z}_{\Lambda,c\to c}(-1)^{-1}}{\omega(a)}-\frac{\omega(\bar{a})-1}{\omega(\bar{a})}.$
On the other hand, let $\Gamma_{a}$ denote the $1$-segment subgraph of $\Gamma$ with $E\Gamma_{a}=\{a,\bar{a}\}$. By Proposition 6.4 and Example 5.14,
(7.24) $\begin{split}\mathcal{Z}_{\Gamma,c\to c}(-1)^{-1}&=\mathcal{Z}_{\Lambda,c\to c}(-1)^{-1}+\mathcal{Z}_{\Gamma_{a},c\to c}(-1)^{-1}-1\\ &=\mathcal{Z}_{\Lambda,c\to c}(-1)^{-1}+\omega(a)\frac{\omega(\bar{a})-1}{\omega(\bar{a})}.\end{split}$
Hence (7.17) follows from (7.23) and (7.24). Finally, assume that both $|o^{-1}(c)|\geq 2$ and $|o^{-1}(d)|\geq 2$. Denote by $\Xi_{1}$ and $\Xi_{2}$ be the connected components of $\Gamma\setminus\{a,\bar{a}\}$ containing $c$ and $d$, respectively. There are exactly two connected components because $\Gamma$ has no cycles of length $\geq 2$. Since $|o^{-1}(c)|\geq 2$ and $|o^{-1}(d)|\geq 2$, both $E\Xi_{1}$ and $E\Xi_{2}$ are non-empty. Moreover, $E\Gamma=E\Xi_{1}\sqcup\{a,\bar{a}\}\sqcup E\Xi_{2}$. For $i\in\{1,2\}$, let $\Gamma_{i}$ be the smallest subgraph of $\Gamma$ containing $\Lambda_{i}\cup\{a,\bar{a}\}$, and note that $k(\Gamma_{i})<k(\Gamma)$. Let also $\Gamma_{a}$ be the $1$-segment subgraph with edge set $\{a,\bar{a}\}$, and observe that $\Gamma_{1}\cap\Gamma_{2}=\Gamma_{a}$. Moreover, if $\Lambda_{1}=\Xi_{1}$ and $\Lambda_{2}=\Gamma_{2}$, we have $\Gamma=\Lambda_{1}\cup\Lambda_{2}$, $\Lambda_{1}\cap\Lambda_{2}=\{c\}$ and $\Lambda_{i}\subseteq\Gamma_{i}$ for every $i\in\{1,2\}$. Hence, Corollary 6.5 and Corollary 6.8 imply
$\begin{split}\mathcal{Z}_{\Gamma,c\to c}(-1)^{-1}&=\mathcal{Z}_{\Gamma_{1},c\to c}(-1)^{-1}+\mathcal{Z}_{\Gamma_{2},c\to c}(-1)^{-1}-\mathcal{Z}_{\Gamma_{a},c\to c}(-1)^{-1};\\ \mathcal{Z}_{\Gamma,a\to a}(-1)^{-1}&=\mathcal{Z}_{\Gamma_{1},a\to a}(-1)^{-1}+\mathcal{Z}_{\Gamma_{2},a\to a}(-1)^{-1}-\mathcal{Z}_{\Gamma_{a},a\to a}(-1)^{-1}.\end{split}$
The induction hypothesis now yields (7.17). ∎
By Lemma 7.6 and Lemma 7.8, we deduce the following.
Corollary 7.9.
Let $(\Gamma,\omega)$ be a unimodular edge-weighted graph satisfying Setting [$\Gamma$] and such that $\Gamma$ has no cycles of length $\geq 2$. If $\chi(\Gamma,u)=\mathcal{Z}_{\Gamma,u\to u}(-1)^{-1}$ for some $u\in\Gamma$, then $\chi(\Gamma,u)=\mathcal{Z}_{\Gamma,u\to u}(-1)^{-1}$ for every $u\in\Gamma$.
Proof.
Let $u\in\Gamma$. Since $\Gamma$ is connected, for all $c\in V\Gamma$ and $a\in E\Gamma$ there are reduced paths $\mathfrak{p}\in\mathcal{P}_{\Gamma}(U\to c)$ and $\mathfrak{q}\in\mathcal{P}_{\Gamma}(U\to A)$ (cf. Notation 2.1). Moreover, note that $\mathcal{Z}_{\Gamma,b\to b}(s)=\mathcal{Z}_{\Gamma,\bar{b}\to\bar{b}}(s)$ and $\chi(\Gamma,b)=\chi(\Gamma,\bar{b})$ for every $b\in E\Gamma$. Hence we may assume that $\mathfrak{p}\in\mathcal{P}_{\Gamma}(u\to c)$ and $\mathfrak{q}\in\mathcal{P}_{\Gamma}(u\to a)$. The statement follows from Lemma 7.6 and Lemma 7.8. ∎
Proof of Theorem E.
By Corollary 7.9, it suffices to prove that
(7.25) $\mathcal{Z}_{\Gamma,c\to c}(-1)^{-1}=\chi(\Gamma,c)$
for some vertex $c\in V\Gamma$. We prove it by induction on $|E\Gamma|/2=:k(\Gamma)\geq 1$. Let first $k(\Gamma)=1$, i.e., $E\Gamma=\{a,\bar{a}\}$. If $o(a)\neq t(a)$, Example 5.14 and Example 7.5(i) yield
$\mathcal{Z}_{\Gamma,c\to c}(-1)^{-1}=1+(1-\omega(\bar{a}))\frac{\omega(a)}{\omega(\bar{a})}=\chi(\Gamma,c).$
If $o(a)=t(a)$, from (5.19) and Example 7.5(ii) we deduce that
$\mathcal{Z}_{\Gamma,c\to c}(-1)^{-1}=1-\omega(a)=\chi(\Gamma,c).$
Let now $k(\Gamma)\geq 2$ and assume that the statement holds for all graphs $\Gamma^{\prime}$ with $k(\Gamma^{\prime})<k(\Gamma)$. Without loss of generality, we may take $c\in V\Gamma$ such that $|o^{-1}(c)|\geq 2$. Note that this vertex exists because $k(\Gamma)\geq 2$. By Lemma 6.6, there are proper connected subgraphs $\Lambda_{1}$ and $\Lambda_{2}$ of $\Gamma$ such that $\Gamma=\Lambda_{1}\cup\Lambda_{2}$ and $\Lambda_{1}\cap\Lambda_{2}=\{c\}$. Then Proposition 6.4 and Lemma 7.7 yield the claim. ∎
In view of the proof of Corollary F, we observe what follows.
Lemma 7.10.
Let $\Delta=(\Gamma,(X_{a}),(G(c)))$ be a local action diagram. Let $G=U(\Delta,\iota,c_{0})$ and $T=T(\Delta,\iota,c_{0})$ be as in Setting [(P)-cl]. Let $\tilde{\Delta}=(\Gamma,(X_{a}),(\tilde{G}(c)))$ be a local action diagram such that $\tilde{G}:=U(\tilde{\Delta},\iota,c_{0})$ acts weakly locally $\infty$-transitively on $T$. Then, for all $t_{1},t_{2}\in T$ we have
$\zeta_{G,G_{t_{1}},G_{t_{2}}}(-1)=\zeta_{\tilde{G},\tilde{G}_{t_{1}},\tilde{G}_{t_{2}}}(-1).$
By Example 3.12(i), note that $\tilde{\Delta}$ as in Lemma 7.10 exists.
Proof.
Both $(G,T)$ and $(\tilde{G},T)$ are (P)-closed actions on trees satisfying Setting [(P)-cl]. The statement now follows by applying Theorem 5.19 to both $\zeta_{G,G_{t_{1}},G_{t_{2}}}(s)$ and $\zeta_{\tilde{G},\tilde{G}_{t_{1}},\tilde{G}_{t_{2}}}(s)$. In detail, in both cases one checks that the matrices $\mathcal{F}(0)$ and $\mathcal{Y}_{\pi(t_{1}),\pi(t_{2})}(0)$ involved in the statement of Theorem 5.19, as well as the integer $\kappa_{\pi(t_{1})}(\pi(t_{2}))$, depend only on $X=\bigsqcup_{a\in E\Gamma}X_{a}$, on $\Gamma$ and its standard edge weight $\omega$, and on the inversion map $\iota$. The latter quantities do not vary by passing from $\Delta$ to $\tilde{\Delta}$, and the statement follows. ∎
Proof of Corollary F.
By Remark 5.15.1 and Lemma 7.10, it suffices to prove the statement for $(G,T)$ being weakly locally $\infty$-transitive. In this case, the claim follows from Remark 5.9.2, Remark 7.4.1 and Theorem E. ∎
7.3. The behaviour at $s=-1$ and the Ihara zeta function of a weighted graph
In [9, §3], a generalisation of the classical Ihara zeta function has been defined for every finite graph $\Gamma$ with a transition weight. Although it is not necessary here, we mention that the finiteness hypothesis on $\Gamma$ can be relaxed. According to [9, Definition 3.3], a transition weight on a finite graph $\Gamma$ is a map $W\colon E\Gamma\times E\Gamma\longrightarrow\mathbb{R}_{\geq 0}$ such that, whenever $W(a,b)\neq 0$, then $t(a)=o(b)$.
According to the definition of graph in [9] (cf. [9, Definition 3.1]), every edge is supposed to be uniquely determined by its endpoints. However, for the results involved below, this hypothesis has no influence and thus we do not assume it.
The Ihara zeta function $Z_{(\Gamma,W)}(x)$ of $(\Gamma,W)$ has been defined in [9, Definition 3.8] as a suitable infinite product of meromorphic functions on $\mathbb{C}$ converging for all $x\in\mathbb{C}$ with $|x|\ll 1$. Here we only need the following characterisation of the reciprocal of $Z_{(\Gamma,W)}(x)$, cf. [9, Theorem 3.11]. Namely,
(7.26) $Z_{(\Gamma,W)}(x)^{-1}=\det(I-xT),$
where $I$ is the identity matrix of dimension $|E\Gamma|$ and $T=[T(a,b)]_{a,b\in E\Gamma}\in\mathrm{Mat}_{n}(\mathbb{R})$ is defined as $T(a,b)=W(a,b)$ for all $a,b\in E\Gamma$ (assuming to have set a total order on $E\Gamma$). The matrix $T$ is called the Bass operator of $(\Gamma,W)$ (cf. [9, Definition 3.10]). Note that (7.26) gives a meromorphic continuation of $Z_{(\Gamma,W)}(x)$ to $\mathbb{C}$.
Example 7.11.
Let $\Gamma$ be a finite graph with an edge weight $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 2}$. Let $N_{\mathrm{edg}}=N_{\mathrm{edg}}^{\omega}$ and $\mathcal{E}(s)$ be as in Definition 4.4 and Definition 5.10, respectively. Consider the map $W=W_{(\Gamma,\omega)}\colon E\Gamma\times E\Gamma\to\mathbb{Z}_{\geq 0}$ defined by
$W(a,b):=\mathcal{E}(-1)(a,b)=\left\{\begin{array}[]{cl}N_{\mathrm{edg}}(a,b),&\text{if }t(a)=o(b);\\ 0,&\text{otherwise}.\end{array}\right.$
Then $W$ yields a transition weight on $\Gamma$. Note that assuming that $\omega(a)\geq 2$ is necessary to have $W(\bar{a},a)\neq 0$, for all $a\in E\Gamma$. In particular, by (7.26) we have
$Z_{(\Gamma,W)}(x)^{-1}=\det(I-x\mathcal{E}(-1)).$
Theorem 7.12.
Let $(\Gamma,\omega)$ be an edge-weighted graph satisfying Setting [$\Gamma$]. Let $\Gamma_{1},\Gamma_{2}$ be subgraphs of $\Gamma$ satisfying $\Gamma=\Gamma_{1}\cup\Gamma_{2}$ and such that $\Gamma_{1}\cap\Gamma_{2}$ is a $1$-segment with edge set $\{a,\bar{a}\}$. Assume also that $t(a)$ and $o(a)$ are terminal vertices in $\Gamma_{1}$ and $\Gamma_{2}$, respectively. Let $W$, $W_{1}$ and $W_{2}$ be the transition weights defined in Example 7.11 on $\Gamma$, $\Gamma_{1}$ and $\Gamma_{2}$, respectively. Then,
$\frac{\mathcal{Z}_{\Gamma,a\to a}(-1)}{\mathcal{Z}_{\Gamma_{1},a\to a}(-1)\cdot\mathcal{Z}_{\Gamma_{2},a\to a}(-1)}=\frac{1}{\omega(a)\omega(\bar{a})}\cdot\frac{Z_{(\Gamma,W)}(1)}{Z_{(\Gamma_{1},W_{1})}(1)\cdot Z_{(\Gamma_{2},W_{2})}(1)}.$
Proof.
Denote by $\mathcal{E}(-1)$, $\mathcal{E}_{1}(-1)$, $\mathcal{E}_{2}(-1)$ the Bass operators of $\Gamma$, $\Gamma_{1}$ and $\Gamma_{2}$ at $-1$. Let also $I,I_{1}$ and $I_{2}$ denote the identity matrices with complex entries of dimension $|E\Gamma|$, $|E\Gamma_{1}|$ and $|E\Gamma_{2}|$, respectively. By Theorem 5.12 and Example 7.11,
(7.27) $\mathcal{Z}_{\Gamma,a\to a}(-1)=Z_{(\Gamma,W)}(1)\cdot\det(I-M).$
where
$M=[M(h,k)]_{h,k\in E\Gamma}:=\mathcal{E}(-1)-\mathcal{U}_{a,a}(-1)=(I-(e_{a}+e_{\bar{a}})^{t}(e_{a}+e_{\bar{a}}))\mathcal{E}(-1).$
For all $h,k\in E\Gamma$, observe that
(7.28) $\begin{array}[]{ccl}M(h,k)&=&e_{h}\Big{(}I-(e_{a}+e_{\bar{a}})^{t}(e_{a}+e_{\bar{a}})\Big{)}\mathcal{E}(-1)e_{k}^{t}\\ &=&e_{h}\mathcal{E}(-1)e_{k}^{t}-e_{h}(e_{a}^{t}+e_{\bar{a}}^{t})\Big{(}(e_{a}+e_{\bar{a}})\mathcal{E}(-1)e_{k}^{t}\Big{)}\\ &=&\mathcal{E}(-1)(h,k)-\mathbbm{1}_{\{a,\bar{a}\}}(h)\Big{(}\mathcal{E}(-1)(a,k)+\mathcal{E}(-1)(\bar{a},k)\Big{)}.\end{array}$
Similarly, for every $i\in\{1,2\}$ we have
(7.29) $\mathcal{Z}_{\Gamma_{i},a\to a}(-1)=Z_{(\Gamma_{i},W_{i})}(1)\cdot\det(I_{i}-M_{i}),$
where $M_{i}=[M_{i}(h,k)]_{h,k\in E\Gamma_{i}}$ is the $|E\Gamma_{i}|$-dimensional given by
(7.30) $M_{i}(h,k)=\mathcal{E}_{i}(-1)(h,k)-\mathbbm{1}_{\{a,\bar{a}\}}(h)\Big{(}\mathcal{E}_{i}(-1)(a,k)+\mathcal{E}_{i}(-1)(\bar{a},k)\Big{)}.$
By (7.28) and (7.30), for every $i\in\{1,2\}$ we deduce that
(7.31) $M(h,k)=M_{i}(h,k),\quad\forall\,h,k\in E\Gamma_{i}.$
Let $\tilde{M}_{1}:=[M(h,k)]_{h,k\in E\Gamma_{1}\setminus\{\bar{a}\}}$ and $\tilde{M}_{2}:=[M(h,k)]_{h,k\in E\Gamma_{2}\setminus\{a\}}$. Set also $\tilde{I}_{1}$ and $\tilde{I}_{2}$ be the identity matrices in $\mathrm{Mat}_{|E\Gamma_{1}|-1}(\mathbb{C})$ and $\mathrm{Mat}_{|E\Gamma_{2}|-1}(\mathbb{C})$, respectively. We claim that $M$, $M_{1}$ and $M_{2}$ have the following decompositions in diagonal blocks:
(7.33) $\begin{array}[]{c}M=\begin{bmatrix}\tilde{M}_{1}&0\\ 0&\tilde{M}_{2}\end{bmatrix};\,M_{1}=\begin{bmatrix}\tilde{M}_{1}&0\\ 0&M(\bar{a},\bar{a})\end{bmatrix};\,M_{2}=\begin{bmatrix}M(a,a)&0\\ 0&\tilde{M}_{2}\end{bmatrix}.\end{array}$
Before proving (7.33), we use it to conclude the argument. From (7.33) we deduce that
(7.34) $\begin{split}\det(I-M)&=\det(\tilde{I}_{1}-\tilde{M}_{1})\cdot\det(\tilde{I}_{2}-\tilde{M}_{2})\\ &=\frac{\det(I_{1}-M_{1})\cdot\det(I_{2}-M_{2})}{(1-M(\bar{a},\bar{a})(1-M(a,a)))}.\end{split}$
Moreover, (7.28) yields $M(a,a)=-\mathcal{E}(-1)(\bar{a},a)=1-\omega(a)$ and $M(\bar{a},\bar{a})=-\mathcal{E}(-1)(a,\bar{a})=1-\omega(\bar{a})$. Combining (7.27), (7.29) and (7.34) we conclude the statement.
It remains to prove (7.33). By (7.30), it suffices to show that $M(h,k)=0$ if either $(h,k)\in(E\Gamma_{1}\setminus\{\bar{a}\})\times(E\Gamma_{2}\setminus\{a\})$ or $(h,k)\in(E\Gamma_{2}\setminus\{a\})\times(E\Gamma_{1}\setminus\{\bar{a}\})$. Recall that the only edge of $\Gamma_{1}$ (resp. $\Gamma_{2}$) ending at $t(a)$ (resp. $o(a)$) is $a$ (resp. $\bar{a}$). Hence, if $h\in E\Gamma_{1}\setminus\{a,\bar{a}\}$ then $t(h)\in V\Gamma_{1}\setminus\{t(a)\}$ and every $k\in E\Gamma_{2}\setminus\{a\}$ satisfies $o(k)\in V\Gamma_{2}\setminus\{o(a)\}$. Since $V\Gamma_{1}\setminus\{t(a)\}$ and $V\Gamma_{2}\setminus\{o(a)\}$ are disjoint, for such $h$ and $k$ we have $t(h)\neq o(k)$ and (7.28) implies that $M(h,k)=\mathcal{E}(-1)(h,k)=0$. Similarly, if $h\in E\Gamma_{2}\setminus\{a,\bar{a}\}$ and $k\in E\Gamma_{1}\setminus\{\bar{a}\}$ we have $M(h,k)=\mathcal{E}(-1)(h,k)=0$. Moreover, $M(a,k)=-\mathcal{E}(-1)(\bar{a},k)=0$ for every $k\in E\Gamma_{2}\setminus\{a\}$ as $t(\bar{a})\neq o(k)$. Similarly, $M(\bar{a},k)=-\mathcal{E}(-1)(a,k)=0$ for every $k\in E\Gamma_{1}\setminus\{\bar{a}\}$. ∎